### v1

# Extragradient methods with complexity guarantees for hierarchical variational inequalities

Abstract
In the framework of a real Hilbert space we consider the problem of approaching solutions to a class of hierarchical variational inequality problems, subsuming several other problem classes including certain mathematical programs under equilibrium constraints, constrained min-max problems, hierarchical game problems, optimal control under VI constraints, and simple bilevel optimization problems. For this general problem formulation, we establish rates of convergence in terms of suitably constructed gap functions, measuring feasibility gaps and optimality gaps. We present worst-case iteration complexity results on both levels of the variational problem, as well as weak convergence under a geometric weak sharpness condition on the lower level solution set. Our results match and improve the state of the art in terms of their iteration complexity and the generality of the problem formulation.
1 Introduction
Solving hierarchical equilibrium problems is an increasingly active area in mathematical programming that consists in finding a solution to an outer variational inequality over the solutions of an inner variational inequality. In this paper we consider the following problem. Let $(\mathcal{Z},\langle\cdot,\cdot\rangle)$ be a real Hilbert space, potentially infinite-dimensional. We consider the problem of solving the hierarchical equilibrium problem, in which a hemi-variational inequality ($\operatorname{HVI}$) is solved over the solution set of another mixed variational inequality. Specifically, the class of equilibrium problems we consider is of the form
 $\text{ Find }z^{\ast}\in\mathcal{S}_{2}\text{ s.t. }\langle{\mathsf{F}}_{1}(z^{\ast}),z-z^{\ast}\rangle+g_{1}(z)-g_{1}(z^{\ast})\geq 0\qquad\forall z\in\mathcal{S}_{2}\triangleq\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2}).$ (P)
where ${\mathsf{F}}_{1},{\mathsf{F}}_{2}:\mathcal{Z}\to\mathcal{Z}$ are monotone operators and $g_{1},g_{2}$ are proper, convex, and lower-semicontinuous functions. We refer to the set $\mathcal{S}_{2}\triangleq\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2})$ (assumed to be non-empty) as the solution set of the lower-level problem. In the same vein, we will call $\mathcal{S}_{1}$ the set of solutions of (P). Obviously, these sets are nested in the sense $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$. Problem (P) is the formulation of a system of nested $\operatorname{HVI}$’s (also known as variational inequalities of the second kind). These families of variational problems are prominent in mechanics [goeleven2013variational] and optimal control [Barbu93, de2011optimal]. In finite dimensions, it has become a useful model template to study a plethora of problems in mathematical programming, machine learning, operations research, game theory, and signal processing [malitsky2020golden, InertialFBF, Facchinei:2014aa]. Some motivating examples are given below.
1.1 Motivating Examples
1.1.1 Simple bilevel optimization
[…]
This model problem has been studied in several very recent papers. Important references include [merchav2023convex], who approach this problem via a disciplined proximal splitting ansatz, as well as [AminiYous19, DoronShtern23, yousefian2021bilevel, merchav2024fast, boct2025accelerating].
1.1.2 Hierarchical variational inequalities
If $g_{1}=0$ and $g_{2}=\operatorname{\iota}_{\mathcal{K}}$ for a non-empty, closed convex set $\mathcal{K}\subseteq\mathcal{Z}$, problem (P) reduces to
$\text{Find }z^{\ast}\in\mathcal{S}_{2}\qquad\text{s.t.: }\langle{\mathsf{F}}_{1}(z^{\ast}),z-z^{\ast}\rangle\geq 0\hskip 17.00024pt\forall z\in\mathcal{S}_{2}=\operatorname{zer}({\mathsf{F}}_{2}+\operatorname{\mathsf{NC}}_{\mathcal{K}}).$
This is the problem investigated in the recent papers [samadi_improved_2025, marschner2025tikhonov] and [alves2025inertial].
1.1.3 Equilibrium selection in Nash games
Selecting solutions from a given variational inequality problem is another important application of the model template (P). If ${\mathsf{F}}_{1}(x)=\nabla f_{1}(x)$, then we recover the optimal equilibrium selection problem as
$\min_{x\in\mathcal{S}}\{f_{1}(x)+g_{1}(x)\},\qquad\mathcal{S}=\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2}).$
[…]
1.1.4 Hierarchical jointly convex generalized Nash equilibrium problem
In very recent work, [LamSagSIOPT25] studied a projection-based method for solving a challenging class of jointly convex generalized Nash equilibrium problems (GNEPs) featuring hierarchy and non-smooth data. We first show that their model is a special case of the problem (P).
The lower level Nash equilibrium problem.
Let $\nu\in\{1,\ldots,N\}$ represent the labels of the lower-level players, who engage in a Nash game in which each player’s parameterized optimization problem is given by
$\min_{y^{\nu}\in\mathcal{Y}^{\nu}}\{h^{\ell}_{\nu}(y^{\nu},y^{-\nu}))+\varphi^{\ell}_{\nu}(y^{\nu})\}.$
The data of this game problem enjoy the following assumptions:
  1. (a)
    Player $\nu$’s feasible set $\mathcal{Y}^{\nu}$ is a real separable Hilbert space;
  2. (b)
[…]
  3. (c)
    The game’s pseudogradient ${\mathsf{F}}_{2}(y)=(\nabla_{y^{1}}h^{\ell}_{1}(y),\ldots,\nabla_{y^{N}}h^{\ell}_{N}(y))$ is monotone on $\mathcal{Y}\triangleq\prod_{\nu=1}^{N}\mathcal{Y}^{\nu}$;
  4. (d)
    $\varphi_{\nu}^{\ell}(\cdot)$ is proper closed convex with compact domain $\operatorname{dom}(\varphi^{\ell}_{\nu})\triangleq\mathcal{A}_{\nu}\subset\mathcal{Y}_{\nu}$.
We can characterize equilibria of the Nash game $\Gamma_{\rm low}=\{h^{\ell}_{\nu},\varphi_{\nu}^{\ell}\}_{1\leq\nu\leq N}$ as solutions to the mixed variational inequality $\mathcal{S}_{2}=\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2})$, where $g_{2}(y^{1},\ldots,y^{N})\triangleq\sum_{\nu=1}^{N}\varphi^{\ell}_{\nu}(y^{\nu}).$ The assumptions on the game problem ensure that $\operatorname{dom}(g_{2})$ is a compact subset of the separable Hilbert space $\mathcal{Y}$. Furthermore, $\mathcal{S}_{2}$ is nonempty, convex, and compact.
The upper level GNEP.
The upper level problem is the Nash game with joint coupling constraint represented by the equilibrium set $\mathcal{S}_{2}$. Let $\mu\in\{1,\ldots,M\}$ denote the labels of upper-level players. Each upper-level player aims to solve the constrained optimization problem
$\min_{x^{\mu}\in\mathcal{X}^{\mu}}\{h^{u}_{\mu}(x^{\mu},x^{-\mu})+\varphi^{u}_{\mu}(x^{\mu})\}\qquad\text{s.t.: }(x^{\mu},x^{-\mu})\in\mathcal{S}_{2}.$
Similar to the upper level NEP, we impose the following assumptions on the problem data:
  1. (d)
[…]
  2. (e)
    The operator ${\mathsf{F}}_{1}(x)=(\nabla_{x^{1}}h^{u}_{1}(x),\ldots,\nabla_{x^{N}}h^{u}_{N}(x))$ is monotone on $\mathcal{X}\triangleq\prod_{\nu=1}^{N}\mathcal{X}^{\nu}$;
  3. (f)
    $\varphi^{u}_{\mu}$ is proper closed convex with closed convex domain $\operatorname{dom}(\varphi^{u}_{\mu})$.
The decision variables $x^{\mu}$ correspond to blocks $(y^{\nu})_{\nu\in\mathcal{N}_{\mu}}$, in which $\mathcal{N}_{\mu}\subseteq\{1,\ldots,N\}$ represent the lower-level players controlled by upper-level player $\mu$. The blocks can well be overlapping, meaning that individual lower level players can have multiple upper level players. In particular, $\mathcal{X}^{\mu}=\prod_{\nu\in\mathcal{N}_{\mu}}\mathcal{Y}^{\nu}$. The entire hierarchical equilibrium problem thus belongs to our template (P), with $\mathcal{S}_{1}=\operatorname{zer}({\mathsf{F}}_{1}+\partial g_{1}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})$, where $g_{1}(x^{1},\ldots,x^{M})\triangleq\sum_{\mu=1}^{M}\varphi^{u}_{\mu}(x^{\mu}).$
1.2 Related works
This paper belongs to a subclass of bilevel problems in which a single decision variable is to be designed in order to optimize a function over the set of minimizerrs of another function. To emphasize this special character of such problems are also known as simple bilevel problems [Dempe:2021aa]. Starting with the seminal contribution of [Solodov:2007aa] and [Cabot2005], Tikhonov regularization has become one of the dominant paradigms in the numerical solution of simple bilevel optimization problems. In the context of the present paper, this means we have to consider the iteration-dependent operators
${\mathsf{V}}_{k}\triangleq{\mathsf{F}}_{2}+\sigma_{k}{\mathsf{F}}_{1},\qquad G_{k}\triangleq g_{2}+\sigma_{k}g_{1},$
where $k$ is the iteration counter and $(\sigma_{k})_{k}$ is a well-chosen sequence. Over the past years, a quite significant list of publications have appeared proving rates of convergence with respect to the inner (lower-level) and the outer (upper-level) problem. Specifically, building on the generalized Tikhonov regularization framework of [Solodov:2007aa], the first results in this direction have been produced in [beck2014first, SabSht17]. [merchav2023convex] develop a proximal subgradient method for convex composite models in the upper and lower level problem. Adaptive versions of a related proximal-gradient strategy are studied in [latafat2025convergence]. Accelerated simple bilevel optimization in the convex non-smooth composite setting have recently been derived in [merchav2024fast]. Beyond proximal-gradient based methods, projection-free schemes are developed in [DoronShtern23] and [giang2024projection]. Moving beyond optimization, there exists a significant literature, rooted in the numerical resolution of inverse problems and signal processing, which studied solution methods for selecting a solution to a variational inequality given a pre-defined selection criterion [ogura2002non, xu2003convergence, marino2011explicit, ono2014hierarchical]. This is known as the equilibrium selection problem and is in fact a special case of the general family of Mathematical Programs under equilibrium constraints (MPEC) [luo1996mathematical, Hintermuller:2014aa]. All these papers focus on asymptotic convergence of the sequence generated by tailor-made numerical algorithms. In this more enhanced setting, the only results on iteration complexity we are aware of are published in the papers [kaushik2023incremental, Kaushik:2021aa, jalilzadeh2024stochastic]. Moving beyond the function-case in the upper-level (equilibrium selection), there exist a few papers on numerical methods for solving nested and hierarchical variational inequalities [van2021regularization, ThongNumerical20, lampariello2022solution]. Among those, only [lampariello2022solution] contains complexity statements.
Recently, the papers [samadi_improved_2025] and [alves2025inertial] have addressed the finite-dimensional variational inequality case, obtained from our model template (P) by imposing $g_{1}=0$ and $g_{2}$ an indicator function over a compact convex set. Both papers derive complexity guarantees for the Korpelevich version of the extragradient method, applied to the Tikhonov regularized variational inequality with the operator ${\mathsf{F}}_{2}+\sigma_{k}{\mathsf{F}}_{1}$ at iteration $k$. The obtained rates are $O(1/k^{\delta})$ and $O(1/k^{1-\delta})$, where $\delta\in(0,1)$ is a parameter defining the strength of the Tikhonov regularization.
In a very recent work, [LamSagSIOPT25] studied a difficult class of game problems featuring hierarchy and non-smooth local cost functions. We show in Section 6 that their model falls into our general problem framework. Their algorithm is not a full splitting method, and thus not directly comparable to our operator splitting approach.
1.3 Contributions
The aim of this paper is to settle the complexity issue for a very general family of nested equilibrium problems formulated in the model framework (P), and thereby extend all the existing results on hierarchical monotone variational inequalities. Our main contributions are as follows:
  • More general problem formulation: Unlike [samadi_improved_2025, alves2025inertial], our hierarchical equilibrium model (P) contains general convex composite terms $g_{1},g_{2}$. Specifically, these papers treat the special case $g_{1}=0$ and $g_{2}=\operatorname{\iota}_{\mathcal{C}}$ for a compact convex and nonempty set $\mathcal{C}$ in a finite-dimensional euclidean vector space. Our analysis performed in potentially infinite-dimensional real Hilbert spaces, and considers general hemi-variational inequalities ($\operatorname{HVI}$’s). This is class contains a rather large class of variational models, which found significant application in optimal control and mechanics [Facchinei:2014aa, goeleven2013variational]. Besides the gain in modelling, adding these convex and potentially non-smooth functions is an important extension, since it allows us to cover the non-smooth composite convex model when specialized to the potential case, as studied in [merchav2024fast].
  • Enhanced Algorithmic design: The recent papers [samadi_improved_2025, alves2025inertial] approach the numerical resolution of problem (P) via double-call version of the extragradient method. Our approach instead uses the ideas of the optimistic extragradient method [hsieh2019convergence] and thus requires only one operator evaluation per iteration. Within this algorithmic setting, we derive sublinear convergence rates for the lower and upper level problem and improved rates under the strong monotonicity of ${\mathsf{F}}_{1}$. Specifically, using a polynomial regularization sequence $\sigma_{k}=O(k^{-\delta})$ for $\delta\in(0,1)$, we demonstrate upper complexity bounds on the order of $O(k^{-\delta})$ for the gap function associated with the lower level equilibrium problem, and a $O(k^{-(1-\delta)})$ complexity bound for the gap function of the entire hierarchical problem. This matches the existing bounds reported in [samadi_improved_2025, alves2025inertial], but for a much larger class of equilibrium problems, and fewer function evaluations (specifically, instead of 2 just 1 per iteration).
  • Unbounded domains: An important assumption in the complexity analysis in [samadi_improved_2025, alves2025inertial] is compactness of the domains over which the variational inequalities are solved. We abandon this assumption. Instead, our proof builds on ideas originating in the convergence analysis of non-autonomous evolution equations due to [attouch2018asymptotic]. In the potential case, these ideas have already been successfully applied to the study of hierarchical optimization problems [Bot:2014aa, Cabot2005]. In particular, [boct2025accelerating] make the role of this technique very transparent by emphasizing the role of the geometry of the lower level problem in establishing a unified complexity statement for both levels simultaneously. A methodological contribution of this work is to show how these ideas naturally translate to the variational setting. We belief that our proof technique is going to be useful also for other hierarchical equilibrium problems as well.
2 Facts on variational inequalities
2.1 Preliminaries
[…]
$\operatorname{\textsf{prox}}_{G}(u)=\operatorname*{argmin}_{z\in\mathcal{Z}}\{G(z)+\frac{1}{2}\norm{z-u}^{2}\}.$
If $x,u\in\mathcal{Z}$, we have
$x=\operatorname{\textsf{prox}}_{G}(u)\Leftrightarrow\langle u-x,z-x\rangle\leq G(z)-G(x)\hskip 17.00024pt\forall z\in\mathcal{Z}.$ (2.1)
The metric projection onto a nonempty closed convex set $\mathcal{C}\subset\mathcal{Z}$ is defined as $\Pi_{\mathcal{C}}(x)=\operatorname*{argmin}_{y\in\mathcal{C}}\norm{x-y}=\operatorname{\textsf{prox}}_{\operatorname{\iota}_{\mathcal{C}}}(x)$, where the convex indicator function $\operatorname{\iota}_{\mathcal{C}}:\mathcal{Z}\to(-\infty,\infty]$ is defined by
$\operatorname{\iota}_{\mathcal{C}}(x)\triangleq\left\{\begin{array}[]{cc}0&\text{if }x\in\mathcal{C},\\ +\infty&\text{ if }x\notin\mathcal{C}.\end{array}\right$.
We denote also $\operatorname{dist}(x,\mathcal{C})\triangleq\norm{x-\Pi_{\mathcal{C}}(x)}$. The polar cone attached to a closed convex set $\mathcal{C}\subset\mathcal{Z}$ is $\mathcal{C}^{\circ}\triangleq\{z\in\mathcal{Z}|\langle z,x\rangle\leq 0\qquad\forall x\in\mathcal{C}\}$. The conjugate function of $f\in\Gamma_{0}(\mathcal{Z})$ is $f^{\ast}(y)\triangleq\sup_{x\in\operatorname{dom}(f)}\{\langle y,x\rangle-f(x)\}.$ A mapping (operator) ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ is $\mu$-strongly-monotone ($\mu\geq 0$) if
$\langle{\mathsf{F}}(z_{1})-{\mathsf{F}}(z_{2}),z_{1}-z_{2}\rangle\geq\mu\norm{z_{1}-z_{2}}^{2}\hskip 17.00024pt\forall z_{1},z_{2}\in\mathcal{Z}.$
A 0-monotone operator is simply called monotone. ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ is $L$-Lipschitz continuous if
$\norm{{\mathsf{F}}(z_{1})-{\mathsf{F}}(z_{2})}\leq L\norm{z_{1}-z_{2}}\hskip 17.00024pt\forall z_{1},z_{2}\in\mathcal{Z}.$
The normal cone of a closed convex set $\mathcal{C}\subset\mathcal{Z}$ at point $z\in\mathcal{C}$ is defined $\operatorname{\mathsf{NC}}_{\mathcal{C}}(z)\triangleq\{\xi\in\mathcal{Z}|\langle\xi,z^{\prime}-z\rangle\leq 0\qquad\forall z^{\prime}\in\mathcal{C}\}$. Under convexity, the tangent cone $\operatorname{\mathsf{TC}}_{\mathcal{C}}(x)$ of $\mathcal{C}$ at $x\in\mathcal{C}$ is polar to the normal cone, and we have the explicit expression $\operatorname{\mathsf{TC}}_{\mathcal{C}}(x)=\operatorname{cl}\left(\bigcup_{\lambda>0}\frac{\mathcal{C}-x}{\lambda}\right)$. The support function of set $\mathcal{C}$ is defined as $\operatorname{\mathtt{s}}(p|\mathcal{C})\triangleq\sup_{z\in\mathcal{C}}\langle p,z\rangle.$ It follows from these definitions that
$\langle p,z\rangle=\operatorname{\mathtt{s}}(p|\mathcal{C})\hskip 17.00024pt\forall p\in\operatorname{\mathsf{NC}}_{\mathcal{C}}(z).$ (2.2)
Lemma 2.1 (Lemma 5.31, [BauCom16]).
[…]
$a_{k+1}\leq a_{k}-b_{k}+c_{k}.$
Then, $\lim_{k\to\infty}a_{k}$ exists and $\sum_{k\geq 1}b_{k}<\infty$.
2.2 Hemi-variational inequalities
Given a monotone and Lipschitz continuous mapping ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ and a function $g\in\Gamma_{0}(\mathcal{Z})$, the hemi-variational inequality problem $\operatorname{HVI}({\mathsf{F}},g)$ is to find $z^{\ast}\in\mathcal{Z}$ such that
$\langle{\mathsf{F}}(z^{\ast}),u-z^{\ast}\rangle+g(u)-g(z^{\ast})\geq 0\hskip 17.00024pt\forall u\in\mathcal{Z}.$ (2.3)
One can rewrite (2.3) as a monotone inclusion $0\in({\mathsf{F}}+\partial g)(z^{\ast})$. If $g=\operatorname{\iota}_{\mathcal{C}}$ for a closed convex set $\mathcal{C}\subset\mathcal{Z}$, we obtain from the above the variational inequality $\operatorname{VI}({\mathsf{F}},\mathcal{C})$:
$\text{Find }z^{\ast}\in\mathcal{C}\text{ s.t. }\langle{\mathsf{F}}(z^{\ast}),u-z^{\ast}\rangle\geq 0\hskip 17.00024pt\forall u\in\mathcal{C}.$
Let $\mathcal{C}\subset\operatorname{dom}(g)$ be a given compact subset of $\operatorname{dom}(g)$. A popular convergence measures for $\operatorname{HVI}({\mathsf{F}},g)$ is the localized gap function
$\Theta(z|{\mathsf{F}},g,\mathcal{C})\triangleq\sup_{y\in\mathcal{C}}\{\langle{\mathsf{F}}(y),z-y\rangle+g(z)-g(y)\}=\sup_{y\in\mathcal{C}}H^{({\mathsf{F}},g)}(z,y),$ (2.4)
where $H^{({\mathsf{F}},g)}:\mathcal{Z}\times\mathcal{Z}\to[-\infty,\infty]$ is defined as $H^{({\mathsf{F}},g)}(z,y)\triangleq\langle{\mathsf{F}}(y),z-y\rangle+g(z)-g(y)$. Historically, this localized version of a gap function can be traced back to [Minty62] (see also [kinderlehrer2000introduction, NesDual07]). Lemma 2.2 below summarizes its role for the numerical resolution of the problem $\operatorname{HVI}({\mathsf{F}},g)$ under continuity assumptions on $\mathsf{F}$}.
Lemma 2.2.
[…]
Proof.
See Appendix B. $\blacksquare$
2.3 Sharpness and error-bound property
Our geometric framework is phrased in terms of weak-sharpness of the lower-level solution set $\mathcal{S}_{2}$. Originally formulated for optimization problems in [burke1993weak], weak-sharpness in the context of variational inequalities has been defined in [patriksson1993unified]. Subsequently, implications in terms of error bounds of primal and dual gap functions have been stated in [marcotte1998weak, liu2016characterization]. The following definition of weak sharpness is from [Huang:2018aa].
Definition 2.3.
Let ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ be continuous and monotone over $\operatorname{dom}(g)\subset\mathcal{Z}$. Assume $\operatorname{dom}(g)$ is closed. The solution set $\mathcal{S}$ of $\operatorname{HVI}({\mathsf{F}},g)$ is weakly sharp if there exists $\tau>0$ such that
[…]
Based on [Huang:2018aa], we can give the following characterization of weak sharpness in terms of an error bound involving the dual gap function of $\operatorname{HVI}({\mathsf{F}},g)$.
Proposition 2.4.
Consider problem $\operatorname{HVI}({\mathsf{F}},g)$ with $\operatorname{dom}(g)$ a closed convex nonempty subset of $\mathcal{Z}$. If $\mathcal{S}=\operatorname{zer}({\mathsf{F}}+\partial g)$ is weakly sharp, then
$(\forall z\in\mathcal{D}_{g}):\qquad\Theta(z|{\mathsf{F}},g,\operatorname{dom}(g))\geq\tau\operatorname{dist}(z,\mathcal{S}).$
Proof.
See Appendix B. $\blacksquare$
This Proposition shows that weak sharpness implies that the gap function satisfies an error bound property [pang1997error]. Motivated by this fact, we propose the following definition:
Definition 2.5 (Weak Sharpness).
Let $\mathcal{S}$ be the nonempty solution set of $\operatorname{HVI}({\mathsf{F}},g)$. We say $\mathcal{S}$ is $(\alpha,\rho)$-weak sharp with $\alpha>0$ and $\rho\geq 1$ if
$(\forall z^{\ast}\in\mathcal{S})(\forall z\in\mathcal{D}_{g}=\operatorname{dom}(g)):\qquad H^{({\mathsf{F}},g)}(z,z^{\ast})\geq\alpha\rho^{-1}\operatorname{dist}(z,\mathcal{S})^{\rho}.$ (2.5)
Remark 2.1.
An important class of examples arises when $\rho=1$, notably in monotone linear complementarity problems in finite dimensions under nondegeneracy conditions [burke1993weak, pang1997error]. In the case where the $\operatorname{HVI}({\mathsf{F}},g)$ reduces to a convex optimization problem, weak-sharpness implies an Hölderian errror bound, an assumption already imposed by [Cabot2005] in the context of hierarchical minimization. Specifically, let us assume that ${\mathsf{F}}=0$ and $g\in\Gamma_{0}(\mathcal{Z})$. Then $\mathcal{S}=\operatorname*{argmin}g$, and acordingly, $H^{(0,g)}(z,z^{*})=g(z)-\min g$ for $z^{*}\in\mathcal{S}$. Hence, (2.5) implies
$\frac{\alpha}{\rho}\operatorname{dist}(u,\operatorname*{argmin}g)^{\rho}\leq g(u)-\min g\hskip 17.00024pt\forall u\in\mathcal{Z}.$
If $\rho=1$, this is the weak-sharpness condition of [burke1993weak]. The case $\rho=2$ corresponds to the "quadratic growth" condition of [drusvyatskiy2018error]. $\Diamond$
2.4 Constraint qualifications
The following set of assumptions shall be in place throughout the paper.
Assumption 1.
For $i=1,2$ the following properties do hold:
  • (i)
    ${\mathsf{F}}_{i}:\mathcal{Z}\to\mathcal{Z}$ is $L_{{\mathsf{F}}_{i}}$-Lipschitz continuous and monotone;
  • (ii)
    $g_{i}\in\Gamma_{0}(\mathcal{Z})$ and $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})\neq\varnothing$;
  • (iii)
    $\mathcal{S}_{2}\triangleq\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2})\neq\varnothing$.
We remark that Assumption 1 implies that $\mathcal{S}_{2}$ is a non-empty closed convex set. The next assumption is essentially a constraint qualification condition on the composite function
$x\mapsto f_{\bar{x}}(x)\triangleq\langle{\mathsf{F}}_{1}(\bar{x}),x-\bar{x}\rangle+g_{1}(x)-g_{1}(\bar{x})+\operatorname{\iota}_{\mathcal{S}_{2}}(x).$
Assumption 2.
The solution set $\mathcal{S}_{1}$ of (P) satisfies
$\mathcal{S}_{1}=\operatorname{zer}({\mathsf{F}}_{1}+\partial g_{1}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}).$
Similar assumptions have been made in [Bot:2014aa] for solving constrained variational inequalities. Assumption 2 is motivated by conditions ensuring a non-smooth sum rule of
$\partial_{x}\left(\langle{\mathsf{F}}_{1}(\bar{x}),\operatorname{Id}_{\mathcal{Z}}(\bullet)-\bar{x}\rangle+g_{1}(\bullet)-g_{1}(\bar{x})+\operatorname{\iota}_{\mathcal{S}_{2}}(\bullet)\right)(\bar{x})={\mathsf{F}}_{1}(\bar{x})+\partial g_{1}(\bar{x})+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}(\bar{x}).$
Common assumptions ensuring this property are [BauCom16, Corollary 16.48]:
  • $\operatorname{int}\operatorname{dom}(g_{1})\cap\mathcal{S}_{2}\neq\varnothing$;
  • $\operatorname{dom}(g_{1})\cap\operatorname{int}(\mathcal{S}_{2})\neq\varnothing$;
  • if $\mathcal{Z}$ is finite-dimensional and $\operatorname{ri}\operatorname{dom}(g_{1})\cap\operatorname{ri}\operatorname{dom}(g_{2})\neq\varnothing$.
Since $\mathcal{S}_{2}\subseteq\operatorname{dom}(\partial g_{2})\subset\operatorname{dom}(g_{2})$, these assumptions ensure the computationally pleasant identity $\partial(g_{1}+g_{2})=\partial g_{1}+\partial g_{2}$.
2.5 Gap functions for hierarchical variational inequalities
Since we are searching for a solution of a hierarchical system of variational inequalities, we have to introduce different merit functions for measuring the feasibility and the optimality of a test point $z\in\mathcal{Z}$.
Definition 2.6 (Feasibility gap).
We define the feasibility gap over a compact nonempty set $\mathcal{U}\subseteq\operatorname{dom}(g_{2})$ as
$\Theta_{\rm Feas}(z|\mathcal{U})\triangleq\Theta(z|{\mathsf{F}}_{2},g_{2},\mathcal{U})=\sup_{y\in\mathcal{U}}\{\langle{\mathsf{F}}_{2}(y),z-y\rangle+g_{2}(z)-g_{2}(y)\}.$ (2.6)
Definition 2.7 (Optimality gap).
The optimality gap over a compact set $\mathcal{U}\subseteq\operatorname{dom}(g_{1})$ with $\mathcal{U}\cap\mathcal{S}_{2}\neq\varnothing$ is defined as
$\Theta_{\rm Opt}(z|\mathcal{U}\cap\mathcal{S}_{2})\triangleq\Theta(z|{\mathsf{F}}_{1},g_{1},\mathcal{U}\cap\mathcal{S}_{2})=\sup_{y\in\mathcal{U}\cap\mathcal{S}_{2}}\{\langle{\mathsf{F}}_{1}(y),z-y\rangle+g_{1}(z)-g_{1}(y)\}.$ (2.7)
We note that if $z\in\mathcal{U}\cap\mathcal{S}_{2}$, and the regularity conditions stated in Lemma 2.2 hold, then by the same Lemma, the inequality $\Theta_{\rm Opt}(z|\mathcal{U}\cap\mathcal{S}_{2})\leq 0$ is equivalent to $z\in\mathcal{S}_{1}$. However, in general an inequality of the form $\Theta_{\rm Opt}(z|\mathcal{U}\cap\mathcal{S}_{2})\leq 0$, does not tell us much about the qualitative properties of the test point if $z\notin\mathcal{S}_{2}$. It is thus important to obtain a lower bound on the optimality gap.
Lemma 2.8.
Consider problem (P). Let Assumption 1 and 2 hold. Let $\mathcal{U}_{1}\subseteq\operatorname{dom}(g_{1})$ be a nonempty compact set with $\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$. Then, there exists a constant $B_{\mathcal{U}_{1}}>0$ such that
$\Theta_{\rm Opt}(z|\mathcal{U}_{1}\cap\mathcal{S}_{1})\geq-B_{\mathcal{U}_{1}}\operatorname{dist}(z,\mathcal{S}_{2}),\qquad\forall z\in\mathcal{Z}.$ (2.8)
Suppose $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp. Then for all nonempty and compact subsets $\mathcal{U}_{2}\subseteq\operatorname{dom}(g_{2})$ and all $z\in\operatorname{dom}(g_{2})$, we have
$\operatorname{dist}(z,\mathcal{S}_{2})\leq\left[\frac{\rho}{\alpha}\Theta_{\rm Feas}(z|\mathcal{U}_{2}\cap\mathcal{S}_{2})\right]^{1/\rho}.$ (2.9)
Proof.
See Appendix B. $\blacksquare$
Remark 2.2.
Lemma 2.8 allows us to establish a lower bound on the optimality gap over the solution set $\mathcal{S}_{2}$. It extends the corresponding result in [boct2025accelerating] to the operator case. [samadi_improved_2025] and [alves2025inertial] prove this result for special case of bilevel VIs on compact domains. $\Diamond$
3 Optimistic extragradient method
Our algorithmic design follows the popular Tikhonov regularization approach and employs a family of regularized problems with the data
${\mathsf{V}}_{\sigma}(z)\triangleq{\mathsf{F}}_{2}(z)+\sigma{\mathsf{F}}_{1}(z):\mathcal{Z}\to\mathcal{Z},\text{ and }G_{\sigma}(z)\triangleq g_{2}(z)+\sigma g_{1}(z)\in\Gamma_{0}(\mathcal{Z}).$
The parameter $\sigma\geq 0$ determines the relative importance of the upper level problem relative to the lower level problem. A decreasing sequence of regularization parameters $(\sigma_{k})_{k\geq 1}$ induces a corresponding sequence of operators ${\mathsf{V}}_{k}\triangleq{\mathsf{V}}_{\sigma_{k}}$ and convex functions $G_{k}\triangleq G_{\sigma_{k}}$. Being a sum of monotone and Lipschitz continuous mappings, ${\mathsf{V}}_{\sigma}$ is $L_{\sigma}\triangleq L_{{\mathsf{F}}_{2}}+\sigma L_{{\mathsf{F}}_{1}}$-Lipschitz and monotone. In the same vein, the combined mapping $G_{\sigma}$ is proper convex and lower semi-continuous. Furthermore, Assumption 1 guarantees the non-smooth sum rule $\partial G_{k}=\sigma_{k}\partial g_{1}+\partial g_{2}$. The conceptual algorithm we propose for solving (P) is described in Algorithm 1. The reader should consider this as a description of a method; Once concrete definitions for the step-size $(t_{k})_{k}$ and the regularization parameters $(\sigma_{k})_{k}$ are given, we obtain a bona-fide algorithm from the pseudo-code. Importantly, it should be observed that our algorithm requires only one evaluation of operators ${\mathsf{F}}_{1}$, ${\mathsf{F}}_{2}$ per iteration, in contrast to previous works [samadi_improved_2025, alves2025inertial].
Algorithm 1 Optimistic extragradient for hierarchical HVI’s (P)
$z^{1}=z^{1/2}$, step-size sequence $(t_{k})_{k\geq 1}$, non-increasing regularization sequence $(\sigma_{k})_{k\geq 1}$ satisfying
[…]
Set $k=1$
while stopping criterion not met do
   Compute
$z^{k+1/2}=\operatorname{\textsf{prox}}_{t_{k}G_{k}}(z^{k}-t_{k}({\mathsf{F}}_{2}(z^{k-1/2})+\sigma_{k}{\mathsf{F}}_{1}(z^{k-1/2}))),$
$z^{k+1}=\operatorname{\textsf{prox}}_{t_{k}G_{k}}(z^{k}-t_{k}({\mathsf{F}}_{2}(z^{k+1/2})+\sigma_{k}{\mathsf{F}}_{1}(z^{k+1/2})).$
end while
3.1 Statement of the main results
The main result of this paper are two complexity statements for the averaged iterates generated by Algorithm 1 in terms of the introduced gap functions. We work in a specific geometric setting, involving key quantities associated with the lower level problem $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$. Associated to the bifunction $H^{({\mathsf{F}}_{2},g_{2})}$, Appendix A defines the mapping
$\varphi^{({\mathsf{F}}_{2},g_{2})}(z,u)\triangleq\sup_{y\in\operatorname{dom}(g_{2})}\{H^{({\mathsf{F}}_{2},g_{2})}(z,y)+\langle y,u\rangle\}.$ (3.2)
This function encodes dual properties of the variational inequality, since it holds that (cf. eq. (A.1))
$\varphi^{({\mathsf{F}}_{2},g_{2})}(z,u)\leq\sup_{y\in\operatorname{dom}(g_{2})}\{\langle y,u\rangle-H^{({\mathsf{F}}_{2},g_{2})}(y,z)\}=\left(H^{({\mathsf{F}}_{2},g_{2})}(\bullet,z)\right)^{\ast}(u).$ (3.3)
The following summability condition is essentially due to [Attouch:2010aa, AttCzarPey11]:
Assumption 3.
[Attouch-Czarnecki condition] The step size sequence $(t_{k})_{k\geq 1}$ and the regularization sequence $(\sigma_{k})_{k\geq 1}$ satisfy
$(\forall p^{*}\in\text{Range}(\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})):\;\sum_{k=1}^{\infty}t_{k}\left[\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})\right]<\infty.$ (3.4)
To understand the meaning of Assumption 3, it is instructive to specialize our setting to the simple bilevel optimization case (cf. Section 1.1.1). In that case, the data of the lower level problem are identified with ${\mathsf{F}}_{2}=\nabla f_{2}$ and $\mathcal{S}_{2}=\operatorname*{argmin}_{z}(f_{2}+g_{2})(z)$, and thus
$H^{({\mathsf{F}}_{2},g_{2})}(x,y)$ $\leq f_{2}(x)-f_{2}(y)+g_{2}(x)-g_{2}(y)\triangleq\hat{f}_{2}(x)-\hat{f}_{2}(y)$
$=(\hat{f}_{2}-\min\hat{f}_{2})(x)-(\hat{f}_{2}-\min\hat{f}_{2})(y).$
Hence, $\varphi^{({\mathsf{F}}_{2},g_{2})}(z,p)\leq(\hat{f}_{2}-\min\hat{f}_{2})(z)+(\hat{f}_{2}-\min\hat{f}_{2})^{*}(p)$. Since for $z\in\mathcal{S}_{2}$ it holds $(\hat{f}_{2}-\min\hat{f}_{2})(z)=0$, this further implies
$\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p)-\operatorname{\mathtt{s}}(\sigma_{k}p|\mathcal{S}_{2})\leq(\hat{f}_{2}-\min\hat{f}_{2})^{*}(\sigma_{k}p)-\operatorname{\mathtt{s}}(\sigma_{k}p|\mathcal{S}_{2}).$
In [AttCzarPey11, Peypouquet:2012aa, boct2025accelerating] the following summability condition is imposed
$\sum_{k=1}^{\infty}t_{k}\left[(\hat{f}_{2}-\min\hat{f}_{2})^{*}(\sigma_{k}p)-\operatorname{\mathtt{s}}(\sigma_{k}p|\mathcal{S}_{2})\right]<\infty.$
We see that this condition is stronger than our condition (3.4).
Remark 3.1.
Assumption 3 looks quite daunting to verify, but as already observed in [boct2025accelerating], it fits very natural to the geometric setting of this paper. Indeed, let us assume that $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp, with $\alpha>0$ and $\rho>1$. According to Definition 2.5, for all $z\in\operatorname{dom}(g_{2})$ and for all $z^{*}\in\mathcal{S}_{2}$, we have $H^{({\mathsf{F}}_{2},g_{2})}(z,z^{*})\geq\alpha\rho^{-1}\operatorname{dist}(z,\mathcal{S}_{2})^{\rho}$. Hence, eq. (3.2) yields for all $z\in\mathcal{S}_{2}$
$\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})$ $\leq\left(H^{({\mathsf{F}}_{2},g_{2})}(\bullet,z)\right)^{\ast}(\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})$
$\leq\alpha^{-\frac{1}{\rho-1}}\left(\frac{\rho-1}{\rho}\right)\sigma_{k}^{\frac{\rho}{\rho-1}}\norm{p^{\ast}}^{\frac{\rho}{\rho-1}}.$
This shows that under the $(\alpha,\rho)$-weak sharpness condition, the summability condition (3.4) is satisfied whenever $\sum_{k\geq 1}t_{k}\sigma^{\frac{\rho}{\rho-1}}_{k}<\infty.$ $\Diamond$
Assumption 3 will allow us to prove that the sequence generated by the algorithm is bounded, an important step towards deriving convergence rates in terms of the averaged trajectory
$\bar{z}^{K}\triangleq\frac{1}{T_{K}}\sum_{k=1}^{K}t_{k}z^{k+1/2},\qquad T_{K}\triangleq\sum_{k=1}^{K}t_{k}.$ (3.5)
Theorem 3.1.
Consider problem (P). Let Assumptions 1-2 hold. Let additionally either Assumption 3 hold or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be compact. Consider the step size $(t_{k})_{k\geq 1}$ with $8t_{k}^{2}L_{k}^{2}\leq 1$ for all $k\geq 1$, and let $(\sigma_{k})_{k\geq 1}$ be a non-increasing sequence satisfying (3.1).
  • (i)
    Consider a nonempty compact set $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$. Then, there exists a constant $C_{\mathcal{U}_{2}}>0$ for which
    $\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2T_{K}}+\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}C_{\mathcal{U}_{2}}.$ (3.6)
    In particular, under (3.1) all weak accumulation points $\tilde{z}$ of $(\bar{z}^{k})$ satisfy $\Theta_{\rm Feas}(\tilde{z}|\mathcal{U}_{2})\leq 0$.
  • (ii)
    Let $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be a compact subset with $\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$. Then there exist constants $B_{\mathcal{U}_{1}},C_{\mathcal{U}_{1}}>0$ such that
    $B_{\mathcal{U}_{1}}\operatorname{dist}(\bar{z}^{K},\mathcal{S}_{2})\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{T_{K}\sigma_{K}}.$ (3.7)
  • (iii)
    If the lower level solution set $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp, then
    $B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2T_{K}(\alpha/\rho)}+\frac{C_{\mathcal{U}_{2}}\sum_{k=1}^{K}t_{k}\sigma_{k}}{(\alpha/\rho)T_{K}}\right]^{1/\rho}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{T_{K}\sigma_{K}}.$ (3.8)
    In particular, all weak accumulation points $\tilde{z}$ of $(\bar{z}^{k})$ satisfy $\Theta_{\rm Opt}(\tilde{z}|\mathcal{U}_{1}\cap\mathcal{S}_{2})=0$.
Remark 3.2.
Theorem 3.1 derives a rate in terms of the feasibility and the optimality gap of the hierarchical VI problem (P). These rates are established relative to arbitrary compact sets $\mathcal{U}_{1},\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$. Such bounds are meaningful because of the boundedness of the sequence $(z^{k})_{k\geq 1}$, a fact we establish with the derivation of (3.6). The constants involved in (3.6) and (3.7) depend on the Lipschitz modulus of the operators ${\mathsf{F}}_{i}$, and diameter-like constants, which can be exhibited thanks to the boundedness of the trajectory. $\Diamond$
Choosing a specific regularization sequence $(\sigma_{k})_{k\geq 1}$ allows us to turn the estimates from Theorem 3.1 into concrete complexity bounds. In detail, we consider the constant step size policy $t_{k}=t$ and regularization sequence
$\sigma_{k}\triangleq\frac{a}{(k+b)^{\delta}}\hskip 17.00024pta,b>0,\delta\in(0,1).$ (3.9)
Corollary 3.2.
Let the same Assumptions as in Theorem 3.1 be in place. Assume that the regularization sequence $(\sigma_{k})_{k\geq 1}$ is chosen according to (3.9), and $t_{k}=t\leq\frac{1}{\sqrt{8}L_{1}}$ for all $k\geq 1$. Then, we have
$\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt}+\frac{aC_{\mathcal{U}_{2}}}{(1-\delta)(K+b)^{\delta}},$ (3.10)
$-B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt(\alpha/\rho)}+\frac{aC_{\mathcal{U}_{2}}}{(\alpha/\rho)(1-\delta)(K+b)^{\delta}}\right]^{1/\rho}\overset{(*)}{\leq}\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{a(K+b)^{1-\delta}},$ (3.11)
where the inequality ()(*) holds under the $(\alpha,\rho)$-weakly sharpness assumption.
Remark $3.$
Under the regularization sequence (3.9) and if $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weak sharp, a sufficient condition for Assumption 3 to hold is $1>\delta>1-\frac{1}{\rho}$ (cf. Remark 3.1). This agrees with the analysis in the potential case of [boct2025accelerating]. $\Diamond$
4 Proofs of the main results
4.1 Energetic bounds
In this section, we develop the necessary energy bounds for the sequence $(z^{k})_{k\geq 1}$ generated by the optimistic extragradient method. Key is the step size condition $8t_{k}^{2}L^{2}_{k}\leq 1$, where $L_{k}\triangleq L_{{\mathsf{F}}_{2}}+\sigma_{k}L_{{\mathsf{F}}_{1}}$ is the combined Lipschitz constant of the operator ${\mathsf{V}}_{k}$. This condition implies $t_{k}\leq\frac{1}{\sqrt{8}}\frac{1}{L_{{\mathsf{F}}_{2}}+\sigma_{k}L_{{\mathsf{F}}_{1}}}\in\frac{1}{\sqrt{8}}\left[\frac{1}{L_{{\mathsf{F}}_{2}}+\sigma_{1}L_{{\mathsf{F}}_{1}}},\frac{1}{L_{{\mathsf{F}}_{2}}}\right].$
Lemma 4.1.
Let $(\sigma_{k})_{k\geq 1}$ be non-increasing and $(t_{k})_{k\geq 1}$ be a step size policy. Let also $z\in\mathcal{Z}$ be arbitrary.
  • (i)
    For all $k\geq 1$ we have
    $\begin{split}\frac{1}{2}\norm{z^{k+1}-z}^{2}&\leq\frac{1}{2}\norm{z^{k}-z}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}\\ &+t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).\end{split}$ (4.1)
  • (ii)
    If $8t^{2}_{k}L_{k}^{2}\leq 1$ then for all $k\geq 2$,
    $\begin{split}\frac{1}{2}\norm{z^{k+1}-z}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}&\leq\frac{1}{2}\norm{z^{k}-z}^{2}+\frac{t^{2}_{k-1}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\\ &-t_{k}\left(\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle+G_{k}(z^{k+1/2})-G_{k}(z)\right)\\ &-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}.\end{split}$ (4.2)
Proof.
(i) Applying (2.1) to the iterates $z^{k+1/2}$ and $z^{k+1}$, respectively, we obtain
$\langle z^{k+1}-z^{k},z^{k+1}-z\rangle\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1})\right)$ (4.3)
$\langle z^{k+1/2}-z^{k},z^{k+1/2}-z\rangle\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k-1/2}),z-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$
Using $z=z^{k+1}$ in the second inequality, we have
$\langle z^{k+1/2}-z^{k},z^{k+1/2}-z^{k+1}\rangle\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z^{k+1})-G_{k}(z^{k+1/2})\right).$ (4.4)
Now, we observe that
$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k+1}-z^{k}+z^{k}-z}^{2}$
$=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+2\langle z^{k+1}-z^{k},z^{k+1}-z\rangle.$
In the same spirit, we see
$\norm{z^{k+1}-z^{k}}^{2}=\norm{z^{k+1}-z^{k+1/2}}^{2}+\norm{z^{k+1/2}-z^{k}}^{2}+2\langle z^{k+1}-z^{k+1/2},z^{k+1/2}-z^{k}\rangle.$
Combining these two identities with (4.3) and (4.4), we can continue the estimation as
$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k+1/2}}^{2}-\norm{z^{k+1/2}-z^{k}}^{2}$
$-2\langle z^{k+1}-z^{k+1/2},z^{k+1/2}-z^{k}\rangle+2\langle z^{k+1}-z^{k},z^{k+1}-z\rangle$
$\leq\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k+1/2}}^{2}-\norm{z^{k+1/2}-z^{k}}^{2}$
$+2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1}\rangle$
$+2t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$ (4.5)
Adding and subtracting $z^{k+1/2}$ in the second inner product, leaves us with
$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k+1/2}}^{2}-\norm{z^{k+1/2}-z^{k}}^{2}$
$+2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle$
$+2t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$
Applying Young’s inequality $\langle a,b\rangle\leq\frac{c}{2}\norm{a}^{2}+\frac{1}{2c}\norm{b}^{2}$ with $c=t_{k}$, we obtain
$2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle\leq t^{2}_{k}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}+\norm{z^{k+1}-z^{k+1/2}}^{2}.$
Hence, we obtain (4.1)
$\frac{1}{2}\norm{z^{k+1}-z}^{2}$ $\leq\frac{1}{2}\norm{z^{k}-z}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$+t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$
(ii) Observe that for $L_{k}>0$,
$\frac{t_{k}^{2}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$ $=t^{2}_{k}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$\leq t_{k}^{2}L_{k}^{2}\norm{z^{k+1/2}-z^{k-1/2}}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$\leq 2t_{k}^{2}L^{2}_{k}\norm{z^{k+1/2}-z^{k}}^{2}+2t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}$
$-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2},$
where the first inequality follows from the $L_{k}$-Lipschitz continuity of ${\mathsf{V}}_{k}$, and the second inequality follows from $(a+b)^{2}\leq 2a^{2}+2b^{2}$. We can therefore continue,
$\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq 2t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$+\frac{1}{2}\left(4t_{k}^{2}L^{2}_{k}-1\right)\norm{z^{k+1/2}-z^{k}}^{2}$ (4.6)
$\leq 2t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2},$ (4.7)
where the last inequality is due to the step size choice satisfying $4t_{k}^{2}L^{2}_{k}\leq\frac{1}{2}$. Moreover, using the definition of $z^{k}$ and $z^{k-1/2}$, the non-expansiveness of the proximal mapping (LABEL:eq:nonexpansive), and the step size assumption $8t^{2}_{k}L_{k}^{2}\leq 1$ yields $4t^{2}_{k}t^{2}_{k-1}L^{2}_{k}\leq\frac{t^{2}_{k-1}}{2}$. This results in
$2t_{k}^{2}L_{k}^{2}\norm{z^{k}-z^{k-1/2}}^{2}\leq 2t^{2}_{k}t^{2}_{k-1}L^{2}_{k}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\leq\frac{t^{2}_{k-1}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}.$ (4.8)
Combining this inequality with (4.7), plugging the result into (4.1), and rearranging terms, we obtain (4.2)
$\frac{1}{2}\norm{z^{k+1}-z}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$ $\leq\frac{1}{2}\norm{z^{k}-z}^{2}+\frac{t^{2}_{k-1}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}$
$-t_{k}\left(\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle+G_{k}(z^{k+1/2})-G_{k}(z)\right)$
$-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}.$
$\blacksquare$
To continue with our energetic bounds, we define the data
$E_{k}(z)\triangleq\frac{1}{2}\norm{z^{k}-z}^{2}\hskip 17.00024pt\forall k\geq 1,$ (4.9)
$\Psi_{k}(z)\triangleq\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle+G_{k}(z^{k+1/2})-G_{k}(z)\hskip 17.00024pt\forall k\geq 1,$ (4.10)
$D_{k}\triangleq\frac{t^{2}_{k-1}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\hskip 17.00024pt\forall k\geq 2.$ (4.11)
This allows us to rewrite the statement in Lemma 4.1(ii) concisely as
$E_{k+1}(z)+D_{k+1}\leq E_{k}(z)+D_{k}-t_{k}\Psi_{k}(z)-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}\hskip 17.00024pt\forall k\geq 2.$ (4.12)
We now extend the validity of (4.12) to $k\geq 1$.
Lemma 4.2.
Define $D_{1}\triangleq 0$. Then, for all $k\geq 1$ relation (4.12) holds.
Proof.
It is only left to show that the inequality holds for $k=1$. For $k=1$ (4.12) reduces to
$E_{2}(z)+D_{2}\leq E_{1}(z)+D_{1}-t_{1}\Psi_{1}(z)-\frac{1}{4}\norm{z^{3/2}-z^{1}}^{2}.$ (4.13)
Using (4.7) we have that
$\frac{t_{1}^{2}}{2}\norm{{\mathsf{V}}_{1}(z^{1/2})-{\mathsf{V}}_{1}(z^{3/2})}^{2}-\frac{1}{2}\norm{z^{3/2}-z^{1}}^{2}$ $\leq 2t_{1}^{2}L^{2}_{1}\norm{z^{1}-z^{1/2}}^{2}-\frac{t^{2}_{1}}{2}\norm{{\mathsf{V}}_{1}(z^{1/2})-{\mathsf{V}}_{1}(z^{3/2})}^{2}-\frac{1}{4}\norm{z^{3/2}-z^{1}}^{2}$
$=-\frac{t^{2}_{1}}{2}\norm{{\mathsf{V}}_{1}(z^{1/2})-{\mathsf{V}}_{1}(z^{3/2})}^{2}-\frac{1}{4}\norm{z^{3/2}-z^{1}}^{2}$
where the last equality is due to $z^{1/2}=z^{1}$. Plugging this result back into (4.1), we arrive at
$\frac{1}{2}\norm{z^{2}-z}^{2}$ $\leq\frac{1}{2}\norm{z^{1}-z}^{2}-t_{1}\left(\langle{\mathsf{V}}_{1}(z^{3/2}),z^{3/2}-z\rangle+G_{1}(z^{3/2})-G_{1}(z)\right)-\frac{t_{1}^{2}}{2}\norm{{\mathsf{V}}_{1}(z^{1/2})-{\mathsf{V}}_{1}(z^{3/2})}^{2}$
$-\frac{1}{4}\norm{z^{3/2}-z^{1}}^{2}.$
Using the definition of $D_{1}$, we arrive at (4.13). $\blacksquare$
4.2 Towards proving Theorem 3.1
Establishing the rate on the feasibility gap (3.6)
To obtain a bound on the feasibility gap, we start with an Opial-like Lemma which holds under Assumption 3.
Lemma 4.3.
Let $z^{*}\in\mathcal{S}_{1}$ arbitrary and $(z^{k},z^{k+1/2})_{k\geq 1}$ be the sequence generated by Algorithm 1. Let also Assumption 3 hold. Define
$W_{k}(z^{*})\triangleq E_{k}(z^{*})+D_{k}\hskip 17.00024pt\forall k\geq 1.$ (4.14)
Then, the following statements hold:
  1. (a)
[…]
  2. (b)
    $\lim_{k\to\infty}\norm{z^{k+1/2}-z^{k}}=0$;
  3. (c)
    $(z^{k})_{k\geq 1}$ and $(z^{k+1/2})_{k\geq 1}$ are bounded.
Proof.
Pick $z^{\ast}\in\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$ so that there exists $p^{\ast}\in\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}(z^{\ast})$ such that ${\mathsf{F}}_{1}(z^{\ast})-p^{\ast}\in\partial g_{1}(z^{\ast})$. Using the monotonicity of ${\mathsf{F}}_{1}$ and the convex subgradient inequality for $g_{1}$, we obtain
$\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle$ $+g_{1}(z^{k+1/2})-g_{1}(z^{\ast})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle$
$\geq\langle{\mathsf{F}}_{1}(z^{\ast}),z^{k+1/2}-z^{\ast}\rangle+g_{1}(z^{k+1/2})-g_{1}(z^{\ast})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle\geq 0.$
Combined with the energy inequality (4.12) we conclude
$E_{k+1}(z^{\ast})+D_{k+1}$ $\leq E_{k+1}(z^{\ast})+D_{k+1}+t_{k}\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle+g_{1}(z^{k+1/2})-g_{1}(z^{\ast})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right)$
$\leq E_{k}(z^{\ast})+D_{k}-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}$
$-t_{k}\left(\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle+g_{2}(z^{k+1/2})-g_{2}(z^{\ast})-\langle\sigma_{k}p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right).$ (4.15)
Since $p^{\ast}\in\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}(z^{\ast})$, we know from (2.2) that $\langle\sigma_{k}p^{*},z^{*}\rangle=\operatorname{\mathtt{s}}(\sigma_{k}p^{*}|\mathcal{S}_{2})$ and
$-\left(\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle+g_{2}(z^{k+1/2})-g_{2}(z^{\ast})-\langle\sigma_{k}p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right)$
$=\langle\sigma_{k}p^{\ast},z^{k+1/2}\rangle+\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{\ast}-z^{k+1/2}\rangle+g_{2}(z^{\ast})-g_{2}(z^{k+1/2})-\langle\sigma_{k}p^{\ast},z^{\ast}\rangle$
$=\langle\sigma_{k}p^{\ast},z^{k+1/2}\rangle+H^{({\mathsf{F}}_{2},g_{2})}(z^{\ast},z^{k+1/2})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})$
$\leq\sup_{z\in\operatorname{dom}(g_{2})}\left(\langle\sigma_{k}p^{\ast},z\rangle+H^{({\mathsf{F}}_{2},g_{2})}(z^{\ast},z)\right)-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})$
$=\varphi^{({\mathsf{F}}_{2},g_{2})}(z^{\ast},\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})$
$\leq\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2}).$
We note that (A.2) in Appendix A guarantees that the upper bound is informative, in the sense that
$\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})\geq 0.$
Plugging this back into (4.15), we obtain
$E_{k+1}(z^{\ast})+D_{k+1}$ $\leq E_{k}(z^{\ast})+D_{k}-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}+t_{k}\left[\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})\right].$ (4.16)
We see from these estimates that the summability condition (3.4) gives a bound on the sequence $E_{k}(z^{\ast})+D_{k}$. Defining $W_{k}(z^{\ast})\triangleq E_{k}(z^{\ast})+D_{k}$, and
$a_{k}\triangleq t_{k}\left[\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})\right],$
we obtain from (4.16)
$W_{k+1}(z^{\ast})-W_{k}(z^{\ast})+\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}\leq a_{k}.$
From this, summability of $a_{k}$, and non-negativity of $W_{k}(z^{*})$ it follows from Lemma 2.1 that $\lim_{k\to\infty}W_{k}(z^{\ast})=W_{\infty}(z^{\ast})$ exists, and $\lim_{k\rightarrow\infty}\norm{z^{k+1/2}-z^{k}}=0$. By the definition of $E_{k}(z^{*})$, we further obtain that the sequences $(z^{k})_{k\geq 1}$ and $(z^{k+1/2})_{k\geq 1}$ are bounded. $\blacksquare$
[…]
We next consider a compact convex set $\mathcal{U}_{2}\subset\operatorname{dom}g_{1}\cap\operatorname{dom}g_{2}$ and choose as anchor point an arbitrary element $z\in\mathcal{U}_{2}$. Then, we have
$\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle$ $\leq\norm{{\mathsf{F}}_{1}(z^{k+1/2})}\cdot\norm{z-z^{k+1/2}}$ (4.17)
$\leq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+\norm{{\mathsf{F}}_{1}(z^{k+1/2})-{\mathsf{F}}_{1}(z^{1})}\right)\cdot\left(\norm{z^{1}-z^{k+1/2}}+\norm{z^{1}}+\norm{z}\right)$ (4.18)
$\leq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+L_{{\mathsf{F}}_{1}}R\right)\cdot(R+\norm{z^{1}}+\max_{z\in\mathcal{U}_{2}}\norm{z}),$ (4.19)
[…]
$C_{\mathcal{U}_{2}}$ $\triangleq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+L_{{\mathsf{F}}_{1}}R\right)(R+\norm{z^{1}}+\max_{z\in\mathcal{U}_{2}}\norm{z})+\Omega_{1}(\mathcal{U}_{2},R)$
to obtain
$\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2})\leq C_{\mathcal{U}_{2}},\qquad\forall z\in\mathcal{U}_{2},k\geq 1.$
It follows by the definition of $\Psi_{k}(z)$ and monotonicity of ${\mathsf{F}}_{2}$ that, for all $z\in\mathcal{U}_{2}$ and $k\geq 1$,
$\begin{split}\Psi_{k}(z)+\sigma_{k}C_{\mathcal{U}_{2}}&\geq\Psi_{k}(z)+\sigma_{k}(\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2}))\\ &=\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)\\ &\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z).\end{split}$ (4.20)
Multiplying by $t_{k}$ and using (4.12), where we now neglect the non-positive term $\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}$, we obtain, for all $z\in\mathcal{U}_{2},k\geq 1$,
$t_{k}(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z))$ $\leq t_{k}\Psi_{k}(z)+t_{k}\sigma_{k}C_{\mathcal{U}_{2}}$
$\leq(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1})+t_{k}\sigma_{k}C_{\mathcal{U}_{2}}.$ (4.21)
Summing (4.21) for $k=1,\ldots,K$ and dividing by $T_{K}$ yields, for all $z\in\mathcal{U}_{2}$,
$\frac{1}{T_{K}}\sum_{k=1}^{K}t_{k}\left(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)\right)\leq\frac{1}{T_{K}}(E_{1}(z)+D_{1})+\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}C_{\mathcal{U}_{2}}.$ (4.22)
Using the definition of the ergodic trajectory (3.5), we can apply the Jensen inequality in order to finally arrive at
$\langle{\mathsf{F}}_{2}(z),\bar{z}^{K}-z\rangle+g_{2}(\bar{z}^{K})-g_{2}(z)\leq\frac{1}{T_{K}}(E_{1}(z)+D_{1})+\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}C_{\mathcal{U}_{2}}\hskip 17.00024pt\forall z\in\mathcal{U}_{2}.$
Thus, for any compact subset $\mathcal{U}_{2}\subset\operatorname{dom}g_{1}\cap\operatorname{dom}g_{2}\subset\mathcal{Z}$ we can use the definition of the localized gap function (2.4) specialized as the feasibility gap (2.6) together with $D_{1}=0$, to obtain (3.6).
Moreover, under the step size condition (3.1) it follows that every weak accumulation point $\tilde{z}$ of the averaged trajectory $(\bar{z}^{k})$ satisfies
$\langle{\mathsf{F}}_{2}(z),\tilde{z}-z\rangle+g_{2}(\tilde{z})-g_{2}(z)\leq 0\hskip 17.00024pt\forall z\in\mathcal{U}_{2}$
In particular, $\Theta_{\rm Feas}(\tilde{z}|\mathcal{U}_{2})\leq 0$.
Proof of (3.10)
Choosing constant step $t_{k}=t$ and penalty sequence as in (3.9), we easily see that $\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}\leq\frac{a}{(1-\delta)(K+b)^{\delta}},$ so that the concrete complexity bound becomes
$\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt}+\frac{aC_{\mathcal{U}_{2}}}{(1-\delta)(K+b)^{\delta}}.$
Remark 4.1.
Corollary 3.2 is formulated for regularization sequences $(\sigma_{k})_{k\geq 1}$ of the form (3.9) with $\delta\in(0,1)$. In the limiting case $\delta=1$ and constant step size $t_{k}=t$, the complexity bound
$\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt}+\frac{a\log(K+b)}{K}C_{\mathcal{U}_{2}}.$
Hence, even in this limiting case, not satisfying condition (3.1) explicitly requested in Algorithm 1, the ergodic average $(\bar{z}^{k})_{k}$ asymptotically approaches the feasible set $\mathcal{S}_{2}$. $\Diamond$
Establishing the rate on the optimality gap (3.7)
Let $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be a compact set with $\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$. Using the monotonicity of ${\mathsf{F}}_{1}$ and ${\mathsf{F}}_{2}$, we see
$\Psi_{k}(z)$ $=\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)$
$+\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)$
$\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)$
$+\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right).$
Hence, when choosing $z\in\mathcal{S}_{2}$, we are able to continue the above bound as
$\Psi_{k}(z)$ $\geq\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)\qquad\forall k\geq 1.$
Plugging this in (4.12) and dropping the non-positive term $\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}$, we arrive at
$\sigma_{k}t_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)$ $\leq(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1})$
$=W_{k}(z)-W_{k+1}(z).$
To continue with our estimation, we proceed with the following simple result.
Lemma 4.4.
Let either Assumption 3 hold or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be compact. Then, the mapping $z\mapsto W_{k}(z)$ is bounded over bounded sets.
Proof.
Since
$W_{k}(z)$ $=\frac{1}{2}\norm{z^{k}-z}^{2}+\frac{t_{k}^{2}}{2}\norm{{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2})}^{2}\leq\frac{1}{2}\norm{z^{k}-z}^{2}+\frac{t^{2}_{k}L^{2}_{k}}{2}\norm{z^{k+1/2}-z^{k-1/2}}^{2}$
the result follows immediately from the boundedness of the sequences $(z^{k})_{k\geq 1}$, $(z^{k-1/2})_{k\geq 1}$ that hold both under Assumption 3 and when $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is compact. $\blacksquare$
Since $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$, it follows from the assumptions made that $\mathcal{U}_{1}\cap\mathcal{S}_{2}\neq\varnothing$. Therefore, we can pick $z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}$ arbitrary, so that we can find a positive constant $\bar{\omega}(z)$ such that $W_{k}(z)\leq\bar{\omega}(z)$ for all $k\geq 1$. Using this bound, and $\sigma_{k+1}\leq\sigma_{k}$, we arrive that
$\sum_{k=1}^{K}t_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)$ $\leq\frac{1}{\sigma_{1}}W_{1}(z)+W_{2}(z)(\frac{1}{\sigma_{2}}-\frac{1}{\sigma_{1}})+\ldots+W_{K}(z)(\frac{1}{\sigma_{K}}-\frac{1}{\sigma_{K-1}})$
$\leq\frac{\bar{\omega}(z)}{\sigma_{K}}.$
Dividing both sides by $T_{K}$ and applying the Jensen inequality, we obtain
$\langle{\mathsf{F}}_{1}(z),\bar{z}^{K}-z\rangle+g_{1}(\bar{z}^{K})-g_{1}(z)\leq\frac{\bar{\omega}(z)}{T_{K}\sigma_{K}}.$
Thus, for any compact subset $\mathcal{U}_{1}\subset\mathcal{Z}$ with $\mathcal{U}_{1}\cap\mathcal{S}_{2}\neq\varnothing$, we can use the definition of the localized gap function (2.4) specialized as the optimality gap (2.7) to obtain
$\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{T_{K}\sigma_{K}},$ (4.23)
where $C_{\mathcal{U}_{1}}\triangleq\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\bar{\omega}(z)$. This is the upper bound in (3.7). The lower bound in (3.7) holds since $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$:
[…]
Establishing the improved rate on the optimality gap (3.8)
If additionally the lower level solution set enjoys $(\alpha,\rho)$ weak-sharpness we have from (2.9) and (3.6)
$\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2}$ $)\geq-B_{\mathcal{U}_{1}}\operatorname{dist}(\bar{z}^{K},\mathcal{S}_{2})$
$\geq-B_{\mathcal{U}_{1}}\left[\frac{\rho}{\alpha}\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\cap\mathcal{S}_{2})\right]^{1/\rho}\geq-B_{\mathcal{U}_{1}}\left[\frac{\rho}{\alpha}\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\right]^{1/\rho}$
$\geq-B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2T_{K}(\alpha/\rho)}+\frac{C_{\mathcal{U}_{2}}\sum_{k=1}^{K}t_{k}\sigma_{k}}{(\alpha/\rho)T_{K}}\right]^{1/\rho}.$
This verifies (3.8).
Proof of (3.11)
Choosing the penalty sequence $(\sigma_{k})_{k\geq 1}$ according to (3.9) and the step size $t_{k}=t$ small enough, the lower and upper complexity bound on the gap function becomes
$B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt(\alpha/\rho)}+\frac{aC_{\mathcal{U}_{2}}}{(\alpha/\rho)(1-\delta)(K+b)^{\delta}}\right]^{1/\rho}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{at(K+b)^{1-\delta}}.$
Note that if $\delta=1$ in the choice of the penalty sequence (3.9), we do not obtain convergence in terms of the gap function to 0, but rather only an $O(1)$ upper bound and $o(1)$ lower bound.
5 Extensions
This section contains several extensions. First, we develop estimates for the case in which the upper level $\operatorname{HVI}$ is strongly monotone, implying that the entire problem (P) has a unique solution. We then move on by explaining how Algorithm 1 can be extended to the Bregman setting, and Tseng’s modified extragradient method.
5.1 Improved rates under stronger monotonicity assumptions
On top the already formulated hypothesis, we impose the following:
Assumption 4.
The operator ${\mathsf{F}}_{1}:\mathcal{Z}\to\mathcal{Z}$ is $\mu$-strongly monotone.
Since we assume that the over hierarchical equilibrium problem has a nonempty feasible set, Assumption 4 implies that $\mathcal{S}_{1}=\{z^{*}\}$ for some $z^{*}\in\mathcal{S}_{2}$. We again establish rates on gap functions in terms of the suitably adapted ergodic average
$\bar{z}^{K}=\frac{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}z^{i+1/2}}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}},$ (5.1)
where $\gamma_{k}\triangleq\frac{1}{\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)}$ and $\gamma_{0}\triangleq 1$.
Theorem 5.1.
Consider problem (P). Let Assumptions 1, 2, 4 be in place. Let additionally either Assumption 3 hold or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be compact. Let $(\sigma_{k})_{k\geq 0}$ be non-increasing and the sequences $(t_{k})_{k\geq 0}$ and $(\sigma_{k})_{k\geq 0}$ satisfy
$4t^{2}_{k}L^{2}_{k}+2t_{k}\sigma_{k}\mu\leq 1,\qquad k\geq 1.$ (5.2)
Let $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be a nonempty compact set. Then, there exists a constant $\tilde{C}_{\mathcal{U}_{2}}>0$ for which
$\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{\sigma_{1}\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}+\tilde{C}_{\mathcal{U}_{2}}\frac{\sum_{i=1}^{k}t_{i}\sigma^{2}_{i}\gamma_{i}}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$ (5.3)
Additionally, for any nonempty compact subset $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ with $\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$, there exists a constant $B_{\mathcal{U}_{1}}>0$ such that
$B_{\mathcal{U}_{1}}\operatorname{dist}(\bar{z}^{K},\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{2\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$ (5.4)
In particular, if the lower level solution set $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp, then
$B_{\mathcal{U}_{1}}\left[\frac{\sigma_{1}\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{2(\alpha/\rho)\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}+\frac{\tilde{C}_{\mathcal{U}_{1}\cap\mathcal{S}_{2}}\sum_{i=1}^{k}t_{i}\sigma^{2}_{i}\gamma_{i}}{(\alpha/\rho)\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}\right]^{1/\rho}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{2\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$ (5.5)
Proof.
We start with a refined version of the recursion (4.12). Recall the notation (4.9) and the bound (4.1) that holds for any $z\in\mathcal{Z}$ and $k\geq 1$:
$E_{k+1}(z)$ $\leq E_{k}(z)-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$+t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$ (5.6)
For the second and third terms in the r.h.s. we have from (4.6)
$\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq 2t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
$+\frac{1}{2}\left(4t_{k}^{2}L^{2}_{k}-1\right)\norm{z^{k+1/2}-z^{k}}^{2}.$ (5.7)
Using the definition of $z^{k}$ and $z^{k-1/2}$, the non-expansiveness of the proximal mapping (LABEL:eq:nonexpansive), and the step size assumption (5.2), we further obtain
$2t^{2}_{k}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}$ $\leq 2t^{2}_{k}t_{k-1}^{2}L_{k}^{2}\norm{{\mathsf{V}}_{k-1}(z^{k-1/2})-{\mathsf{V}}_{k-1}(z^{k-3/2})}^{2}\leq(1-t_{k}\sigma_{k}\mu)\frac{t_{k-1}^{2}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-1/2})-{\mathsf{V}}_{k-1}(z^{k-3/2})}^{2}.$ (5.8)
Combining (5.7), (5.8), and using the notation (4.11), we obtain
$\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq(1-t_{k}\sigma_{k}\mu)D_{k}-D_{k+1}+\frac{1}{2}\left(4t_{k}^{2}L^{2}_{k}-1\right)\norm{z^{k+1/2}-z^{k}}^{2}.$ (5.9)
By the strong monotonicity, we obtain
$\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle$ $=\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z\rangle+\sigma_{k}\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z\rangle$
$\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+\sigma_{k}\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+\sigma_{k}\mu\norm{z^{k+1/2}-z}^{2}$
$\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+\sigma_{k}\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+\frac{\sigma_{k}\mu}{2}\norm{z^{k}-z}^{2}-\sigma_{k}\mu\norm{z^{k+1/2}-z^{k}}^{2},$ (5.10)
where the last inequality follows from the triangle inequality and $(a+b)^{2}/2\leq a^{2}+b^{2}\Leftrightarrow a^{2}\geq(a+b)^{2}/2-b^{2}$. Plugging (5.9) and (5.10) multiplied by 1-1 into (5.6), recalling the notation (4.9) and $G_{k}(\cdot)=g_{2}(\cdot)+\sigma_{k}g_{1}(\cdot)$, we obtain
$E_{k+1}(z)$ $+D_{k+1}\leq(1-t_{k}\sigma_{k}\mu)(E_{k}(z)+D_{k})+(2t^{2}_{k}L^{2}_{k}+t_{k}\mu\sigma_{k}-\frac{1}{2})\norm{z^{k+1/2}-z^{k}}^{2}$
$-t_{k}\left(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)+\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)\right)$
$\leq(1-t_{k}\sigma_{k}\mu)(E_{k}(z)+D_{k})$
$-t_{k}\left(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)+\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)\right)$ (5.11)
where the last inequality uses (5.2).
We now use this recursion to establish rates on the feasibility and optimality gap, starting with the feasibility gap.
Since the analysis for the monotone setting holds also for strongly monotone case, we have that either Lemma 4.3 still holds under the same assumptions made there (i.e. either Assumption 3 or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ compact). Thus, in any case, for a compact subset $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$, there exists a constant $\tilde{C}_{\mathcal{U}_{2}}$ such that
$\langle{\mathsf{F}}_{1}(z),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2})\leq\tilde{C}_{\mathcal{U}_{2}},\qquad\forall z\in\mathcal{U}_{2},k\geq 1.$
Combining this with (5.11), we get
$t_{k}(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z))$ $\leq\left(1-t_{k}\sigma_{k}\mu\right)(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1})+t_{k}\sigma_{k}\widetilde{C}_{\mathcal{U}_{2}}.$
Multiplying both sides by $\sigma_{k}\gamma_{k}$ (see (5.1) for the definition of $\gamma_{k}$) and using that $\sigma_{k}\geq\sigma_{k+1}$, we obtain
$t_{k}\sigma_{k}\gamma_{k}(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z))$ $\leq\sigma_{k}\gamma_{k-1}(E_{k}(z)+D_{k})-\sigma_{k+1}\gamma_{k}(E_{k+1}(z)+D_{k+1})+t_{k}\sigma_{k}^{2}\gamma_{k}\widetilde{C}_{\mathcal{U}_{2}}.$
Thus, telescoping these inequalities and using the convexity of $g_{2}$, for the ergodic average (5.1), we obtain
$\langle{\mathsf{F}}_{2}(z),\bar{z}^{K}-z\rangle+g_{2}(\bar{z}^{K})-g_{2}(z)$ $\leq\frac{\sigma_{1}(E_{1}(z)+D_{1})+\widetilde{C}_{\mathcal{U}_{2}}\sum_{i=1}^{K}t_{i}\sigma_{i}^{2}\gamma_{i}}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$
Taking the supremum over $\mathcal{U}_{2}$ on both sides and using the definition of the localized gap function (2.4) specialized as the feasibility gap (2.6) together with $D_{1}=0$, we obtain the upper bound (5.3).
To obtain a bound in terms of the optimality gap, consider an arbitrary reference point $z\in\mathcal{S}_{2}$ and rearrange (5.11) to arrive at
$t_{k}\sigma_{k}(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z))$ $\leq\left(1-t_{k}\sigma_{k}\mu\right)(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1}).$
Multiplying the previous inequality by $\gamma_{k}$ yields
$t_{k}\sigma_{k}\gamma_{k}(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z))$ $\leq\gamma_{k-1}(E_{k}(z)+D_{k})-\gamma_{k}(E_{k+1}(z)+D_{k+1}).$
Telescoping these inequalities, using the convexity of $g_{1}$, $D_{1}=0$, $\gamma_{0}=1$, for the ergodic average (5.1), we obtain
$\langle{\mathsf{F}}_{1}(z),\bar{z}^{K}-z\rangle+g_{1}(\bar{z}^{K})-g_{1}(z)$ $\leq\frac{E_{1}(z)}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$
Since $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is a nonempty compact set with $\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$, we have that $\mathcal{U}_{1}\cap\mathcal{S}_{2}\neq\varnothing$. Taking supremum over $\mathcal{U}_{1}\cap\mathcal{S}_{2}$, we arrive at the upper bound in (5.4). The lower bound in (5.4) is the same as in (3.7).
If additionally the lower level solution set enjoys $(\alpha,\rho)$ weak-sharpness we obtain (5.5) by combining (5.4), (2.9), and (5.3). $\blacksquare$
To have a better understanding of the obtained rates, we now make a particular choice of the sequences $(t_{k})_{k\geq 1}$, $(\sigma_{k})_{k\geq 1}$.
Corollary 5.2.
Let the same Assumptions as in Theorem 5.1 be in place. Assume that the regularization sequences $(t_{k})_{k\geq 1}$ and $(\sigma_{k})_{k\geq 1}$ are chosen as
$t_{k}=\frac{1}{4(L_{k}+\sigma_{k}\mu)}=\frac{1}{4(L_{{\mathsf{F}}_{2}}+\sigma_{k}(L_{{\mathsf{F}}_{1}}+\mu))},\qquad\sigma_{k}=\frac{4L_{{\mathsf{F}}_{2}}}{\mu k},\qquad k\geq 1.$ (5.12)
Then, we have
$\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{8L_{{\mathsf{F}}_{2}}(L_{{\mathsf{F}}_{1}}+\mu)\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{\mu K}+\tilde{C}_{\mathcal{U}_{2}}\frac{4L_{{\mathsf{F}}_{2}}(1+\ln K)}{\mu K},$ (5.13)
$-B_{\mathcal{U}_{1}}\left[\frac{8L_{{\mathsf{F}}_{2}}(L_{{\mathsf{F}}_{1}}+\mu)\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{\mu K(\alpha/\rho)}+\tilde{C}_{\mathcal{U}_{2}}\frac{4L_{{\mathsf{F}}_{2}}(1+\ln K)}{\mu K(\alpha/\rho)}\right]^{1/\rho}\overset{(*)}{\leq}\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})$ (5.14)
$\leq\frac{2(L_{{\mathsf{F}}_{1}}+\mu)\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{K},$ (5.15)
where the inequality ()(*) holds under the $(\alpha,\rho)$-weakly sharpness assumption.
Proof.
First, we show that (5.2) is indeed satisfied:
$4t_{k}^{2}L_{k}^{2}+2t_{k}\sigma_{k}\mu=4\frac{L_{k}^{2}}{16(L_{k}+\sigma_{k}\mu)^{2}}+\frac{2\sigma_{k}\mu}{4(L_{k}+\sigma_{k}\mu)}\leq\frac{1}{4}+\frac{1}{2}\leq 1.$ (5.16)
Further,
$t_{i}\sigma_{i}=\frac{1}{4(L_{{\mathsf{F}}_{2}}+\sigma_{i}(L_{{\mathsf{F}}_{1}}+\mu))}\cdot\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}=\frac{1}{L_{{\mathsf{F}}_{2}}+\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}(L_{{\mathsf{F}}_{1}}+\mu)}\cdot\frac{L_{{\mathsf{F}}_{2}}}{\mu i}=\frac{1}{i+\frac{4}{\mu}(L_{{\mathsf{F}}_{1}}+\mu)}\cdot\frac{1}{\mu}=\frac{1}{i+\varkappa}\cdot\frac{1}{\mu},$ (5.17)
where we denoted $\varkappa=\frac{4}{\mu}(L_{{\mathsf{F}}_{1}}+\mu)$. This gives us
$\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)=\prod_{i=1}^{k}(1-\frac{1}{i+\varkappa}\cdot\frac{1}{\mu}\cdot\mu)=\prod_{i=1}^{k}(1-\frac{1}{i+\varkappa})=\prod_{i=1}^{k}(\frac{i+\varkappa-1}{i+\varkappa})=\frac{\varkappa}{k+\varkappa}$ (5.18)
and $\gamma_{k}=1/\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)=\frac{k+\varkappa}{\varkappa}$. Finally, we have
$\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}=\sum_{i=1}^{K}\frac{1}{i+\varkappa}\cdot\frac{1}{\mu}\cdot\frac{i+\varkappa}{\varkappa}=\frac{K}{\mu\varkappa}=\frac{K}{4(L_{{\mathsf{F}}_{1}}+\mu)},$ (5.19)
$\sum_{i=1}^{K}t_{i}\sigma_{i}^{2}\gamma_{i}=\sum_{i=1}^{K}\frac{1}{i+\varkappa}\cdot\frac{1}{\mu}\cdot\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}\cdot\frac{i+\varkappa}{\varkappa}=\frac{L_{{\mathsf{F}}_{2}}}{\mu(L_{{\mathsf{F}}_{1}}+\mu)}\sum_{i=1}^{K}\frac{1}{i}\leq\frac{L_{{\mathsf{F}}_{2}}(1+\ln K)}{\mu(L_{{\mathsf{F}}_{1}}+\mu)}.$ (5.20)
Substituting this into (5.3), we obtain (5.13). Substituting this into (5.4) and (5.5), we obtain (5.14). $\blacksquare$
5.2 Tseng splitting
Optimistic extragradient (also known as Popov’s or past extragradient) method saves one evaluation of the time-varying operator $V_{k}$ per iteration compared to the standard extragradient method. Yet, it still requires two evaluations of the proximal operator of $G_{k}$ per iteration. An attractive modified extragradient scheme, which uses past values of the operator but saves on one evaluation of the proximal operator, is the following optimistic version of Tseng’s Forward-Backward-Forward method [Tsen00, tran2024revisiting]
$z^{k+1/2}$ $=\operatorname{\textsf{prox}}_{t_{k}G_{k}}(z^{k}-t_{k}{\mathsf{V}}_{k}(z^{k-1/2})),$
$z^{k+1}$ $=z^{k+1/2}-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2})).$
By construction, there exits a $\xi^{k}\in\partial G_{k}(z^{k+1/2})$ such that
$z^{k+1/2}$ $=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k-1/2})+\xi^{k}),$
$z^{k+1}$ $=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k-1/2})+\xi^{k})-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}))$
$=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k}).$
This further gives us
$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k+1}-z^{k}+z^{k}-z}^{2}$
$=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+2\langle z^{k+1}-z^{k},z^{k+1}-z\rangle$
$=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+\langle z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k})-z^{k},z^{k+1}-z\rangle$
$=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1}-z^{k+1/2}\rangle$
$\hskip 17.00024pt-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1/2}-z\rangle.$ (5.21)
Now observe that $t_{k}({\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k})=z^{k}-z^{k+1}=z^{k}-z^{k+1/2}+t_{k}({\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2})).$ As a consequence, we get
$2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})$ $+\xi^{k},z^{k+1}-z^{k+1/2}\rangle$
$=2\langle z^{k}-z^{k+1/2},z^{k+1}-z^{k+1/2}\rangle+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle$
$=\norm{z^{k}-z^{k+1/2}}^{2}+\norm{z^{k+1}-z^{k+1/2}}^{2}-\norm{z^{k+1}-z^{k}}$
$\hskip 17.00024pt2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle.$
Combining this estimate with (5.21) gives
$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}-\norm{z^{k}-z^{k+1/2}}^{2}+\norm{z^{k+1}-z^{k+1/2}}^{2}+\norm{z^{k+1}-z^{k}}^{2}$
$\hskip 17.00024pt2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle$
$\hskip 17.00024pt-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi_{k},z^{k+1/2}-z\rangle$
Since $\xi^{k}\in\partial G_{k}(z^{k+1/2})$ it holds that $G_{k}(z^{k+1/2})-G_{k}(z)\leq\langle\xi^{k},z^{k+1/2}-z\rangle$ and therefore
$-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1/2}-z\rangle$ $=-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle-2t_{k}\langle\xi^{k},z^{k+1/2}-z\rangle$
$\leq 2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+2t_{k}(G_{k}(z)-G_{k}(z^{k+1/2}))$ (5.22)
Moreover, Young’s inequality yields
$2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle\leq t_{k}^{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}+\norm{z^{k+1}-z^{k+1/2}}.$
With this estimate, we arrive at (4.1)
$\frac{1}{2}\norm{z^{k+1}-z}^{2}$ $\leq\frac{1}{2}\norm{z^{k}-z}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+\frac{t_{k}^{2}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$ (5.23)
$\hskip 17.00024pt+t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+t_{k}(G(z)-G(z^{k+1/2})).$
Repeating the same steps as in the rest of Section 4.1, we obtain the same energy bound (4.12). Using this energy bound and repeating the same arguments as in Section 4, we obtain the same results as in Theorem 3.1 and Corollary 3.2.
6 Numerical Experiments
We verify the performance of the optimistic extragradient method with two numerical examples.
6.1 Hierarchical Nash equilibria
To provide a simple illustration on the practical performance of our method, we’ve tested Algorithm 1 on a simple version of the hierarchical equilibrium problem described Section 1.1.4, taken from [LamSagSIOPT25]. We’ve chosen a deliberately simple example for which the equilibrium set of the game can be computed explicitly and we can monitor the evolution of the inexactness of the algorithm. Let us consider $N=4$ lower-level players and $M=2$ upper-level players, with $x^{1}=(y^{2},y^{4}),x^{2}=(y^{1},y^{3})$. The player-cost functions are
$h^{\ell}_{1}(y^{1},y^{-1})=0.5(y^{1})^{2}+y^{1}(y^{2}+2y^{3}+y^{4}-100),\;\varphi^{\ell}_{1}(y^{1})=\delta_{[-100,50]}(y^{1}),$
$h^{\ell}_{2}(y^{2},y^{-2})=0.5(y^{2})^{2}+y^{2}(y^{1}+y^{3}+y^{4}-50),\;\varphi^{\ell}_{2}(y^{2})=\max\{-10(y^{2}-15),0\}+\delta_{[0,50]}(y^{2}),$
$h^{\ell}_{3}(y^{3},y^{-3})=0.5(y^{3})^{2}+y^{3}(y^{2}+y^{4}-100),\;\varphi^{\ell}_{3}(y^{3})=\delta_{[0,100]}(y^{3})$
$h^{\ell}_{4}(y^{4},y^{-4})=0.5(y^{4})^{2}+y^{4}(y^{1}+y^{2}+y^{2}+y^{3}-50),\;\varphi^{\ell}_{4}(y^{4})=\delta_{[0,50]}(y^{4}).$
For the upper-level players, we assume
$h^{u}_{1}(x_{1},x_{2})=(y^{2}-20)^{2}+(y^{4}-50)^{2}+(y^{2}+y^{4})(y^{1}+y^{3}),\varphi^{u}_{1}(x^{1})=0,$
$h^{u}_{2}(x_{1},x_{2})=(y^{1})^{2}+y^{1}(y^{2}+y^{3})+(y^{3})^{2}+y^{3}(y^{2}+y^{4}),\varphi^{u}_{2}(x^{1})=0.$
One can obtain an explicit expression for the lower-level equilibrium set: $\mathcal{S}_{2}=\{(-50,y^{2},50,50-y^{2}):15\leq y^{2}\leq 50\}$. Thus, the unique variational equilibrium of this game is (50,15,50,35)(-50,15,50,35). We’ve solved this game problem with Algorithm 1 using the regularization sequence $\sigma_{k}=\frac{1}{(k+3)^{1/2}}$ and constant step size. Figure 1 displays the temporal evolution of the strategies of the four players, as well as the distance to the unique variational equilibrium
Figure 1: Evolution of the strategy profile and error plot.
6.2 Constrained Min-Max problem
We consider min-max problems with joint linear constraints,
$\min_{x}\max_{y}\mathcal{L}(x,y)\hskip 17.00024pt\text{s.t. }(x,y)\in\operatorname*{argmin}_{(x^{\prime},y^{\prime})\in\mathcal{X}\times\mathcal{Y}}\frac{1}{2}\norm{{\mathbf{A}}x^{\prime}+{\mathbf{B}}y^{\prime}-c}^{2}_{\mathcal{Z}}$ (6.1)
where ${\mathbf{A}}:\mathcal{X}\to\mathcal{Z}?$ and ${\mathbf{B}}:\mathcal{Y}\to\mathcal{Z}?$ are bounded linear operators. The saddle function $\mathcal{L}:\mathcal{Z}\triangleq\mathcal{X}\times\mathcal{Y}\to(-\infty,\infty]$ is of the form
$\mathcal{L}(x,y)=\varphi(x)+f_{1}(x)+\langle x,\mathbf{K}y\rangle-f_{2}(y)-\psi(y),$
[…]
${\mathsf{F}}_{1}(x,y)=\begin{pmatrix}\nabla f_{1}(x)\\ \nabla f_{2}(y)\end{pmatrix}+\begin{pmatrix}\mathbf{K}y\\ -\mathbf{K}^{\ast}x\end{pmatrix},\;{\mathsf{F}}_{2}(x,y)=\begin{pmatrix}{\mathbf{A}}^{\ast}({\mathbf{A}}x+{\mathbf{B}}y-c)\\ {\mathbf{B}}^{\ast}({\mathbf{A}}x+{\mathbf{B}}y-c)\end{pmatrix}$,
$\partial g_{1}(x,y)=\partial\varphi(x)\times\partial\psi(y),\;g_{2}(x,y)=0.$
Using the product space structure $z\triangleq(x,y)\in\mathcal{X}\times\mathcal{Y}\triangleq\mathcal{Z}$, the optimistic extragradient method can be applied to solve this family of equilibrium problems.
Generalized Absolute value equations (GAVEs).
The Generalized Absolute value equations (GAVEs) [mangasarian2006absolute] is an important non-smooth NP-hard problem in the form of
[…]
We approach this problem via the relaxed formulation
[…]
For our numerical investigations, we’ve implemented the following example from [Ebadi24]: Consider the boundary value problem
[…]
$u(0)=-1,u(1)=0$
where $f(x)=x^{2}-1$. This second-order ODE has the exact solution
$u(x)=0.1961\sin(x)-4\cos(x)-x^{2}+3\qquad 0\leq x\leq 1.$
Let $h=\frac{1}{n+1}$ be the mesh size generating the grid $x_{0}=0,x_{i}=ih,1\leq i\leq n,x_{n+1}=1$, we obtain the discretized equation
[…]
Figure 2: Evolution of the approximate solution $u(k)$ after completion of the $k$-th iteration of the Algorithm (left), and the evolution of the residual relative to the true solution (right).
Using the five-point central difference formula for $u^{\prime\prime}(x_{i})$, we obtain a problem of the form (6.2) with
$A=\frac{1}{12h^{2}}\begin{pmatrix}20&-6&-4&1&&&&&\\ -16&30&-16&1&&&&&\\  1&-16&30&-16&1&&&&\\  &\ddots&\ddots&\ddots&&&&&\\ &&&&1&-16&30&-16&1\\  &&&&&1&-16&30&-16\\  &&&&&1&-4&-6&20\\  \end{pmatrix},\;b=\begin{pmatrix}f(x_{1})+\frac{11u_{0}}{12h^{2}}\\ f(x_{2})-\frac{u_{0}}{12h^{2}}\\ f(x_{3})\\ \vdots\\ f(x_{n-2})\\ f(x_{n-1})-\frac{u_{n-1}}{12h^{2}}\\ f(x_{n})+\frac{11u_{n+1}}{12h^{2}}\end{pmatrix}$.
We’ve reconstructed the solution $u(x)$ using Algorithm 1 with regularization sequence $\sigma_{k}=\frac{1}{(k+3)^{1/2}}$ and constant step size. Figure LABEL:fig:GAVEODESolution displays the evolution of approximate solutions obtained with our method measured at the indicated snapshots of the algorithm.
7 Conclusion
This paper proves seminal results on the iteration complexity of hierarchical hemi-variational inequalities. Constructing a suitably defined optimistic extragradient method, we generalize the state-of-the art along several dimensions. In particular, we derive complexity guarantees in terms of suitably defined feasibility and optimality gaps. Our analysis reveals close connections to geometric conditions imposed on the operator defining the lower level equilibrium problems, conditions which have been so-far dominantly been used in the potential case. We belief that our proof strategy is going to be useful for several other problems currently under investigation in the literature. In particular, it will be very interesting to derive rates for stochastic formulation of our model template and derive rates in terms of expected gap functions. Further interesting directions for future research inlude the consideration of accelerated methods. We leave these important extensions to future research.
Acknowledgements
This research benefited from the support of the FMJH Program PGMO. MST’s research is supported by the Deutsche Forschungsgemeinschaft (DFG) - Projektnummer 556222748 "non-stationary hierarchical minimization". We thank Johannes-Carl Schnebel for his help on conducting the numerical experiments, and Enis Chenchene, Radu I. Bot, as well as David A. Hulett for fruitful discussions on this topic.
Appendix
Appendix A The Fitzpatrick function
Let $\mathsf{M}:\mathcal{Z}\to 2^{\mathcal{Z}}$ be a maximally monotone operator. The Fitzpatrick function [fitzpatrick1988representing, BauCom16] $\mathcal{F}_{\mathsf{M}}:\mathcal{Z}\to(-\infty,+\infty]$, associated with the operator $\mathsf{M}$, is defined as
$\mathcal{F}_{\mathsf{M}}(x,u)\triangleq\sup_{(y,v)\in\operatorname{graph}(\mathsf{M})}\{\langle x,v\rangle+\langle y,u\rangle-\langle y,v\rangle\}.$
Following [Borwein:2016aa], we define the gap function $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)\triangleq\mathcal{F}_{\mathsf{M}}(x,0)$. $\operatorname{\mathsf{Gap}}_{\mathsf{M}}$ is convex, and in fact the smallest translation invariant gap function associated with the monotone operator $\mathsf{M}$ [Borwein:2016aa, Theorem 3.1]. Importantly, this gives the properties $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)\geq 0$ and $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)=0$ if and only if $x\in\operatorname{zer}(\mathsf{M})$. To make this concept concrete, observe that if $\mathsf{M}={\mathsf{F}}+\operatorname{\mathsf{NC}}_{\mathcal{C}}$, then the above definition of the gap function reduces to the well-known Auslender dual gap function [FacPan03, auslender1976optimisation]
$\operatorname{\mathsf{Gap}}_{{\mathsf{F}}+\operatorname{\mathsf{NC}}_{\mathcal{C}}}(x)=\sup_{y\in\mathcal{C}}\langle{\mathsf{F}}(y),x-y\rangle.$
If $\mathsf{M}={\mathsf{F}}+\partial g$ for a function $g\in\Gamma_{0}(\mathcal{Z})$, we easily obtain
$\operatorname{\mathsf{Gap}}_{{\mathsf{F}}+\partial g}(x)\leq\sup_{y\in\operatorname{dom}(g)}\langle{\mathsf{F}}(y),x-y\rangle+g(x)-g(y).$
To the data $({\mathsf{F}},g)$, we thus associate the bifunction $H^{({\mathsf{F}},g)}:\mathcal{Z}\times\mathcal{Z}\to[-\infty,\infty]$ defined as
$H^{({\mathsf{F}},g)}(x,y)\triangleq\langle{\mathsf{F}}(y),x-y\rangle+g(x)-g(y).$
If ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ is monotone and continuous on $\operatorname{dom}(g)$, it is easy to verify that $H^{({\mathsf{F}},g)}(x,y)\leq-H^{({\mathsf{F}},g)}(y,x)$ for all $(x,y)\in\operatorname{dom}(g)\times\operatorname{dom}(g)$ (i.e. $H^{({\mathsf{F}},g)}$ is a monotone bifunction [iusem2011maximal]). In the structured setting $\operatorname{HVI}({\mathsf{F}},g)$, we obtain the following bounds on the Fitzpatrick function, showing its close connection to the bifunction $H^{({\mathsf{F}},g)}$ and the Auslender dual gap function. First, the convex subgradient inequality yields the relation
$\mathcal{F}_{{\mathsf{F}}+\partial g}(x,u)$ $=\sup_{y\in\operatorname{dom}(\partial g),\xi\in\partial g(y)}\{\langle x-y,{\mathsf{F}}(y)\rangle+\langle\xi,x-y\rangle+\langle y,u\rangle\}$
$\leq\sup_{y\in\operatorname{dom}(g)}\{\langle{\mathsf{F}}(y),x-y\rangle+g(x)-g(y)+\langle y,u\rangle\}$
$=\sup_{y\in\operatorname{dom}(g)}\{H^{({\mathsf{F}},g)}(x,y)+\langle y,u\rangle\}\triangleq\varphi^{({\mathsf{F}},g)}(x,u).$
The function $\varphi^{({\mathsf{F}},g)}(x,u)$ is thus seen as the Fitzpatrick transform of $H^{({\mathsf{F}},g)}$. Second, since the function $x\mapsto H^{({\mathsf{F}},g)}(x,y)$ is convex, we see
$\varphi^{({\mathsf{F}},g)}(x,u)\leq\sup_{y\in\operatorname{dom}(g)}\{\langle y,u\rangle-H^{({\mathsf{F}},g)}(y,x)\}=\left(H^{({\mathsf{F}},g)}(\bullet,x)\right)^{\ast}(u).$ (A.1)
In particular,
$\varphi^{({\mathsf{F}},g)}(x,u)\geq\langle x,u\rangle\hskip 17.00024pt\forall x\in\operatorname{dom}(g),$ (A.2)
$\mathcal{F}_{{\mathsf{F}}+\partial g}(x,0)\leq\sup_{y\in\operatorname{dom}(g)}H^{({\mathsf{F}},g)}(x,y)=\varphi^{({\mathsf{F}},g)}(x,0).$ (A.3)
Following this notation, we define for a given a subset $\mathcal{C}\subseteq\mathcal{Z}$, the restricted dual gap function as
$\Theta(x|{\mathsf{F}},g,\mathcal{C})\triangleq\sup_{y\in\mathcal{C}}H^{({\mathsf{F}},g)}(x,y).$ (A.4)
Appendix B Omitted proofs
Proof of Lemma 2.2.
The lower bound $\Theta(x|{\mathsf{F}},g,\mathcal{C})\geq 0$ is clear for $x\in\mathcal{C}$. Let’s assume that $x\in\mathcal{C}$ is a solution to (2.3). Then, by the monotonicity of $\mathsf{F}$} we have
$\langle{\mathsf{F}}(y),y-x\rangle+g(y)-g(x)\geq\langle{\mathsf{F}}(x),y-x\rangle+g(y)-g(x)\geq 0\hskip 17.00024pt\forall y\in\mathcal{Z}.$
Hence,
$\Theta(x|{\mathsf{F}},g,\mathcal{C})=\sup_{y\in\mathcal{C}}\{\langle{\mathsf{F}}(y),x-y\rangle+g(x)-g(y)\}\leq 0.$
This, together with already obtained inequality $\Theta(x|{\mathsf{F}},g,\mathcal{C})\geq 0$ implies $\Theta(x|{\mathsf{F}},g,\mathcal{C})=0$.
Now assume $\Theta(x|{\mathsf{F}},g,\mathcal{C})=0$ for $x\in\mathcal{C}$. Then,
$\langle{\mathsf{F}}(x^{\prime}),x-x^{\prime}\rangle+g(x)-g(x^{\prime})\leq 0\hskip 17.00024pt\forall x^{\prime}\in\mathcal{C}.$
Equivalently,
$\langle{\mathsf{F}}(x^{\prime}),x^{\prime}-x\rangle+g(x^{\prime})-g(x)\geq 0\hskip 17.00024pt\forall x^{\prime}\in\mathcal{C}.$
This means $x\in\mathcal{C}$ is a Minty (weak) solution to the hemi-variational inequality with the data $({\mathsf{F}},g)$ over the set $\mathcal{C}$. Let $w\in\mathcal{C}$ be arbitrary and consider v=tw+(1t)xv=tw+(1-t)x for $t\in[0,1]$. Since $\mathcal{C}$ is convex, we have $v\in\mathcal{C}$. It follows
$0$ $\leq\langle{\mathsf{F}}(v),v-x\rangle+g(v)-g(x)$
$=\langle{\mathsf{F}}(x+t(w-x)),t(w-x)\rangle+g(tw+(1-t)x)-g(x)$
$\leq t\langle{\mathsf{F}}(x+t(w-x)),w-x\rangle+t(g(w)-g(x))$
Dividing both sides by $t$ and then letting $t\to 0^{+}$, the weak continuity of $\mathsf{F}$} implies
$0\leq\langle{\mathsf{F}}(x),w-x\rangle+g(w)-g(x).$ (B.1)
Since $w\in\mathcal{C}$ has been chosen arbitrarily, we see that $x$ is a Stampacchia (strong) solution of the hemivariational inequality with the additional constraint $x\in\mathcal{C}$.
We now show that if $\mathcal{C}\subset\operatorname{dom}(g)$, and there exists $\epsilon>0$ such that $\mathcal{U}\triangleq B(x,\epsilon)\cap\mathcal{C}=B(x,\epsilon)\cap\operatorname{dom}(g)$, then (B.1) holds also for $w\in\operatorname{dom}(g)$. To that end, assume to the contrary, there exits $z\in\operatorname{dom}(g)$ for which
$\langle{\mathsf{F}}(x),z-x\rangle+g(z)-g(x)<0.$
Then, by convexity of $\operatorname{dom}(g)$ and definition of $\mathcal{U}$ there exists $\lambda>0$ small enough for which $w=x+\lambda(z-x)\in\mathcal{U}$ and by convexity
$0\leq\langle{\mathsf{F}}(x),w-x\rangle+g(w)-g(x)\leq\lambda(\langle F(x),z-x\rangle+g(z)-g(x))<0,$
thus arriving at a contradiction. Hence, $\langle F(x),z-x\rangle+g(z)-g(x)\geq 0$ for any $z\in\operatorname{dom}(g)$ and moreover for any $z\in\mathcal{Z}$, which finishes the proof that $x$ is a solution to (2.3). $\blacksquare$
Proof of Proposition 2.4.
Pick $z\in\operatorname{dom}(g)$. Without loss of generality, assume $z\notin\mathcal{S}$. Let $\bar{z}=\Pi_{\mathcal{S}}(z)$, so that there exists $p\in\operatorname{\mathsf{NC}}_{\mathcal{S}}(\bar{z})$ satisfying $p=z-\bar{z}$. Hence, $p^{\ast}\triangleq\frac{z-\bar{z}}{\norm{z-\bar{z}}}$ is a unit norm element of $\operatorname{\mathsf{NC}}_{\mathcal{S}}(\bar{z})$ satisfying
$\langle p^{\ast},z-\bar{z}\rangle=\norm{z-\bar{z}}.$
[…]
$\tau p^{\ast}-{\mathsf{F}}(\bar{z})-\xi^{\ast}\in[\operatorname{\mathsf{TC}}_{\operatorname{dom}(g)}(\bar{z})\cap\operatorname{\mathsf{NC}}_{\mathcal{S}}(\bar{z})]^{\circ}.$
Hence,
$\tau\langle p^{\ast},z-\bar{z}\rangle\leq\langle{\mathsf{F}}(\bar{z})+\xi^{\ast},z-\bar{z}\rangle\leq\langle{\mathsf{F}}(\bar{z}),z-\bar{z}\rangle+g(z)-g(\bar{z})$
$\Rightarrow\tau\norm{z-\bar{z}}\leq H^{({\mathsf{F}},g)}(z,\bar{z})$
$\Rightarrow\tau\operatorname{dist}(z,\mathcal{S})\leq H^{({\mathsf{F}},g)}(z,\bar{z})\leq\sup_{y\in\operatorname{dom}(g)}H^{({\mathsf{F}},g)}(z,y)=\Theta(z|{\mathsf{F}},g,\operatorname{dom}(g)).$
$\blacksquare$
Proof of Lemma 2.8.
Let $z^{*}\in\mathcal{S}_{1}=\operatorname{zer}({\mathsf{F}}_{1}+\partial g_{1}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})\subset\mathcal{S}_{2}$. Then, there exists $p^{*}\in\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}(z^{*})$ such that ${\mathsf{F}}_{1}(z^{*})-p^{*}\in\partial g_{1}(z^{*})$. By the convex sugradient inequality, this implies
$\langle{\mathsf{F}}_{1}(z^{*}),z-z^{*}\rangle+g_{1}(z)-g_{1}(z^{*})\geq\langle-p^{*},z-z^{*}\rangle\hskip 17.00024pt\forall z\in\mathcal{Z}.$ (B.2)
Take $z\in\mathcal{Z}$ and let $\hat{z}=\Pi_{\mathcal{S}_{2}}(z)$ be the orthogonal projector of $z$ onto $\mathcal{S}_{2}$. Thus, $\langle p^{*},\hat{z}-z^{*}\rangle\leq 0$, resulting in
$\langle p^{*},\hat{z}-z\rangle\leq\langle p^{*},z^{*}-z\rangle,$
which implies when combined with (B.2)
$\langle{\mathsf{F}}_{1}(z^{*}),z-z^{*}\rangle+$ $g_{1}(z)-g_{1}(z^{*})\geq\langle p^{*},\hat{z}-z\rangle$
$\geq-\norm{p^{*}}\cdot\norm{\hat{z}-z}=-\norm{p^{*}}\operatorname{dist}(z,\mathcal{S}_{2}).$
Hence, for all compact $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})$ with $z^{*}\in\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$, we conclude $\Theta_{\rm Opt}(z|\mathcal{U}_{1}\cap\mathcal{S}_{1})\geq-B_{\mathcal{U}_{1}}\operatorname{dist}(z,\mathcal{S}_{2})$, where $B_{\mathcal{U}_{1}}=\sup_{\mathcal{U}_{1}\cap\mathcal{S}_{1}}\norm{p^{*}}$.
To show (2.9), we can directly use Definition 2.5, to conclude
$\Theta_{\rm Feas}(z|\mathcal{U}_{2}\cap\mathcal{S}_{2})=\sup_{z^{*}\in\mathcal{U}_{2}\cap\mathcal{S}_{2}}H^{({\mathsf{F}}_{2},g_{2})}(z,z^{*})\geq\frac{\alpha}{\rho}\operatorname{dist}(z,\mathcal{S}_{2})^{\rho}.$
$\blacksquare$

### v2

@@
 1 Introduction
-Solving hierarchical equilibrium problems is an increasingly active area in mathematical programming that consists in finding a solution to an outer variational inequality over the solutions of an inner variational inequality.
-In this paper we consider the following problem.
-Let $(\mathcal{Z},\langle\cdot,\cdot\rangle)$ be a real Hilbert space, potentially infinite-dimensional.
-We consider the problem of solving the hierarchical equilibrium problem, in which a hemi-variational inequality ($\operatorname{HVI}$) is solved over the solution set of another mixed variational inequality.
-Specifically, the class of equilibrium problems we consider is of the form
+Let $(\mathcal{Z},\langle\cdot,\cdot\rangle)$ be a real separable Hilbert space.
+We consider the problem of solving the hierarchical equilibrium problem, in which a hemi-variational inequality ($\operatorname{HVI}$), also known as variational inequality of the second kind, is solved over the solution set of another $\operatorname{HVI}$:
  $\text{ Find }z^{\ast}\in\mathcal{S}_{2}\text{ s.t.
-}\langle{\mathsf{F}}_{1}(z^{\ast}),z-z^{\ast}\rangle+g_{1}(z)-g_{1}(z^{\ast})\geq 0\qquad\forall z\in\mathcal{S}_{2}\triangleq\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2}).$ (P)
+}\langle{\mathsf{F}}_{1}(z^{\ast}),z-z^{\ast}\rangle+g_{1}(z)-g_{1}(z^{\ast})\geq 0\qquad\forall z\in\mathcal{S}_{2}\triangleq\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2}),$ (P)
 where ${\mathsf{F}}_{1},{\mathsf{F}}_{2}:\mathcal{Z}\to\mathcal{Z}$ are monotone operators and $g_{1},g_{2}$ are proper, convex, and lower-semicontinuous functions.
-We refer to the set $\mathcal{S}_{2}\triangleq\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2})$ (assumed to be non-empty) as the solution set of the lower-level problem.
+We refer to the set $\mathcal{S}_{2}$ (assumed to be non-empty) as the solution set of the lower-level problem.
 In the same vein, we will call $\mathcal{S}_{1}$ the set of solutions of (P).
-Obviously, these sets are nested in the sense $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$.
-Problem (P) is the formulation of a system of nested $\operatorname{HVI}$’s (also known as variational inequalities of the second kind).
-These families of variational problems are prominent in mechanics [goeleven2013variational] and optimal control [Barbu93, de2011optimal].
-In finite dimensions, it has become a useful model template to study a plethora of problems in mathematical programming, machine learning, operations research, game theory, and signal processing [malitsky2020golden, InertialFBF, Facchinei:2014aa].
+Obviously, these sets are nested in the sense that $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$.
+Hierarchical $\operatorname{HVI}$’s as in (P) are prominent in mechanics [28] and optimal control [8, 19].
+In finite dimensions, it has become a useful model template to study a plethora of problems in mathematical programming, machine learning, control, game theory, and signal processing [42, 16, 24].
 Some motivating examples are given below.
@@
 […]
-This model problem has been studied in several very recent papers.
-Important references include [merchav2023convex], who approach this problem via a disciplined proximal splitting ansatz, as well as [AminiYous19, DoronShtern23, yousefian2021bilevel, merchav2024fast, boct2025accelerating].
+This model has been studied in several very recent papers, e.g., in [48], which approach this problem via a disciplined proximal splitting ansatz, as well as in [3, 21, 65, 47, 13].
 1.1.2 Hierarchical variational inequalities
-If $g_{1}=0$ and $g_{2}=\operatorname{\iota}_{\mathcal{K}}$ for a non-empty, closed convex set $\mathcal{K}\subseteq\mathcal{Z}$, problem (P) reduces to
+If $g_{1}=0$ and $g_{2}$ is the convex indicator of a non-empty, closed convex set $\mathcal{K}\subseteq\mathcal{Z}$, problem (P) reduces to
 $\text{Find }z^{\ast}\in\mathcal{S}_{2}\qquad\text{s.t.: }\langle{\mathsf{F}}_{1}(z^{\ast}),z-z^{\ast}\rangle\geq 0\hskip 17.00024pt\forall z\in\mathcal{S}_{2}=\operatorname{zer}({\mathsf{F}}_{2}+\operatorname{\mathsf{NC}}_{\mathcal{K}}).$
-This is the problem investigated in the recent papers [samadi_improved_2025, marschner2025tikhonov] and [alves2025inertial].
-1.1.3 Equilibrium selection in Nash games
-Selecting solutions from a given variational inequality problem is another important application of the model template (P).
-If ${\mathsf{F}}_{1}(x)=\nabla f_{1}(x)$, then we recover the optimal equilibrium selection problem as
-$\min_{x\in\mathcal{S}}\{f_{1}(x)+g_{1}(x)\},\qquad\mathcal{S}=\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2}).$
-[…]
+This is a challenging formulation of nested variational inequalities, which has been recently investigated in [57, 46, 2].
 1.1.3 Hierarchical jointly convex generalized Nash equilibrium problem
-In very recent work, [LamSagSIOPT25] studied a projection-based method for solving a challenging class of jointly convex generalized Nash equilibrium problems (GNEPs) featuring hierarchy and non-smooth data.
-We first show that their model is a special case of the problem (P).
-The lower level Nash equilibrium problem.
+A very recent work [38] studied a projection method for solving a challenging class of jointly convex generalized Nash equilibrium problems (GNEPs) featuring hierarchy and non-smooth data (see [17] for a stochastic formulation of a similar problem).
+The lower-level Nash Equilibrium Problem (NEP)
 Let $\nu\in\{1,\ldots,N\}$ represent the labels of the lower-level players, who engage in a Nash game in which each player’s parameterized optimization problem is given by
 $\min_{y^{\nu}\in\mathcal{Y}^{\nu}}\{h^{\ell}_{\nu}(y^{\nu},y^{-\nu}))+\varphi^{\ell}_{\nu}(y^{\nu})\}.$
-The data of this game problem enjoy the following assumptions:
+The data of this game problem enjoys the following assumptions:
   1. (a)
@@
   3. (c)
-    The game’s pseudogradient ${\mathsf{F}}_{2}(y)=(\nabla_{y^{1}}h^{\ell}_{1}(y),\ldots,\nabla_{y^{N}}h^{\ell}_{N}(y))$ is monotone on $\mathcal{Y}\triangleq\prod_{\nu=1}^{N}\mathcal{Y}^{\nu}$;
+    The game’s pseudogradient ${\mathsf{F}}_{2}(y)\triangleq(\nabla_{y^{1}}h^{\ell}_{1}(y),\ldots,\nabla_{y^{N}}h^{\ell}_{N}(y))$ is monotone on the separable Hilbert space $\mathcal{Y}\triangleq\prod_{\nu=1}^{N}\mathcal{Y}^{\nu}$;
   4. (d)
-    $\varphi_{\nu}^{\ell}(\cdot)$ is proper closed convex with compact domain $\operatorname{dom}(\varphi^{\ell}_{\nu})\triangleq\mathcal{A}_{\nu}\subset\mathcal{Y}_{\nu}$.
-We can characterize equilibria of the Nash game $\Gamma_{\rm low}=\{h^{\ell}_{\nu},\varphi_{\nu}^{\ell}\}_{1\leq\nu\leq N}$ as solutions to the mixed variational inequality $\mathcal{S}_{2}=\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2})$, where $g_{2}(y^{1},\ldots,y^{N})\triangleq\sum_{\nu=1}^{N}\varphi^{\ell}_{\nu}(y^{\nu}).$ The assumptions on the game problem ensure that $\operatorname{dom}(g_{2})$ is a compact subset of the separable Hilbert space $\mathcal{Y}$.
+    The functions $\varphi_{\nu}^{\ell}(\cdot)$ are proper, closed, and convex, with compact domain.
+We can characterize equilibria of the Nash game $\Gamma_{\rm low}\triangleq\{h^{\ell}_{\nu},\varphi_{\nu}^{\ell}\}_{1\leq\nu\leq N}$ as solutions to the $\operatorname{HVI}$ with data ${\mathsf{F}}_{2}$, $g_{2}(y^{1},\ldots,y^{N})\triangleq\sum_{\nu=1}^{N}\varphi^{\ell}_{\nu}(y^{\nu})$ (see e.g.
+[25]).
+Hence, $\mathcal{S}_{2}=\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2})$.
+The above assumptions ensure that $\operatorname{dom}(g_{2})$ is a compact subset of $\mathcal{Y}$.
 Furthermore, $\mathcal{S}_{2}$ is nonempty, convex, and compact.
-The upper level GNEP.
-The upper level problem is the Nash game with joint coupling constraint represented by the equilibrium set $\mathcal{S}_{2}$.
+The upper level GNEP
+The upper-level problem is the Nash game with joint coupling constraint represented by the equilibrium set $\mathcal{S}_{2}$.
 Let $\mu\in\{1,\ldots,M\}$ denote the labels of upper-level players.
@@
 $\min_{x^{\mu}\in\mathcal{X}^{\mu}}\{h^{u}_{\mu}(x^{\mu},x^{-\mu})+\varphi^{u}_{\mu}(x^{\mu})\}\qquad\text{s.t.: }(x^{\mu},x^{-\mu})\in\mathcal{S}_{2}.$
-Similar to the upper level NEP, we impose the following assumptions on the problem data:
-  1. (d)
+As for the lower-level NEP, we make the following assumptions on the problem data:
+  1. (e)
 […]
-  2. (e)
-    The operator ${\mathsf{F}}_{1}(x)=(\nabla_{x^{1}}h^{u}_{1}(x),\ldots,\nabla_{x^{N}}h^{u}_{N}(x))$ is monotone on $\mathcal{X}\triangleq\prod_{\nu=1}^{N}\mathcal{X}^{\nu}$;
   2. (f)
+    The operator ${\mathsf{F}}_{1}(x)\triangleq(\nabla_{x^{1}}h^{u}_{1}(x),\ldots,\nabla_{x^{N}}h^{u}_{N}(x))$ is monotone on the product space $\mathcal{X}\triangleq\prod_{\nu=1}^{N}\mathcal{X}^{\nu}$;
+  3. (g)
     $\varphi^{u}_{\mu}$ is proper closed convex with closed convex domain $\operatorname{dom}(\varphi^{u}_{\mu})$.
 The decision variables $x^{\mu}$ correspond to blocks $(y^{\nu})_{\nu\in\mathcal{N}_{\mu}}$, in which $\mathcal{N}_{\mu}\subseteq\{1,\ldots,N\}$ represent the lower-level players controlled by upper-level player $\mu$.
-The blocks can well be overlapping, meaning that individual lower level players can have multiple upper level players.
 In particular, $\mathcal{X}^{\mu}=\prod_{\nu\in\mathcal{N}_{\mu}}\mathcal{Y}^{\nu}$.
-The entire hierarchical equilibrium problem thus belongs to our template (P), with $\mathcal{S}_{1}=\operatorname{zer}({\mathsf{F}}_{1}+\partial g_{1}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})$, where $g_{1}(x^{1},\ldots,x^{M})\triangleq\sum_{\mu=1}^{M}\varphi^{u}_{\mu}(x^{\mu}).$
+The entire hierarchical equilibrium problem thus belongs to our template (P), with $\mathcal{S}_{1}=\operatorname{zer}({\mathsf{F}}_{1}+\partial g_{1}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})$, where $g_{1}(x^{1},\ldots,x^{M})\triangleq\sum_{\mu=1}^{M}\varphi^{u}_{\mu}(x^{\mu}).$ Note that if $M=1$ (i.e., there is only one player on the upper level), then we recover the problem of equilibrium selection in Nash games, which is an important problem in game theory [65, 11].
 1.2 Related works
-This paper belongs to a subclass of bilevel problems in which a single decision variable is to be designed in order to optimize a function over the set of minimizerrs of another function.
-To emphasize this special character of such problems are also known as simple bilevel problems [Dempe:2021aa].
-Starting with the seminal contribution of [Solodov:2007aa] and [Cabot2005], Tikhonov regularization has become one of the dominant paradigms in the numerical solution of simple bilevel optimization problems.
-In the context of the present paper, this means we have to consider the iteration-dependent operators
-${\mathsf{V}}_{k}\triangleq{\mathsf{F}}_{2}+\sigma_{k}{\mathsf{F}}_{1},\qquad G_{k}\triangleq g_{2}+\sigma_{k}g_{1},$
-where $k$ is the iteration counter and $(\sigma_{k})_{k}$ is a well-chosen sequence.
-Over the past years, a quite significant list of publications have appeared proving rates of convergence with respect to the inner (lower-level) and the outer (upper-level) problem.
-Specifically, building on the generalized Tikhonov regularization framework of [Solodov:2007aa], the first results in this direction have been produced in [beck2014first, SabSht17].
-[merchav2023convex] develop a proximal subgradient method for convex composite models in the upper and lower level problem.
-Adaptive versions of a related proximal-gradient strategy are studied in [latafat2025convergence].
-Accelerated simple bilevel optimization in the convex non-smooth composite setting have recently been derived in [merchav2024fast].
-Beyond proximal-gradient based methods, projection-free schemes are developed in [DoronShtern23] and [giang2024projection].
-Moving beyond optimization, there exists a significant literature, rooted in the numerical resolution of inverse problems and signal processing, which studied solution methods for selecting a solution to a variational inequality given a pre-defined selection criterion [ogura2002non, xu2003convergence, marino2011explicit, ono2014hierarchical].
-This is known as the equilibrium selection problem and is in fact a special case of the general family of Mathematical Programs under equilibrium constraints (MPEC) [luo1996mathematical, Hintermuller:2014aa].
+This paper focuses on a subclass of bilevel problems, usually termed as simple bilevel problems [20], in which a single decision variable is to be designed in order to solve the upper-level problem over the solution set of a lower-level problem.
+Starting with the seminal contributions [59, 1], Tikhonov regularization has become one of the dominant paradigms in the numerical solution of simple bilevel optimization problems.
+In our $\operatorname{HVI}$ context, this idea leads to the definition of the data that combines both levels in (P) as
+${\mathsf{V}}_{k}\triangleq{\mathsf{F}}_{2}+\sigma_{k}{\mathsf{F}}_{1},\qquad G_{k}\triangleq g_{2}+\sigma_{k}g_{1},$ (1.1)
+where $k$ is the iteration counter and $(\sigma_{k})_{k\geq 1}$ is a well-chosen regularization sequence.
+Over the past years, quite a significant list of publications appeared, proving rates of convergence with respect to the inner (lower-level) and the outer (upper-level) problem in the setting of hierarchical optimization as in Section 1.1.1.
+The first results in this direction are [10, 56] for strongly convex upper level functions.
+[48] develop an alternating proximal subgradient method for convex composite models in the upper and lower level problem.
+Building on the iterative regularization structure, adaptive versions of a related proximal-gradient strategy are studied in [39] without rates on the upper level problem.
+[58] and [13] present rates for both upper and lower level problems for the iterative proximal gradient method and its accelerated version, which was first analyzed in [47].
+Beyond proximal-gradient based methods, projection-free schemes are developed in [21, 27].
+Moving beyond optimization, there exists a significant body of literature, rooted in the numerical resolution of inverse problems and signal processing, that studied solution methods for selecting a solution to a variational inequality given a pre-defined selection criterion [51, 64, 45, 52].
+This is known as the equilibrium selection problem, a special case of the general family of mathematical programs under equilibrium constraints [41, 29].
 All these papers focus on asymptotic convergence of the sequence generated by tailor-made numerical algorithms.
-In this more enhanced setting, the only results on iteration complexity we are aware of are published in the papers [kaushik2023incremental, Kaushik:2021aa, jalilzadeh2024stochastic].
-Moving beyond the function-case in the upper-level (equilibrium selection), there exist a few papers on numerical methods for solving nested and hierarchical variational inequalities [van2021regularization, ThongNumerical20, lampariello2022solution].
-Among those, only [lampariello2022solution] contains complexity statements.
-Recently, the papers [samadi_improved_2025] and [alves2025inertial] have addressed the finite-dimensional variational inequality case, obtained from our model template (P) by imposing $g_{1}=0$ and $g_{2}$ an indicator function over a compact convex set.
-Both papers derive complexity guarantees for the Korpelevich version of the extragradient method, applied to the Tikhonov regularized variational inequality with the operator ${\mathsf{F}}_{2}+\sigma_{k}{\mathsf{F}}_{1}$ at iteration $k$.
-The obtained rates are $O(1/k^{\delta})$ and $O(1/k^{1-\delta})$, where $\delta\in(0,1)$ is a parameter defining the strength of the Tikhonov regularization.
-In a very recent work, [LamSagSIOPT25] studied a difficult class of game problems featuring hierarchy and non-smooth local cost functions.
-We show in Section 6 that their model falls into our general problem framework.
-Their algorithm is not a full splitting method, and thus not directly comparable to our operator splitting approach.
+The only results on iteration complexity we are aware of are published in the papers [34, 35, 33].
+Moving beyond the function-case in the upper-level (equilibrium selection), there exist a few papers on numerical methods for solving hierarchical VIs [63, 60, 37].
+Among those, only [37] contains complexity statements.
+Recently, the papers [57] and [2] have addressed the finite-dimensional variational inequality case, obtained from our model template (P) by imposing $g_{1}=0$ and $g_{2}$ to be an indicator function over a compact convex set.
+Both papers derive complexity guarantees for the Korpelevich extragradient method, applied to the Tikhonov regularized variational inequality with the operator ${\mathsf{F}}_{2}+\sigma_{k}{\mathsf{F}}_{1}$ at iteration $k$.
+The obtained rates are $O(1/k^{\delta})$ (lower level) and $O(1/k^{1-\delta})$ (upper level), where $\delta\in(0,1)$ is a parameter defining the strength of the Tikhonov regularization.
 1.3 Contributions
@@
 Our main contributions are as follows:
-  • More general problem formulation: Unlike [samadi_improved_2025, alves2025inertial], our hierarchical equilibrium model (P) contains general convex composite terms $g_{1},g_{2}$.
-Specifically, these papers treat the special case $g_{1}=0$ and $g_{2}=\operatorname{\iota}_{\mathcal{C}}$ for a compact convex and nonempty set $\mathcal{C}$ in a finite-dimensional euclidean vector space.
-Our analysis performed in potentially infinite-dimensional real Hilbert spaces, and considers general hemi-variational inequalities ($\operatorname{HVI}$’s).
-This is class contains a rather large class of variational models, which found significant application in optimal control and mechanics [Facchinei:2014aa, goeleven2013variational].
-Besides the gain in modelling, adding these convex and potentially non-smooth functions is an important extension, since it allows us to cover the non-smooth composite convex model when specialized to the potential case, as studied in [merchav2024fast].
-  • Enhanced Algorithmic design: The recent papers [samadi_improved_2025, alves2025inertial] approach the numerical resolution of problem (P) via double-call version of the extragradient method.
-Our approach instead uses the ideas of the optimistic extragradient method [hsieh2019convergence] and thus requires only one operator evaluation per iteration.
-Within this algorithmic setting, we derive sublinear convergence rates for the lower and upper level problem and improved rates under the strong monotonicity of ${\mathsf{F}}_{1}$.
-Specifically, using a polynomial regularization sequence $\sigma_{k}=O(k^{-\delta})$ for $\delta\in(0,1)$, we demonstrate upper complexity bounds on the order of $O(k^{-\delta})$ for the gap function associated with the lower level equilibrium problem, and a $O(k^{-(1-\delta)})$ complexity bound for the gap function of the entire hierarchical problem.
-This matches the existing bounds reported in [samadi_improved_2025, alves2025inertial], but for a much larger class of equilibrium problems, and fewer function evaluations (specifically, instead of 2 just 1 per iteration).
-  • Unbounded domains: An important assumption in the complexity analysis in [samadi_improved_2025, alves2025inertial] is compactness of the domains over which the variational inequalities are solved.
+More general problem formulation: Unlike [57, 2], our hierarchical equilibrium model (P) contains general convex composite terms $g_{1},g_{2}$.
+Specifically, these papers treat the special case $g_{1}=0$ and $g_{2}=\operatorname{\iota}_{\mathcal{C}}$, the indicator function of a compact convex and nonempty set $\mathcal{C}$ in a finite-dimensional Euclidean vector space.
+Our analysis is performed in potentially infinite-dimensional real Hilbert spaces, and considers general $\operatorname{HVI}$’s.
+This class contains a rather large set of variational models, which have found significant application in optimal control and mechanics [24, 8, 28].
+Besides the gain in modeling, adding these convex and potentially non-smooth functions is an important extension, as it allows treating the non-smooth composite convex model when specialized to the potential case, as studied in [47].
+Enhanced Algorithmic design and combined rate statements: The recent papers [57, 2] approach the numerical resolution of problem (P) via double-call version of the extragradient method.
+Our approach instead builds on the optimistic extragradient method [30], requiring only one operator evaluation per iteration.
+Within this algorithmic setting, we derive sublinear convergence rates for the lower- and upper-level problem and improved rates under the strong monotonicity of ${\mathsf{F}}_{1}$.
+Specifically, using a polynomial regularization sequence $\sigma_{k}=O(k^{-\delta})$ for $\delta\in(0,1)$, we demonstrate upper complexity bounds on the order of $O(k^{-\delta})$ for the gap function associated with the lower-level equilibrium problem (feasibility gap), and a $O(k^{-(1-\delta)})$ complexity bound for the gap function of the entire hierarchical problem (optimality gap).
+These bounds match the existing bounds reported in [57, 2], while applying to a much broader class of equilibrium problems and requiring fewer function evaluations.
+Unbounded domains: An important assumption in the complexity analysis in [57, 2] is compactness of the domains over which the variational inequalities are solved.
 We abandon this assumption.
-Instead, our proof builds on ideas originating in the convergence analysis of non-autonomous evolution equations due to [attouch2018asymptotic].
-In the potential case, these ideas have already been successfully applied to the study of hierarchical optimization problems [Bot:2014aa, Cabot2005].
-In particular, [boct2025accelerating] make the role of this technique very transparent by emphasizing the role of the geometry of the lower level problem in establishing a unified complexity statement for both levels simultaneously.
+Instead, our proof builds on ideas originating from the asymptotic analysis of non-autonomous evolution equations due to [4].
+In the potential case, these ideas have already been successfully applied to the study of hierarchical optimization problems [14, 1].
+In particular, [13] make the role of this technique very transparent by emphasizing the importance of the geometry of the lower level problem in establishing a unified complexity statement at both levels simultaneously.
 A methodological contribution of this work is to show how these ideas naturally translate to the variational setting.
-We belief that our proof technique is going to be useful also for other hierarchical equilibrium problems as well.
+We believe that our proof technique is going to be useful for other hierarchical equilibrium problems as well.
+The outline of this paper is as follows.
+Section 2 collects relevant facts on VIs, defining the geometric properties we assume for the solution set of the lower-level problem, and introducing the merit functions we employ in our complexity statements.
+Section 3 contains the main algorithmic paradigm we employ to prove our rates.
+Section 4 details the proof of our two main theorems.
+Section 5 reports important extensions.
+We establish improved rates under strong monotonicity of the upper-level $\operatorname{HVI}$ and show that the main statements extend to optimistic versions of the forward-backward-forward splitting, enjoying only one proximal operator evaluation per iteration.
+Section 6 presents numerical experiments demonstrating our results.
 2 Facts on variational inequalities
@@
 […]
-$\operatorname{\textsf{prox}}_{G}(u)=\operatorname*{argmin}_{z\in\mathcal{Z}}\{G(z)+\frac{1}{2}\norm{z-u}^{2}\}.$
-If $x,u\in\mathcal{Z}$, we have
 $x=\operatorname{\textsf{prox}}_{G}(u)\Leftrightarrow\langle u-x,z-x\rangle\leq G(z)-G(x)\hskip 17.00024pt\forall z\in\mathcal{Z}.$ (2.1)
-The metric projection onto a nonempty closed convex set $\mathcal{C}\subset\mathcal{Z}$ is defined as $\Pi_{\mathcal{C}}(x)=\operatorname*{argmin}_{y\in\mathcal{C}}\norm{x-y}=\operatorname{\textsf{prox}}_{\operatorname{\iota}_{\mathcal{C}}}(x)$, where the convex indicator function $\operatorname{\iota}_{\mathcal{C}}:\mathcal{Z}\to(-\infty,\infty]$ is defined by
-$\operatorname{\iota}_{\mathcal{C}}(x)\triangleq\left\{\begin{array}[]{cc}0&\text{if }x\in\mathcal{C},\\ +\infty&\text{ if }x\notin\mathcal{C}.\end{array}\right$.
-We denote also $\operatorname{dist}(x,\mathcal{C})\triangleq\norm{x-\Pi_{\mathcal{C}}(x)}$.
-The polar cone attached to a closed convex set $\mathcal{C}\subset\mathcal{Z}$ is $\mathcal{C}^{\circ}\triangleq\{z\in\mathcal{Z}|\langle z,x\rangle\leq 0\qquad\forall x\in\mathcal{C}\}$.
-The conjugate function of $f\in\Gamma_{0}(\mathcal{Z})$ is $f^{\ast}(y)\triangleq\sup_{x\in\operatorname{dom}(f)}\{\langle y,x\rangle-f(x)\}.$ A mapping (operator) ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ is $\mu$-strongly-monotone ($\mu\geq 0$) if
+For a closed convex set $\mathcal{C}\subset\mathcal{Z}$, the metric projection is $\Pi_{\mathcal{C}}(x)\triangleq\operatorname{\textsf{prox}}_{\operatorname{\iota}_{\mathcal{C}}}(x)$, where the convex indicator function $\operatorname{\iota}_{\mathcal{C}}:\mathcal{Z}\to(-\infty,\infty]$ is defined by $\operatorname{\iota}_{\mathcal{C}}(x)=0$ if $x\in\mathcal{C}$, and $\operatorname{\iota}_{\mathcal{C}}(x)=+\infty$ else.
+We also denote $\operatorname{dist}(x,\mathcal{C})\triangleq\norm{x-\Pi_{\mathcal{C}}(x)}$.
+The polar cone attached to a set $\mathcal{C}\subset\mathcal{Z}$ is $\mathcal{C}^{\circ}\triangleq\{z\in\mathcal{Z}|\langle z,x\rangle\leq 0\qquad\forall x\in\mathcal{C}\}$.
+The normal cone at point $z\in\mathcal{C}$ is defined as $\operatorname{\mathsf{NC}}_{\mathcal{C}}(z)\triangleq\{p\in\mathcal{Z}|\langle p,z^{\prime}-z\rangle\leq 0\qquad\forall z^{\prime}\in\mathcal{C}\}$.
+For any convex $\mathcal{C}\subset\mathcal{Z}$ (not neccesarily closed), the tangent cone $\operatorname{\mathsf{TC}}_{\mathcal{C}}(x)$ of $\mathcal{C}$ at $x\in\mathcal{C}$ is polar to the normal cone, and we have the explicit expression $\operatorname{\mathsf{TC}}_{\mathcal{C}}(x)=\operatorname{cl}\left(\bigcup_{\lambda>0}\frac{\mathcal{C}-x}{\lambda}\right)$.
+The Fenchel conjugate of $f\in\Gamma_{0}(\mathcal{Z})$ is $f^{\ast}(y)\triangleq\sup_{x\in\operatorname{dom}(f)}\{\langle y,x\rangle-f(x)\}.$ Specifically, given a closed convex set $\mathcal{C}\subset\mathcal{Z}$, the Fenchel conjugate of $\operatorname{\iota}_{\mathcal{C}}$ is the support function of set $\mathcal{C}$ defined as $\operatorname{\mathtt{s}}(p|\mathcal{C})\triangleq\sup_{z\in\mathcal{C}}\langle p,z\rangle.$ Note that
+$\langle p,z\rangle=\operatorname{\mathtt{s}}(p|\mathcal{C})\hskip 17.00024pt\forall p\in\operatorname{\mathsf{NC}}_{\mathcal{C}}(z).$ (2.2)
+A mapping (operator) ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ is $\mu$-strongly-monotone ($\mu\geq 0$) if
 $\langle{\mathsf{F}}(z_{1})-{\mathsf{F}}(z_{2}),z_{1}-z_{2}\rangle\geq\mu\norm{z_{1}-z_{2}}^{2}\hskip 17.00024pt\forall z_{1},z_{2}\in\mathcal{Z}.$
-A 0-monotone operator is simply called monotone. ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ is $L$-Lipschitz continuous if
+A 0-monotone operator is simply called monotone. $\mathsf{F}$} is $L$-Lipschitz continuous if
 $\norm{{\mathsf{F}}(z_{1})-{\mathsf{F}}(z_{2})}\leq L\norm{z_{1}-z_{2}}\hskip 17.00024pt\forall z_{1},z_{2}\in\mathcal{Z}.$
-The normal cone of a closed convex set $\mathcal{C}\subset\mathcal{Z}$ at point $z\in\mathcal{C}$ is defined $\operatorname{\mathsf{NC}}_{\mathcal{C}}(z)\triangleq\{\xi\in\mathcal{Z}|\langle\xi,z^{\prime}-z\rangle\leq 0\qquad\forall z^{\prime}\in\mathcal{C}\}$.
-Under convexity, the tangent cone $\operatorname{\mathsf{TC}}_{\mathcal{C}}(x)$ of $\mathcal{C}$ at $x\in\mathcal{C}$ is polar to the normal cone, and we have the explicit expression $\operatorname{\mathsf{TC}}_{\mathcal{C}}(x)=\operatorname{cl}\left(\bigcup_{\lambda>0}\frac{\mathcal{C}-x}{\lambda}\right)$.
-The support function of set $\mathcal{C}$ is defined as $\operatorname{\mathtt{s}}(p|\mathcal{C})\triangleq\sup_{z\in\mathcal{C}}\langle p,z\rangle.$ It follows from these definitions that
-$\langle p,z\rangle=\operatorname{\mathtt{s}}(p|\mathcal{C})\hskip 17.00024pt\forall p\in\operatorname{\mathsf{NC}}_{\mathcal{C}}(z).$ (2.2)
-Lemma 2.1 (Lemma 5.31, [BauCom16]).
-[…]
-$a_{k+1}\leq a_{k}-b_{k}+c_{k}.$
-Then, $\lim_{k\to\infty}a_{k}$ exists and $\sum_{k\geq 1}b_{k}<\infty$.
-2.2 Hemi-variational inequalities
-Given a monotone and Lipschitz continuous mapping ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ and a function $g\in\Gamma_{0}(\mathcal{Z})$, the hemi-variational inequality problem $\operatorname{HVI}({\mathsf{F}},g)$ is to find $z^{\ast}\in\mathcal{Z}$ such that
+2.2 Hemi-variational inequalities (HVI’s)
+Given a monotone and Lipschitz continuous mapping ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ and a function $g\in\Gamma_{0}(\mathcal{Z})$, the $\operatorname{HVI}$ problem $\operatorname{HVI}({\mathsf{F}},g)$ is to find $z^{\ast}\in\mathcal{Z}$ such that
 $\langle{\mathsf{F}}(z^{\ast}),u-z^{\ast}\rangle+g(u)-g(z^{\ast})\geq 0\hskip 17.00024pt\forall u\in\mathcal{Z}.$ (2.3)
-One can rewrite (2.3) as a monotone inclusion $0\in({\mathsf{F}}+\partial g)(z^{\ast})$.
-If $g=\operatorname{\iota}_{\mathcal{C}}$ for a closed convex set $\mathcal{C}\subset\mathcal{Z}$, we obtain from the above the variational inequality $\operatorname{VI}({\mathsf{F}},\mathcal{C})$:
+One can rewrite (2.3) as a monotone inclusion $0\in({\mathsf{F}}+\partial g)(z^{\ast})$, or equivalently the solution set is $\mathcal{S}=\operatorname{zer}({\mathsf{F}}+\partial g)$.
+If $g=\operatorname{\iota}_{\mathcal{C}}$ for a closed convex set $\mathcal{C}\subset\mathcal{Z}$, we recover the variational inequality problem $\operatorname{VI}({\mathsf{F}},\mathcal{C})$:
 $\text{Find }z^{\ast}\in\mathcal{C}\text{ s.t. }\langle{\mathsf{F}}(z^{\ast}),u-z^{\ast}\rangle\geq 0\hskip 17.00024pt\forall u\in\mathcal{C}.$
-Let $\mathcal{C}\subset\operatorname{dom}(g)$ be a given compact subset of $\operatorname{dom}(g)$.
-A popular convergence measures for $\operatorname{HVI}({\mathsf{F}},g)$ is the localized gap function
+Let $\mathcal{C}\subset\operatorname{dom}(g)$ be a given compact subset. A popular merit function for $\operatorname{HVI}({\mathsf{F}},g)$ is the localized (dual) gap function
 $\Theta(z|{\mathsf{F}},g,\mathcal{C})\triangleq\sup_{y\in\mathcal{C}}\{\langle{\mathsf{F}}(y),z-y\rangle+g(z)-g(y)\}=\sup_{y\in\mathcal{C}}H^{({\mathsf{F}},g)}(z,y),$ (2.4)
-where $H^{({\mathsf{F}},g)}:\mathcal{Z}\times\mathcal{Z}\to[-\infty,\infty]$ is defined as $H^{({\mathsf{F}},g)}(z,y)\triangleq\langle{\mathsf{F}}(y),z-y\rangle+g(z)-g(y)$.
-Historically, this localized version of a gap function can be traced back to [Minty62] (see also [kinderlehrer2000introduction, NesDual07]).
-Lemma 2.2 below summarizes its role for the numerical resolution of the problem $\operatorname{HVI}({\mathsf{F}},g)$ under continuity assumptions on $\mathsf{F}$}.
+where $H^{({\mathsf{F}},g)}:\mathcal{Z}\times\mathcal{Z}\to[-\infty,\infty]$ is defined as
+$H^{({\mathsf{F}},g)}(z,y)\triangleq\langle{\mathsf{F}}(y),z-y\rangle+g(z)-g(y).$ (2.5)
+Historically, this localized version of a gap function can be traced back to [49, 36].
+Lemma 2.1 below, whose proof is given in Appendix C, extends a result proved in [50], and summarizes its role for the numerical resolution of the problem $\operatorname{HVI}({\mathsf{F}},g)$ under continuity assumptions on $\mathsf{F}$}.
 Lemma 2.1.
 […]
-Proof.
-See Appendix B. $\blacksquare$
 2.3 Sharpness and error-bound property
-Our geometric framework is phrased in terms of weak-sharpness of the lower-level solution set $\mathcal{S}_{2}$.
-Originally formulated for optimization problems in [burke1993weak], weak-sharpness in the context of variational inequalities has been defined in [patriksson1993unified].
-Subsequently, implications in terms of error bounds of primal and dual gap functions have been stated in [marcotte1998weak, liu2016characterization].
-The following definition of weak sharpness is from [Huang:2018aa].
+Our geometric framework is phrased in terms of weak sharpness of the lower-level solution set $\mathcal{S}_{2}$.
+Originally formulated for optimization problems in [15], weak sharpness in the context of variational inequalities has been defined in [54].
+Subsequently, implications in terms of error bounds of primal and dual gap functions have been developed in [44, 40].
+The following definition of weak sharpness is from [31].
 Definition 2.2.
@@
 […]
-Based on [Huang:2018aa], we can give the following characterization of weak sharpness in terms of an error bound involving the dual gap function of $\operatorname{HVI}({\mathsf{F}},g)$.
+Based on [31], we give the following characterization of weak sharpness in terms of an error bound for the dual gap function of $\operatorname{HVI}({\mathsf{F}},g)$; see proof in Appendix C.
 Proposition 2.3.
-Consider problem $\operatorname{HVI}({\mathsf{F}},g)$ with $\operatorname{dom}(g)$ a closed convex nonempty subset of $\mathcal{Z}$.
-If $\mathcal{S}=\operatorname{zer}({\mathsf{F}}+\partial g)$ is weakly sharp, then
-$(\forall z\in\mathcal{D}_{g}):\qquad\Theta(z|{\mathsf{F}},g,\operatorname{dom}(g))\geq\tau\operatorname{dist}(z,\mathcal{S}).$
-Proof.
-See Appendix B. $\blacksquare$
-This Proposition shows that weak sharpness implies that the gap function satisfies an error bound property [pang1997error]. Motivated by this fact, we propose the following definition:
+Consider problem $\operatorname{HVI}({\mathsf{F}},g)$ with $\operatorname{dom}(g)$ a closed, convex, and nonempty subset of $\mathcal{Z}$.
+If the solution set $\mathcal{S}$ of $\operatorname{HVI}({\mathsf{F}},g)$ is weakly sharp, then
+$(\forall z\in\operatorname{dom}(g)):\qquad\Theta(z|{\mathsf{F}},g,\operatorname{dom}(g))\geq\tau\operatorname{dist}(z,\mathcal{S}).$
+Proposition 2.3 shows that weak sharpness implies that the gap function satisfies an error bound property [53], motivating the following definition:
 Definition 2.4 (Weak Sharpness).
-Let $\mathcal{S}$ be the nonempty solution set of $\operatorname{HVI}({\mathsf{F}},g)$. We say $\mathcal{S}$ is $(\alpha,\rho)$-weak sharp with $\alpha>0$ and $\rho\geq 1$ if
-$(\forall z^{\ast}\in\mathcal{S})(\forall z\in\mathcal{D}_{g}=\operatorname{dom}(g)):\qquad H^{({\mathsf{F}},g)}(z,z^{\ast})\geq\alpha\rho^{-1}\operatorname{dist}(z,\mathcal{S})^{\rho}.$ (2.5)
+Let $\mathcal{S}$ be the solution set of $\operatorname{HVI}({\mathsf{F}},g)$. We say $\mathcal{S}$ is $(\alpha,\rho)$-weakly sharp with $\alpha>0$ and $\rho\geq 1$ if
+$(\forall z^{\ast}\in\mathcal{S})(\forall z\in\operatorname{dom}(g)):\qquad H^{({\mathsf{F}},g)}(z,z^{\ast})\geq\frac{\alpha}{\rho}\operatorname{dist}(z,\mathcal{S})^{\rho}.$ (2.6)
 Remark 2.1.
-An important class of examples arises when $\rho=1$, notably in monotone linear complementarity problems in finite dimensions under nondegeneracy conditions [burke1993weak, pang1997error].
-In the case where the $\operatorname{HVI}({\mathsf{F}},g)$ reduces to a convex optimization problem, weak-sharpness implies an Hölderian errror bound, an assumption already imposed by [Cabot2005] in the context of hierarchical minimization.
+An important class of examples arises when $\rho=1$, notably in monotone linear complementarity problems in finite dimensions under nondegeneracy conditions [15, 53].
+In the case where the $\operatorname{HVI}({\mathsf{F}},g)$ reduces to a convex optimization problem, weak sharpness implies an Hölderian error bound, an assumption already imposed by [1] in the context of hierarchical minimization.
 Specifically, let us assume that ${\mathsf{F}}=0$ and $g\in\Gamma_{0}(\mathcal{Z})$.
-Then $\mathcal{S}=\operatorname*{argmin}g$, and acordingly, $H^{(0,g)}(z,z^{*})=g(z)-\min g$ for $z^{*}\in\mathcal{S}$.
+Then $\mathcal{S}=\operatorname*{argmin}g$, and accordingly, $H^{(0,g)}(z,z^{*})=g(z)-\min g$ for $z^{*}\in\mathcal{S}$.
 Hence, (2.6) implies
 $\frac{\alpha}{\rho}\operatorname{dist}(u,\operatorname*{argmin}g)^{\rho}\leq g(u)-\min g\hskip 17.00024pt\forall u\in\mathcal{Z}.$
-If $\rho=1$, this is the weak-sharpness condition of [burke1993weak]. The case $\rho=2$ corresponds to the "quadratic growth" condition of [drusvyatskiy2018error]. $\Diamond$
+If $\rho=1$, this is the weak-sharpness condition of [15]. The case $\rho=2$ corresponds to the "quadratic growth" condition of [22], a relaxation of strong convexity. $\Diamond$
 2.4 Constraint qualifications
@@
     $\mathcal{S}_{2}\triangleq\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2})\neq\varnothing$.
-We remark that Assumption 1 implies that $\mathcal{S}_{2}$ is a non-empty closed convex set. The next assumption is essentially a constraint qualification condition on the composite function
-$x\mapsto f_{\bar{x}}(x)\triangleq\langle{\mathsf{F}}_{1}(\bar{x}),x-\bar{x}\rangle+g_{1}(x)-g_{1}(\bar{x})+\operatorname{\iota}_{\mathcal{S}_{2}}(x).$
+We remark that Assumption 1 implies that $\mathcal{S}_{2}$ is a non-empty, closed, and convex set.
+The next assumption is essentially a constraint qualification condition on the composite function $x\mapsto\langle{\mathsf{F}}_{1}(\bar{x}),x-\bar{x}\rangle+g_{1}(x)-g_{1}(\bar{x})+\operatorname{\iota}_{\mathcal{S}_{2}}(x).$
 Assumption 2.
-The solution set $\mathcal{S}_{1}$ of (P) satisfies
-$\mathcal{S}_{1}=\operatorname{zer}({\mathsf{F}}_{1}+\partial g_{1}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}).$
-Similar assumptions have been made in [Bot:2014aa] for solving constrained variational inequalities. Assumption 2 is motivated by conditions ensuring a non-smooth sum rule of
+The solution set $\mathcal{S}_{1}$ of (P) satisfies $\mathcal{S}_{1}=\operatorname{zer}({\mathsf{F}}_{1}+\partial g_{1}+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})\neq\varnothing$.
+Similar assumptions have been made in [14] for solving constrained VI’s. Assumption 2 is motivated by conditions ensuring a non-smooth sum rule of the form
 $\partial_{x}\left(\langle{\mathsf{F}}_{1}(\bar{x}),\operatorname{Id}_{\mathcal{Z}}(\bullet)-\bar{x}\rangle+g_{1}(\bullet)-g_{1}(\bar{x})+\operatorname{\iota}_{\mathcal{S}_{2}}(\bullet)\right)(\bar{x})={\mathsf{F}}_{1}(\bar{x})+\partial g_{1}(\bar{x})+\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}(\bar{x}).$
-Common assumptions ensuring this property are [BauCom16, Corollary 16.48]:
-  • $\operatorname{int}\operatorname{dom}(g_{1})\cap\mathcal{S}_{2}\neq\varnothing$;
-  • $\operatorname{dom}(g_{1})\cap\operatorname{int}(\mathcal{S}_{2})\neq\varnothing$;
-  • if $\mathcal{Z}$ is finite-dimensional and $\operatorname{ri}\operatorname{dom}(g_{1})\cap\operatorname{ri}\operatorname{dom}(g_{2})\neq\varnothing$.
-Since $\mathcal{S}_{2}\subseteq\operatorname{dom}(\partial g_{2})\subset\operatorname{dom}(g_{2})$, these assumptions ensure the computationally pleasant identity $\partial(g_{1}+g_{2})=\partial g_{1}+\partial g_{2}$.
-2.5 Gap functions for hierarchical variational inequalities
-Since we are searching for a solution of a hierarchical system of variational inequalities, we have to introduce different merit functions for measuring the feasibility and the optimality of a test point $z\in\mathcal{Z}$.
+Common assumptions ensuring this property can be found in [9, Corollary 16.48].
+2.5 Gap functions for hierarchical HVI’s
+Since we are searching for a solution of a hierarchical system of VIs, we have to introduce different merit functions for measuring the feasibility and the optimality of a test point $z\in\mathcal{Z}$.
 Definition 2.5 (Feasibility gap).
@@
 It is thus important to obtain a lower bound on the optimality gap.
+This is achieved by Lemma 2.7 below, with proof provided in Appendix C.
+It extends the corresponding result in [13] to the operator case.
+[57] and [2] prove this result for the special case of bilevel VIs on compact domains.
 Lemma 2.7.
@@
 Suppose $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp.
-Then for all nonempty and compact subsets $\mathcal{U}_{2}\subseteq\operatorname{dom}(g_{2})$ and all $z\in\operatorname{dom}(g_{2})$, we have
+Then for all nonempty and compact subsets $\mathcal{U}_{2}\subseteq\operatorname{dom}(g_{2})$ s.t.
+$\mathcal{U}_{2}\cap\mathcal{S}_{2}\neq\varnothing$, and all $z\in\operatorname{dom}(g_{2})$, we have
 $\operatorname{dist}(z,\mathcal{S}_{2})\leq\left[\frac{\rho}{\alpha}\Theta_{\rm Feas}(z|\mathcal{U}_{2}\cap\mathcal{S}_{2})\right]^{1/\rho}.$ (2.10)
-Proof.
-See Appendix B. $\blacksquare$
-Remark 2.2.
-Lemma 2.8 allows us to establish a lower bound on the optimality gap over the solution set $\mathcal{S}_{2}$.
-It extends the corresponding result in [boct2025accelerating] to the operator case.
-[samadi_improved_2025] and [alves2025inertial] prove this result for special case of bilevel VIs on compact domains.
-$\Diamond$
 3 Optimistic extragradient method
-Our algorithmic design follows the popular Tikhonov regularization approach and employs a family of regularized problems with the data
-${\mathsf{V}}_{\sigma}(z)\triangleq{\mathsf{F}}_{2}(z)+\sigma{\mathsf{F}}_{1}(z):\mathcal{Z}\to\mathcal{Z},\text{ and }G_{\sigma}(z)\triangleq g_{2}(z)+\sigma g_{1}(z)\in\Gamma_{0}(\mathcal{Z}).$
-The parameter $\sigma\geq 0$ determines the relative importance of the upper level problem relative to the lower level problem.
-A decreasing sequence of regularization parameters $(\sigma_{k})_{k\geq 1}$ induces a corresponding sequence of operators ${\mathsf{V}}_{k}\triangleq{\mathsf{V}}_{\sigma_{k}}$ and convex functions $G_{k}\triangleq G_{\sigma_{k}}$.
-Being a sum of monotone and Lipschitz continuous mappings, ${\mathsf{V}}_{\sigma}$ is $L_{\sigma}\triangleq L_{{\mathsf{F}}_{2}}+\sigma L_{{\mathsf{F}}_{1}}$-Lipschitz and monotone.
-In the same vein, the combined mapping $G_{\sigma}$ is proper convex and lower semi-continuous.
-Furthermore, Assumption 1 guarantees the non-smooth sum rule $\partial G_{k}=\sigma_{k}\partial g_{1}+\partial g_{2}$.
+Our algorithmic design follows the popular Tikhonov regularization approach and, at iteration $k\geq 1$, uses the operators ${\mathsf{V}}_{k}$ and $G_{k}$ defined in (1.1), which combine the upper- and lower-level data.
+A decreasing sequence of regularization parameters $\sigma_{k}\geq 0$ determines the relative importance of the upper-level problem relative to the lower-level problem.
+Being a sum of monotone and Lipschitz continuous mappings, ${\mathsf{V}}_{k}$ is monotone and $L_{k}$-Lipschitz with $L_{k}\triangleq L_{{\mathsf{F}}_{2}}+\sigma_{k}L_{{\mathsf{F}}_{1}}$.
 The conceptual algorithm we propose for solving (P) is described in Algorithm 1.
-The reader should consider this as a description of a method; Once concrete definitions for the step-size $(t_{k})_{k}$ and the regularization parameters $(\sigma_{k})_{k}$ are given, we obtain a bona-fide algorithm from the pseudo-code.
-Importantly, it should be observed that our algorithm requires only one evaluation of operators ${\mathsf{F}}_{1}$, ${\mathsf{F}}_{2}$ per iteration, in contrast to previous works [samadi_improved_2025, alves2025inertial].
-Algorithm 1 Optimistic extragradient for hierarchical HVI’s (P)
-$z^{1}=z^{1/2}$, step-size sequence $(t_{k})_{k\geq 1}$, non-increasing regularization sequence $(\sigma_{k})_{k\geq 1}$ satisfying
+Once concrete definitions for the step-sizes $(t_{k})_{k\geq 1}$ and the regularization parameters $(\sigma_{k})_{k\geq 1}$ are given, we obtain a bona-fide algorithm.
+Importantly, it should be observed that our method only requires one evaluation of operators ${\mathsf{F}}_{1}$, ${\mathsf{F}}_{2}$ per iteration, in contrast to previous works [57, 2].
+Algorithm 1 Optimistic extragradient method for hierarchical HVI’s (P)
 […]
+$\begin{split}&\lim_{k\to\infty}\sigma_{k}=0,\qquad\lim_{T\to\infty}\sum_{k=1}^{T}t_{k}=\infty,\\ &\lim_{T\to\infty}\sigma_{T}\sum_{k=1}^{T}t_{k}=\infty\text{, and }\lim_{T\to\infty}\frac{\sum_{k=1}^{T}t_{k}\sigma_{k}}{\sum_{k=1}^{T}t_{k}}=0.\end{split}$ (3.1)
 Set $k=1$
@@
 end while
+Remark 3.1.
+To simplify the presentation, we implicitly assume that the proximal mapping of $G_{k}$ can be evaluated efficiently.
+In many practically relevant situations this is indeed the case.
+Trivially, if $g_{1}$ or $g_{2}$ is zero, we obtain an easier implementation.
+Other examples include the cases where either $g_{1}$ is an $L_{1}$ norm and $g_{2}$ is an indicator function of a simple set, such as a box or an $L_{2}$-ball [9].
+In general though, the evaluation of the proximal operator of the composite mapping $G_{k}=\sigma_{k}g_{1}+g_{2}$ is not an easy task.
+We can circumvent this problem, by adopting a lifted formulation [21].
+To obtain such a formulation, let $x=[z_{1},z_{2}]\in\mathcal{Z}\otimes\mathcal{Z}\triangleq\mathcal{X}$ and consider the operators
+$\mathsf{A}_{1}(x)=\begin{pmatrix}\partial g_{1}(z_{1})\\  0\end{pmatrix},\mathsf{A}_{2}(x)=\begin{pmatrix}0\\  \partial g_{2}(z_{2})\end{pmatrix}:\mathcal{X}\to 2^{\mathcal{X}},\text{ and }$
+$\mathsf{B}_{1}(x)=\begin{pmatrix}{\mathsf{F}}_{1}(z_{1})\\ 0\end{pmatrix},\mathsf{B}_{2}(x)=\begin{pmatrix}z_{1}-z_{2}\\  {\mathsf{F}}_{2}(z_{2})-z_{1}+z_{2}\end{pmatrix}:\mathcal{X}\to\mathcal{X}.$
+It is easy to see that $\tilde{\mathcal{S}}_{2}=\operatorname{zer}(\mathsf{A}_{2}+\mathsf{B}_{2})=\{(z_{1},z_{2})\in\mathcal{Z}\times\mathcal{Z}|z_{1}=z_{2},\text{ and }z_{2}\in\operatorname{zer}({\mathsf{F}}_{2}+\partial g_{2})\}.$ This means we can obtain $\mathcal{S}_{2}$ as the projection of points in $\tilde{\mathcal{S}}_{2}$ onto its first coordinate.
+Moreover, the problem
+$\text{Find }\bar{x}=[\bar{z}_{1},\bar{z}_{2}]\in\mathcal{X}\text{ s.t.
+}\langle{\mathsf{F}}_{1}(\bar{z}_{1}),z_{1}-\bar{z}_{1}\rangle+g_{1}(z_{1})-g_{1}(\bar{z}_{1})\geq 0\qquad\forall\bar{x}\in\tilde{\mathcal{S}}_{2}$
+shares the same solution as the non-lifted formulation (P) after projection. We can thus approach the solution of the lifted problem by performing the parallel computations
+$x^{k+1/2}=(\operatorname{Id}_{\mathcal{X}}+t_{k}\mathcal{A}_{k})^{-1}(x^{k}-t_{k}\mathcal{B}_{k}(x^{k-1/2})),$
+$x^{k+1}=(\operatorname{Id}_{\mathcal{X}}+t_{k}\mathcal{A}_{k})^{-1}(x^{k}-t_{k}\mathcal{B}_{k}(x^{k+1/2})),$
+where $\mathcal{A}_{k}\triangleq\sigma_{k}\mathsf{A}_{1}+\mathsf{A}_{2}$ and $\mathcal{B}_{k}\triangleq\sigma_{k}\mathsf{B}_{1}+\mathsf{B}_{2}$.
+Thanks to the direct product structure of the lifting, it is easy to see that the resolvents indeed involve parallel computations on the individual factors.
+As an illustration, the update of $x^{k+1/2}=[z_{1}^{k+1/2},z_{2}^{k+1/2}]$ reads explicitly as
+$z_{1}^{k+1/2}$ $=\operatorname{\textsf{prox}}_{t_{k}\sigma_{k}g_{1}}(z_{1}^{k}-t_{k}(z_{1}^{k-1/2}-z_{2}^{k-1/2}+\sigma_{k}{\mathsf{F}}_{1}(z_{1}^{k-1/2}))),$
+$z^{k+1/2}_{2}$ $=\operatorname{\textsf{prox}}_{t_{k}g_{2}}(z_{2}^{k}-t_{k}(z_{2}^{k-1/2}-z_{1}^{k-1/2}+{\mathsf{F}}_{2}(z_{2}^{k-1/2})).$
+$\Diamond$
 3.1 Statement of the main results
-The main result of this paper are two complexity statements for the averaged iterates generated by Algorithm 1 in terms of the introduced gap functions.
-We work in a specific geometric setting, involving key quantities associated with the lower level problem $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$.
-Associated to the bifunction $H^{({\mathsf{F}}_{2},g_{2})}$, Appendix A defines the mapping
+The main result of this paper are two complexity statements for the averaged iterates generated by Algorithm 1 in terms of the gap functions (2.7) and (2.8).
+We work in a specific geometric setting in which we assume some structure on the solution set of the lower level problem $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$.
+Associated to the bifunction $H^{({\mathsf{F}}_{2},g_{2})}$, Appendix B introduces the mapping
 $\varphi^{({\mathsf{F}}_{2},g_{2})}(z,u)\triangleq\sup_{y\in\operatorname{dom}(g_{2})}\{H^{({\mathsf{F}}_{2},g_{2})}(z,y)+\langle y,u\rangle\}.$ (3.2)
-This function encodes dual properties of the variational inequality, since it holds that (cf. eq. (A.1))
+This function encodes dual properties of $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$, since it holds that (cf. eq. (B.1))
 $\varphi^{({\mathsf{F}}_{2},g_{2})}(z,u)\leq\sup_{y\in\operatorname{dom}(g_{2})}\{\langle y,u\rangle-H^{({\mathsf{F}}_{2},g_{2})}(y,z)\}=\left(H^{({\mathsf{F}}_{2},g_{2})}(\bullet,z)\right)^{\ast}(u).$ (3.3)
-The following summability condition is essentially due to [Attouch:2010aa, AttCzarPey11]:
+The following summability condition is essentially due to [6, 5]:
 Assumption 3.
 [Attouch-Czarnecki condition] The step size sequence $(t_{k})_{k\geq 1}$ and the regularization sequence $(\sigma_{k})_{k\geq 1}$ satisfy
-$(\forall p^{*}\in\text{Range}(\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})):\;\sum_{k=1}^{\infty}t_{k}\left[\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})\right]<\infty.$ (3.4)
-To understand the meaning of Assumption 3, it is instructive to specialize our setting to the simple bilevel optimization case (cf.
+$(\forall p\in\text{Range}(\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}})):\;\sum_{k=1}^{\infty}t_{k}\left[\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p)-\operatorname{\mathtt{s}}(\sigma_{k}p|\mathcal{S}_{2})\right]<\infty.$ (3.4)
+To understand the meaning of Assumption 3, it is instructive to specialize our setting to the simple convex bilevel optimization case (cf.
 Section 1.1.1).
-In that case, the data of the lower level problem are identified with ${\mathsf{F}}_{2}=\nabla f_{2}$ and $\mathcal{S}_{2}=\operatorname*{argmin}_{z}(f_{2}+g_{2})(z)$, and thus
-$H^{({\mathsf{F}}_{2},g_{2})}(x,y)$ $\leq f_{2}(x)-f_{2}(y)+g_{2}(x)-g_{2}(y)\triangleq\hat{f}_{2}(x)-\hat{f}_{2}(y)$
-$=(\hat{f}_{2}-\min\hat{f}_{2})(x)-(\hat{f}_{2}-\min\hat{f}_{2})(y).$
+In that case, we have ${\mathsf{F}}_{2}=\nabla f_{2}$ and $\mathcal{S}_{2}=\operatorname*{argmin}_{z}(f_{2}+g_{2})(z)$, and thus
+$H^{({\mathsf{F}}_{2},g_{2})}(y,z)$ $\leq f_{2}(y)-f_{2}(z)+g_{2}(y)-g_{2}(z)\triangleq\hat{f}_{2}(y)-\hat{f}_{2}(z)$
+$=(\hat{f}_{2}-\min\hat{f}_{2})(y)-(\hat{f}_{2}-\min\hat{f}_{2})(z).$
 Hence, $\varphi^{({\mathsf{F}}_{2},g_{2})}(z,p)\leq(\hat{f}_{2}-\min\hat{f}_{2})(z)+(\hat{f}_{2}-\min\hat{f}_{2})^{*}(p)$.
@@
 $\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p)-\operatorname{\mathtt{s}}(\sigma_{k}p|\mathcal{S}_{2})\leq(\hat{f}_{2}-\min\hat{f}_{2})^{*}(\sigma_{k}p)-\operatorname{\mathtt{s}}(\sigma_{k}p|\mathcal{S}_{2}).$
-In [AttCzarPey11, Peypouquet:2012aa, boct2025accelerating] the following summability condition is imposed
+In [5, 55, 13] the following summability condition is imposed
 $\sum_{k=1}^{\infty}t_{k}\left[(\hat{f}_{2}-\min\hat{f}_{2})^{*}(\sigma_{k}p)-\operatorname{\mathtt{s}}(\sigma_{k}p|\mathcal{S}_{2})\right]<\infty.$
-We see that this condition is stronger than our condition (3.4).
+We see that this condition is more restrictive than our condition (3.4).
 Remark 3.2.
-Assumption 3 looks quite daunting to verify, but as already observed in [boct2025accelerating], it fits very natural to the geometric setting of this paper.
+Assumption 3 looks quite daunting to verify, but as already observed in [13], it fits very natural to the geometric setting of this paper.
 Indeed, let us assume that $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp, with $\alpha>0$ and $\rho>1$.
@@
 This shows that under the $(\alpha,\rho)$-weak sharpness condition, the summability condition (3.4) is satisfied whenever $\sum_{k\geq 1}t_{k}\sigma^{\frac{\rho}{\rho-1}}_{k}<\infty.$ $\Diamond$
-Assumption 3 will allow us to prove that the sequence generated by the algorithm is bounded, an important step towards deriving convergence rates in terms of the averaged trajectory
+Next, we assume the non-smooth part $g_{1}$ of $\operatorname{HVI}({\mathsf{F}}_{1},g_{1})$ to satisfy a sort of finite variation property.
+Specifically, we define the variation over the set $\mathcal{A}\times\mathcal{B}\subseteq\mathcal{Z}\times\mathcal{Z}$ as
+$|.\operatorname{\mathsf{Var}}(g_{1}|\mathcal{A}\times\mathcal{B})\triangleq\sup_{(x,y)\in\mathcal{A}\times\mathcal{B}}\absolutevalue{g_{1}(x)-g_{1}(y)}$.
+Assumption 4.
+The set $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is closed, and for any compact sets $\mathcal{A},\mathcal{B}\subset\mathcal{Z}$ satisfying $\mathcal{A},\mathcal{B}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$, we have $\operatorname{\mathsf{Var}}(g_{1}|\mathcal{A}\times\mathcal{B})<\infty$.
+Assumption 3 will allow us to prove that the sequences generated Algorithm 1 are bounded, an important step towards deriving convergence rates in terms of the averaged trajectory
 $\bar{z}^{K}\triangleq\frac{1}{T_{K}}\sum_{k=1}^{K}t_{k}z^{k+1/2},\qquad T_{K}\triangleq\sum_{k=1}^{K}t_{k}.$ (3.5)
@@
 Consider problem (P).
-Let Assumptions 1-2 hold.
-Let additionally either Assumption 3 hold or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be compact.
-Consider the step size $(t_{k})_{k\geq 1}$ with $8t_{k}^{2}L_{k}^{2}\leq 1$ for all $k\geq 1$, and let $(\sigma_{k})_{k\geq 1}$ be a non-increasing sequence satisfying (3.1).
+Let Assumptions 1, 2, and 4 hold.
+Assume additionally that either Assumptions 3 hold, or that $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is compact.
+Let $(\bar{z}^{k})_{k\geq 1}$ as in (3.5) be generated by Algorithm 1 with $16t_{k}^{2}L_{k}^{2}\leq 1$ for all $k\geq 1$.
   • (i)
-    Consider a nonempty compact set $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$. Then, there exists a constant $C_{\mathcal{U}_{2}}>0$ for which
+    Let $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be non empty and compact. Then, there exists a constant $C_{\mathcal{U}_{2}}>0$ for which
     $\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2T_{K}}+\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}C_{\mathcal{U}_{2}}.$ (3.6)
-    In particular, under (3.1) all weak accumulation points $\tilde{z}$ of $(\bar{z}^{k})$ satisfy $\Theta_{\rm Feas}(\tilde{z}|\mathcal{U}_{2})\leq 0$.
   • (ii)
@@
   • (iii)
-    If the lower level solution set $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp, then
-    $B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2T_{K}(\alpha/\rho)}+\frac{C_{\mathcal{U}_{2}}\sum_{k=1}^{K}t_{k}\sigma_{k}}{(\alpha/\rho)T_{K}}\right]^{1/\rho}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{T_{K}\sigma_{K}}.$ (3.8)
-    In particular, all weak accumulation points $\tilde{z}$ of $(\bar{z}^{k})$ satisfy $\Theta_{\rm Opt}(\tilde{z}|\mathcal{U}_{1}\cap\mathcal{S}_{2})=0$.
+    If the lower level solution set $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp and $\mathcal{U}_{2}$ in (i) is s.t.
+$\mathcal{U}_{2}\cap\mathcal{S}_{2}\neq\varnothing$, then for the same set $\mathcal{U}_{1}$ as in (ii), we have the lower bound
+    $-B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2T_{K}(\alpha/\rho)}+\frac{C_{\mathcal{U}_{2}}\sum_{k=1}^{K}t_{k}\sigma_{k}}{(\alpha/\rho)T_{K}}\right]^{\frac{1}{\rho}}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2}).$ (3.8)
 Remark 3.3.
-Theorem 3.1 derives a rate in terms of the feasibility and the optimality gap of the hierarchical VI problem (P).
+Theorem 3.1 derives a rate in terms of the feasibility and the optimality gap of the hierarchical HVI problem (P).
 These rates are established relative to arbitrary compact sets $\mathcal{U}_{1},\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$.
@@
 The constants involved in (3.6) and (3.7) depend on the Lipschitz modulus of the operators ${\mathsf{F}}_{i}$, and diameter-like constants, which can be exhibited thanks to the boundedness of the trajectory.
+The stepsize condition is satisfied whenever $t_{k}\leq\frac{1}{4}\frac{1}{L_{{\mathsf{F}}_{2}}+\sigma_{k}L_{{\mathsf{F}}_{1}}}$.
+This can be achieved by a time-varying stepsize or, since $\sigma_{k}$ is decreasing, by a fixed stepsize, e.g., $t_{k}=t\leq\frac{1}{4}\frac{1}{L_{{\mathsf{F}}_{2}}+\sigma_{1}L_{{\mathsf{F}}_{1}}}$.
 $\Diamond$
-Choosing a specific regularization sequence $(\sigma_{k})_{k\geq 1}$ allows us to turn the estimates from Theorem 3.1 into concrete complexity bounds.
+Choosing a specific regularization sequence $(\sigma_{k})_{k\geq 1}$ allows us to turn the estimates from Theorem 3.1 into concrete convergence rates.
 In detail, we consider the constant step size policy $t_{k}=t$ and regularization sequence
@@
 Corollary 3.2.
-Let the same Assumptions as in Theorem 3.1 be in place.
-Assume that the regularization sequence $(\sigma_{k})_{k\geq 1}$ is chosen according to (3.9), and $t_{k}=t\leq\frac{1}{\sqrt{8}L_{1}}$ for all $k\geq 1$.
+Suppose the same assumptions as in Theorem 3.1 hold.
+Assume that the regularization sequence $(\sigma_{k})_{k\geq 1}$ is chosen according to (3.9), and $t_{k}=t\leq\frac{1}{4(L_{{\mathsf{F}}_{2}}+\sigma_{1}L_{{\mathsf{F}}_{2}})}$ for all $k\geq 1$.
 Then, we have
 $\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt}+\frac{aC_{\mathcal{U}_{2}}}{(1-\delta)(K+b)^{\delta}},$ (3.10)
-$-B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt(\alpha/\rho)}+\frac{aC_{\mathcal{U}_{2}}}{(\alpha/\rho)(1-\delta)(K+b)^{\delta}}\right]^{1/\rho}\overset{(*)}{\leq}\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{a(K+b)^{1-\delta}},$ (3.11)
-where the inequality ()(*) holds under the $(\alpha,\rho)$-weakly sharpness assumption.
-Remark $3.$
-Under the regularization sequence (3.9) and if $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weak sharp, a sufficient condition for Assumption 3 to hold is $1>\delta>1-\frac{1}{\rho}$ (cf.
+$-B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt(\alpha/\rho)}+\frac{aC_{\mathcal{U}_{2}}}{(\alpha/\rho)(1-\delta)(K+b)^{\delta}}\right]^{1/\rho}$
+$\hskip 68.00098pt\overset{(*)}{\leq}\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{a(K+b)^{1-\delta}},$ (3.11)
+where the inequality ()(*) holds under the $(\alpha,\rho)$-weak sharpness assumption.
+Remark 3.4.
+Under the regularization sequence (3.9) and $(\alpha,\rho)$-weak sharpness of the lower-level solution set $\mathcal{S}_{2}$, a sufficient condition for Assumption 3 to hold is $1>\delta>1-\frac{1}{\rho}$ (cf.
 Remark 3.2).
-This agrees with the analysis in the potential case of [boct2025accelerating].
+This coincides with the analysis for the optimization case in [13, Theorem 2.3].
 $\Diamond$
+Theorem 3.1 is instrumental to draw conclusions about the asymptotic behavior of the sequences generated by Algorithm 1.
+However, this statement requires one more (mild) assumption on the data defining problem (P).
+Assumption 5.
+The non-smooth functions in the definition of (P) satisfy the hierarchical coherence condition $\operatorname{dom}(g_{2})\subseteq\operatorname{dom}(g_{1}).$
+Theorem 3.3.
+Consider problem (P).
+Let Assumptions 1-3 and the hierarchic coherence condition Assumption 5 hold.
+Let $(z^{k})_{k\geq 1}$ and $(z^{k+1/2})_{k\geq 1}$ be generated by Algorithm 1 with $16t_{k}^{2}L_{k}^{2}\leq 1$ for all $k\geq 1$.
+Then, the following holds true:
+  • (i)
+    Asymptotic feasibility: All weak accumulation points of $(z^{k+1/2})_{k\geq 1}$ are contained in $\mathcal{S}_{2}$.
+  • (ii)
+    Ergodic average optimality: If additionally Assumption 4 holds, then the sequence $(\bar{z}^{k})_{k\geq 1}$ converges weakly to an element in $\mathcal{S}_{1}$.
 4 Proofs of the main results
 4.1 Energetic bounds
-In this section, we develop the necessary energy bounds for the sequence $(z^{k})_{k\geq 1}$ generated by the optimistic extragradient method.
-Key is the step size condition $8t_{k}^{2}L^{2}_{k}\leq 1$, where $L_{k}\triangleq L_{{\mathsf{F}}_{2}}+\sigma_{k}L_{{\mathsf{F}}_{1}}$ is the combined Lipschitz constant of the operator ${\mathsf{V}}_{k}$.
-This condition implies $t_{k}\leq\frac{1}{\sqrt{8}}\frac{1}{L_{{\mathsf{F}}_{2}}+\sigma_{k}L_{{\mathsf{F}}_{1}}}\in\frac{1}{\sqrt{8}}\left[\frac{1}{L_{{\mathsf{F}}_{2}}+\sigma_{1}L_{{\mathsf{F}}_{1}}},\frac{1}{L_{{\mathsf{F}}_{2}}}\right].$
+In this section, we develop the energy bounds for the sequences $(z^{k})_{k\geq 1}$, $(z^{k+1/2})_{k\geq 1}$ generated by the optimistic extragradient method.
+To simplify the derivations, we introduce the shortcut notation
+$E_{k}(z)\triangleq\frac{1}{2}\norm{z^{k}-z}^{2}\hskip 17.00024pt\forall k\geq 1,$ (4.1)
+$\Psi_{k}(z)\triangleq\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle+G_{k}(z^{k+1/2})-G_{k}(z)\hskip 17.00024pt\forall k\geq 1,$ (4.2)
+$D_{k}\triangleq t^{2}_{k-1}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\hskip 17.00024pt\forall k\geq 2.$ (4.3)
 Lemma 4.1.
-Let $(\sigma_{k})_{k\geq 1}$ be non-increasing and $(t_{k})_{k\geq 1}$ be a step size policy. Let also $z\in\mathcal{Z}$ be arbitrary.
-  • (i)
-    For all $k\geq 1$ we have
-    $\begin{split}\frac{1}{2}\norm{z^{k+1}-z}^{2}&\leq\frac{1}{2}\norm{z^{k}-z}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}\\ &+t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).\end{split}$ (4.1)
-  • (ii)
-    If $8t^{2}_{k}L_{k}^{2}\leq 1$ then for all $k\geq 2$,
-    $\begin{split}\frac{1}{2}\norm{z^{k+1}-z}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}&\leq\frac{1}{2}\norm{z^{k}-z}^{2}+\frac{t^{2}_{k-1}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\\ &-t_{k}\left(\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle+G_{k}(z^{k+1/2})-G_{k}(z)\right)\\ &-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}.\end{split}$ (4.2)
+[…]
+$\begin{split}E_{k+1}(z)\leq E_{k}(z)-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}-\frac{1}{4}\norm{z^{k+1}-z^{k+1/2}}^{2}+D_{k+1}-t_{k}\Psi_{k}(z)\end{split}$ (4.4)
+If additionally $16t^{2}_{k}L_{k}^{2}\leq 1$ and defining $D_{1}\triangleq 0$, we have, for all $k\geq 1$,
+$\begin{split}E_{k+1}(z)+D_{k+1}&\leq E_{k}(z)+D_{k}-t_{k}\Psi_{k}(z)-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}\\ &-\frac{1}{4}\norm{z^{k+1}-z^{k+1/2}}^{2}.\end{split}$ (4.5)
 Proof.
-(i) Applying (2.1) to the iterates $z^{k+1/2}$ and $z^{k+1}$, respectively, we obtain
-$\langle z^{k+1}-z^{k},z^{k+1}-z\rangle\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1})\right)$ (4.3)
-$\langle z^{k+1/2}-z^{k},z^{k+1/2}-z\rangle\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k-1/2}),z-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$
-Using $z=z^{k+1}$ in the second inequality, we have
-$\langle z^{k+1/2}-z^{k},z^{k+1/2}-z^{k+1}\rangle\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z^{k+1})-G_{k}(z^{k+1/2})\right).$ (4.4)
+Proof of (4.4). Applying (2.1) to the iterate $z^{k+1}$, we obtain $\forall z\in\mathcal{Z}$
+$\langle z^{k+1}-z^{k},z^{k+1}-z\rangle\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1})\right).$ (4.6)
+Doing the same for the iterate $z^{k+1/2}$, but substituting $z=z^{k+1}$, we get $\forall z\in\mathcal{Z}$
+$\begin{split}\langle z^{k+1/2}-z^{k},z^{k+1/2}-z^{k+1}\rangle&\leq\langle t_{k}{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle\\ &+t_{k}\left(G_{k}(z^{k+1})-G_{k}(z^{k+1/2})\right).\end{split}$ (4.7)
 Now, we observe that
-$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k+1}-z^{k}+z^{k}-z}^{2}$
-$=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+2\langle z^{k+1}-z^{k},z^{k+1}-z\rangle.$
-In the same spirit, we see
-$\norm{z^{k+1}-z^{k}}^{2}=\norm{z^{k+1}-z^{k+1/2}}^{2}+\norm{z^{k+1/2}-z^{k}}^{2}+2\langle z^{k+1}-z^{k+1/2},z^{k+1/2}-z^{k}\rangle.$
-Combining these two identities with (4.3) and (4.4), we can continue the estimation as
+$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+2\langle z^{k+1}-z^{k},z^{k+1}-z\rangle,$ (4.8)
+$\norm{z^{k+1}-z^{k}}^{2}$ $=\norm{z^{k+1}-z^{k+1/2}}^{2}+\norm{z^{k+1/2}-z^{k}}^{2}$
+$+2\langle z^{k+1}-z^{k+1/2},z^{k+1/2}-z^{k}\rangle.$
+Combining these two identities with (4.6) and (4.7), we can continue as
 $\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k+1/2}}^{2}-\norm{z^{k+1/2}-z^{k}}^{2}$
@@
 Adding and subtracting $z^{k+1/2}$ in the second inner product, leaves us with
-$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k+1/2}}^{2}-\norm{z^{k+1/2}-z^{k}}^{2}$
-$+2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle$
-$+2t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$
-Applying Young’s inequality $\langle a,b\rangle\leq\frac{c}{2}\norm{a}^{2}+\frac{1}{2c}\norm{b}^{2}$ with $c=t_{k}$, we obtain
-$2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle\leq t^{2}_{k}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}+\norm{z^{k+1}-z^{k+1/2}}^{2}.$
-Hence, we obtain (4.1)
-$\frac{1}{2}\norm{z^{k+1}-z}^{2}$ $\leq\frac{1}{2}\norm{z^{k}-z}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
-$+t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$
-(ii) Observe that for $L_{k}>0$,
-$\frac{t_{k}^{2}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$ $=t^{2}_{k}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
-$\leq t_{k}^{2}L_{k}^{2}\norm{z^{k+1/2}-z^{k-1/2}}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
-$\leq 2t_{k}^{2}L^{2}_{k}\norm{z^{k+1/2}-z^{k}}^{2}+2t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}$
-$-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2},$
+$\norm{z^{k+1}-z}^{2}\leq\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k+1/2}}^{2}-\norm{z^{k+1/2}-z^{k}}^{2}$
+$\qquad+2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle$
+$\qquad+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+2t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$ (4.10)
+Applying Young’s inequality $\langle a,b\rangle\leq\frac{c}{2}\norm{a}^{2}+\frac{1}{2c}\norm{b}^{2}$ with $c=2t_{k}$ and (4.3), we obtain
+$2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle$
+$\leq 2t^{2}_{k}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}+\frac{1}{2}\norm{z^{k+1}-z^{k+1/2}}^{2}$
+$=2D_{k+1}+\frac{1}{2}\norm{z^{k+1}-z^{k+1/2}}^{2}.$
+Plugging this inequality and definitions (4.1), (4.2) into the previous inequality gives us (4.4).
+Proof of (4.5). Observe that by (4.3)
+$D_{k+1}$ $=2t^{2}_{k}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-D_{k+1}\leq 2t_{k}^{2}L_{k}^{2}\norm{z^{k+1/2}-z^{k-1/2}}^{2}-D_{k+1}$
+$\leq 4t_{k}^{2}L^{2}_{k}\norm{z^{k+1/2}-z^{k}}^{2}+4t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-D_{k+1},$
 where the first inequality follows from the $L_{k}$-Lipschitz continuity of ${\mathsf{V}}_{k}$, and the second inequality follows from $(a+b)^{2}\leq 2a^{2}+2b^{2}$. We can therefore continue,
-$\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq 2t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
-$+\frac{1}{2}\left(4t_{k}^{2}L^{2}_{k}-1\right)\norm{z^{k+1/2}-z^{k}}^{2}$ (4.6)
-$\leq 2t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
-$-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2},$ (4.7)
-where the last inequality is due to the step size choice satisfying $4t_{k}^{2}L^{2}_{k}\leq\frac{1}{2}$.
-Moreover, using the definition of $z^{k}$ and $z^{k-1/2}$, the non-expansiveness of the proximal mapping (LABEL:eq:nonexpansive), and the step size assumption $8t^{2}_{k}L_{k}^{2}\leq 1$ yields $4t^{2}_{k}t^{2}_{k-1}L^{2}_{k}\leq\frac{t^{2}_{k-1}}{2}$.
-This results in
-$2t_{k}^{2}L_{k}^{2}\norm{z^{k}-z^{k-1/2}}^{2}\leq 2t^{2}_{k}t^{2}_{k-1}L^{2}_{k}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\leq\frac{t^{2}_{k-1}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}.$ (4.8)
-Combining this inequality with (4.7), plugging the result into (4.1), and rearranging terms, we obtain (4.2)
-$\frac{1}{2}\norm{z^{k+1}-z}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$ $\leq\frac{1}{2}\norm{z^{k}-z}^{2}+\frac{t^{2}_{k-1}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}$
-$-t_{k}\left(\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle+G_{k}(z^{k+1/2})-G_{k}(z)\right)$
-$-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}.$
+$D_{k+1}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$
+$\leq 4t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-D_{k+1}+\left(4t_{k}^{2}L^{2}_{k}-\frac{1}{2}\right)\norm{z^{k+1/2}-z^{k}}^{2}$ (4.11)
+$\leq 4t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-D_{k+1}-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2},$ (4.12)
+where the last inequality is due to the step size choice satisfying $4t_{k}^{2}L^{2}_{k}\leq\frac{1}{4}$.
+If $k\geq 2$, using the definition of $z^{k}$ and $z^{k-1/2}$, the non-expansiveness of the proximal mapping, and that the step size assumption $16t^{2}_{k}L_{k}^{2}\leq 1$ yields $4t^{2}_{k}L^{2}_{k}\leq 1$, we obtain
+$4t_{k}^{2}L_{k}^{2}\norm{z^{k}-z^{k-1/2}}^{2}\leq 4t^{2}_{k}L^{2}_{k}t^{2}_{k-1}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\leq D_{k}.$ (4.13)
+The same inequality holds for $k=1$ since $D_{1}=0$ and $z^{1/2}=z^{1}$.
+Combining (4.13) with (4.12), plugging the result into (4.4), using the notation (4.2), and rearranging terms, we obtain (4.5).
 $\blacksquare$
-To continue with our energetic bounds, we define the data
-$E_{k}(z)\triangleq\frac{1}{2}\norm{z^{k}-z}^{2}\hskip 17.00024pt\forall k\geq 1,$ (4.9)
-$\Psi_{k}(z)\triangleq\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle+G_{k}(z^{k+1/2})-G_{k}(z)\hskip 17.00024pt\forall k\geq 1,$ (4.10)
-$D_{k}\triangleq\frac{t^{2}_{k-1}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-3/2})-{\mathsf{V}}_{k-1}(z^{k-1/2})}^{2}\hskip 17.00024pt\forall k\geq 2.$ (4.11)
-This allows us to rewrite the statement in Lemma 4.1(ii) concisely as
-$E_{k+1}(z)+D_{k+1}\leq E_{k}(z)+D_{k}-t_{k}\Psi_{k}(z)-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}\hskip 17.00024pt\forall k\geq 2.$ (4.12)
-We now extend the validity of (4.12) to $k\geq 1$.
-Lemma 4.2.
-Define $D_{1}\triangleq 0$. Then, for all $k\geq 1$ relation (4.12) holds.
-Proof.
-It is only left to show that the inequality holds for $k=1$. For $k=1$ (4.12) reduces to
-$E_{2}(z)+D_{2}\leq E_{1}(z)+D_{1}-t_{1}\Psi_{1}(z)-\frac{1}{4}\norm{z^{3/2}-z^{1}}^{2}.$ (4.13)
-Using (4.7) we have that
-$\frac{t_{1}^{2}}{2}\norm{{\mathsf{V}}_{1}(z^{1/2})-{\mathsf{V}}_{1}(z^{3/2})}^{2}-\frac{1}{2}\norm{z^{3/2}-z^{1}}^{2}$ $\leq 2t_{1}^{2}L^{2}_{1}\norm{z^{1}-z^{1/2}}^{2}-\frac{t^{2}_{1}}{2}\norm{{\mathsf{V}}_{1}(z^{1/2})-{\mathsf{V}}_{1}(z^{3/2})}^{2}-\frac{1}{4}\norm{z^{3/2}-z^{1}}^{2}$
-$=-\frac{t^{2}_{1}}{2}\norm{{\mathsf{V}}_{1}(z^{1/2})-{\mathsf{V}}_{1}(z^{3/2})}^{2}-\frac{1}{4}\norm{z^{3/2}-z^{1}}^{2}$
-where the last equality is due to $z^{1/2}=z^{1}$. Plugging this result back into (4.1), we arrive at
-$\frac{1}{2}\norm{z^{2}-z}^{2}$ $\leq\frac{1}{2}\norm{z^{1}-z}^{2}-t_{1}\left(\langle{\mathsf{V}}_{1}(z^{3/2}),z^{3/2}-z\rangle+G_{1}(z^{3/2})-G_{1}(z)\right)-\frac{t_{1}^{2}}{2}\norm{{\mathsf{V}}_{1}(z^{1/2})-{\mathsf{V}}_{1}(z^{3/2})}^{2}$
-$-\frac{1}{4}\norm{z^{3/2}-z^{1}}^{2}.$
-Using the definition of $D_{1}$, we arrive at (4.13). $\blacksquare$
 4.2 Towards proving Theorem 3.1
 Establishing the rate on the feasibility gap (3.6)
-To obtain a bound on the feasibility gap, we start with an Opial-like Lemma which holds under Assumption 3.
+To obtain a bound on the feasibility gap, we start with an Opial-like Lemma.
 Lemma 4.2.
-Let $z^{*}\in\mathcal{S}_{1}$ arbitrary and $(z^{k},z^{k+1/2})_{k\geq 1}$ be the sequence generated by Algorithm 1. Let also Assumption 3 hold. Define
-$W_{k}(z^{*})\triangleq E_{k}(z^{*})+D_{k}\hskip 17.00024pt\forall k\geq 1.$ (4.14)
+Let Assumptions 1-3 hold.
+Let $z^{*}\in\mathcal{S}_{1}$ be arbitrary and let $(z^{k})_{k\geq 1}$ and $(z^{k+1/2})_{k\geq 1}$ be generated by Algorithm 1 with $16t^{2}_{k}L_{k}^{2}\leq 1$.
+Define $D_{1}=0$ and
+$W_{k}(z)\triangleq E_{k}(z)+D_{k}\hskip 17.00024pt\forall k\geq 1.$ (4.14)
 Then, the following statements hold:
@@
   2. (b)
-    $\lim_{k\to\infty}\norm{z^{k+1/2}-z^{k}}=0$;
+    $\lim_{k\to\infty}\norm{z^{k+1/2}-z^{k}}=\lim_{k\to\infty}\norm{z^{k+1}-z^{k+1/2}}=\lim_{k\to\infty}\norm{z^{k+1/2}-z^{k-1/2}}=0$;
   3. (c)
     $(z^{k})_{k\geq 1}$ and $(z^{k+1/2})_{k\geq 1}$ are bounded.
+  4. (d)
+    $\lim_{k\to\infty}\norm{z^{k}-z^{\ast}}$ and $\lim_{k\to\infty}\norm{z^{k+1/2}-z^{*}}$ both exist and are equal.
 Proof.
-Pick $z^{\ast}\in\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$ so that there exists $p^{\ast}\in\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}(z^{\ast})$ such that ${\mathsf{F}}_{1}(z^{\ast})-p^{\ast}\in\partial g_{1}(z^{\ast})$.
+Pick $z^{\ast}\in\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$ so that there exists $p^{\ast}\in\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}(z^{\ast})$ s.t.
+${\mathsf{F}}_{1}(z^{\ast})-p^{\ast}\in\partial g_{1}(z^{\ast})$.
 Using the monotonicity of ${\mathsf{F}}_{1}$ and the convex subgradient inequality for $g_{1}$, we obtain
-$\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle$ $+g_{1}(z^{k+1/2})-g_{1}(z^{\ast})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle$
+$-H^{({\mathsf{F}}_{1},g_{1})}(z^{\ast},z^{k+1/2})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle$
+$=\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle+g_{1}(z^{k+1/2})-g_{1}(z^{\ast})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle$
 $\geq\langle{\mathsf{F}}_{1}(z^{\ast}),z^{k+1/2}-z^{\ast}\rangle+g_{1}(z^{k+1/2})-g_{1}(z^{\ast})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle\geq 0.$
-Combined with the energy inequality (4.12) we conclude
-$E_{k+1}(z^{\ast})+D_{k+1}$ $\leq E_{k+1}(z^{\ast})+D_{k+1}+t_{k}\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle+g_{1}(z^{k+1/2})-g_{1}(z^{\ast})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right)$
-$\leq E_{k}(z^{\ast})+D_{k}-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}$
-$-t_{k}\left(\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle+g_{2}(z^{k+1/2})-g_{2}(z^{\ast})-\langle\sigma_{k}p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right).$ (4.15)
+Combining with the energy inequality (4.5) and the fact that
+$\Psi_{k}(z)=-\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z,z^{k+1/2})-H^{({\mathsf{F}}_{2},g_{2})}(z,z^{k+1/2}),$ (4.15)
+we continue
+$E_{k+1}(z^{\ast})+D_{k+1}$
+$\leq E_{k+1}(z^{\ast})+D_{k+1}+t_{k}\sigma_{k}\left(-H^{({\mathsf{F}}_{1},g_{1})}(z^{\ast},z^{k+1/2})+\langle p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right)$
+$\leq E_{k}(z^{\ast})+D_{k}-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}-\frac{1}{4}\norm{z^{k+1}-z^{k+1/2}}^{2}$
+$+t_{k}\left(H^{({\mathsf{F}}_{2},g_{2})}(z^{\ast},z^{k+1/2})+\langle\sigma_{k}p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right).$ (4.16)
 Since $p^{\ast}\in\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}(z^{\ast})$, we know from (2.2) that $\langle\sigma_{k}p^{*},z^{*}\rangle=\operatorname{\mathtt{s}}(\sigma_{k}p^{*}|\mathcal{S}_{2})$ and
-$-\left(\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z^{\ast}\rangle+g_{2}(z^{k+1/2})-g_{2}(z^{\ast})-\langle\sigma_{k}p^{\ast},z^{k+1/2}-z^{\ast}\rangle\right)$
-$=\langle\sigma_{k}p^{\ast},z^{k+1/2}\rangle+\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{\ast}-z^{k+1/2}\rangle+g_{2}(z^{\ast})-g_{2}(z^{k+1/2})-\langle\sigma_{k}p^{\ast},z^{\ast}\rangle$
+$H^{({\mathsf{F}}_{2},g_{2})}(z^{\ast},z^{k+1/2})+\langle\sigma_{k}p^{\ast},z^{k+1/2}-z^{\ast}\rangle$
 $=\langle\sigma_{k}p^{\ast},z^{k+1/2}\rangle+H^{({\mathsf{F}}_{2},g_{2})}(z^{\ast},z^{k+1/2})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})$
 $\leq\sup_{z\in\operatorname{dom}(g_{2})}\left(\langle\sigma_{k}p^{\ast},z\rangle+H^{({\mathsf{F}}_{2},g_{2})}(z^{\ast},z)\right)-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})$
-$=\varphi^{({\mathsf{F}}_{2},g_{2})}(z^{\ast},\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})$
-$\leq\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2}).$
-We note that (A.2) in Appendix A guarantees that the upper bound is informative, in the sense that
-$\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})\geq 0.$
-Plugging this back into (4.15), we obtain
-$E_{k+1}(z^{\ast})+D_{k+1}$ $\leq E_{k}(z^{\ast})+D_{k}-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}+t_{k}\left[\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})\right].$ (4.16)
-We see from these estimates that the summability condition (3.4) gives a bound on the sequence $E_{k}(z^{\ast})+D_{k}$. Defining $W_{k}(z^{\ast})\triangleq E_{k}(z^{\ast})+D_{k}$, and
-$a_{k}\triangleq t_{k}\left[\sup_{z\in\mathcal{S}_{2}}\varphi^{({\mathsf{F}}_{2},g_{2})}(z,\sigma_{k}p^{\ast})-\operatorname{\mathtt{s}}(\sigma_{k}p^{\ast}|\mathcal{S}_{2})\right],$
-we obtain from (4.16)
-$W_{k+1}(z^{\ast})-W_{k}(z^{\ast})+\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}\leq a_{k}.$
-From this, summability of $a_{k}$, and non-negativity of $W_{k}(z^{*})$ it follows from Lemma 2.1 that $\lim_{k\to\infty}W_{k}(z^{\ast})=W_{\infty}(z^{\ast})$ exists, and $\lim_{k\rightarrow\infty}\norm{z^{k+1/2}-z^{k}}=0$.
+[…]
+We note that (B.2) in Appendix B guarantees that the upper bound is informative, in the sense that $c_{k}\geq 0$. Plugging this back into (4.16) and using (4.14), we obtain
+$W_{k+1}(z^{\ast})\leq W_{k}(z^{\ast})-\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}-\frac{1}{4}\norm{z^{k+1}-z^{k+1/2}}^{2}+t_{k}c_{k}.$
+The summability condition (3.4) implies $\sum_{k}t_{k}c_{k}<\infty$.
+Hence, the non-negativity of $W_{k}(z^{*})$ allows us to deduce from Lemma A.1 that $\lim_{k\to\infty}W_{k}(z^{\ast})=W_{\infty}(z^{\ast})$ exists, and $\lim_{k\rightarrow\infty}\norm{z^{k+1/2}-z^{k}}=\lim_{k\to\infty}\norm{z^{k+1}-z^{k+1/2}}=0$.
+By the triangle inequality, the latter also implies that $\lim_{k\to\infty}\norm{z^{k+1/2}-z^{k-1/2}}=0$.
 By the definition of $E_{k}(z^{*})$, we further obtain that the sequences $(z^{k})_{k\geq 1}$ and $(z^{k+1/2})_{k\geq 1}$ are bounded.
-$\blacksquare$
+By (4.3), $L_{k}$-Lipschitzness of ${\mathsf{V}}_{k}$ and item (b), we have, as $k\to\infty$,
+$0\leq D_{k+1}$ $=t_{k}^{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}\leq t_{k}^{2}L^{2}_{k}\norm{z^{k+1/2}-z^{k-1/2}}^{2}\to 0.$
+Since $E_{k}(z^{*})=W_{k}(z^{*})-D_{k}$, it thus follos $\lim_{k\rightarrow\infty}E_{k}(z^{*})=W_{\infty}(z^{*}).$ Moreover, from $\norm{z^{k}-z^{k+1/2}}\to 0$, and (4.1), we immediately deduce
+$\lim_{k\rightarrow\infty}\norm{z^{k+1/2}-z^{*}}^{2}$ $=\lim_{k\rightarrow\infty}\norm{z^{k}-z^{*}}^{2}+\norm{z^{k}-z^{k+1/2}}^{2}+\langle z^{k}-z^{k+1/2},z^{k}-z^{*}\rangle$
+$=\lim_{k\rightarrow\infty}\norm{z^{k}-z^{*}}^{2}=2W_{\infty}(z^{*}).$
+This proves (d). $\blacksquare$
 […]
-We next consider a compact convex set $\mathcal{U}_{2}\subset\operatorname{dom}g_{1}\cap\operatorname{dom}g_{2}$ and choose as anchor point an arbitrary element $z\in\mathcal{U}_{2}$. Then, we have
-$\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle$ $\leq\norm{{\mathsf{F}}_{1}(z^{k+1/2})}\cdot\norm{z-z^{k+1/2}}$ (4.17)
-$\leq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+\norm{{\mathsf{F}}_{1}(z^{k+1/2})-{\mathsf{F}}_{1}(z^{1})}\right)\cdot\left(\norm{z^{1}-z^{k+1/2}}+\norm{z^{1}}+\norm{z}\right)$ (4.18)
-$\leq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+L_{{\mathsf{F}}_{1}}R\right)\cdot(R+\norm{z^{1}}+\max_{z\in\mathcal{U}_{2}}\norm{z}),$ (4.19)
+Consider a point $z\in\mathcal{D}_{1}\cap\mathcal{D}_{2}$. Invoking Assumption 1, we have
+$\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle\leq\norm{{\mathsf{F}}_{1}(z^{k+1/2})}\cdot\norm{z-z^{k+1/2}}$
+$\leq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+\norm{{\mathsf{F}}_{1}(z^{k+1/2})-{\mathsf{F}}_{1}(z^{1})}\right)\cdot\left(\norm{z^{1}-z^{k+1/2}}+\norm{z^{1}}+\norm{z}\right)$
+$\leq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+L_{{\mathsf{F}}_{1}}R\right)\cdot(R+\norm{z^{1}}+\norm{z})\triangleq a(z),$
 […]
-[…]
-$C_{\mathcal{U}_{2}}$ $\triangleq\left(\norm{{\mathsf{F}}_{1}(z^{1})}+L_{{\mathsf{F}}_{1}}R\right)(R+\norm{z^{1}}+\max_{z\in\mathcal{U}_{2}}\norm{z})+\Omega_{1}(\mathcal{U}_{2},R)$
-to obtain
-$\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2})\leq C_{\mathcal{U}_{2}},\qquad\forall z\in\mathcal{U}_{2},k\geq 1.$
-It follows by the definition of $\Psi_{k}(z)$ and monotonicity of ${\mathsf{F}}_{2}$ that, for all $z\in\mathcal{U}_{2}$ and $k\geq 1$,
-$\begin{split}\Psi_{k}(z)+\sigma_{k}C_{\mathcal{U}_{2}}&\geq\Psi_{k}(z)+\sigma_{k}(\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2}))\\ &=\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)\\ &\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z).\end{split}$ (4.20)
-Multiplying by $t_{k}$ and using (4.12), where we now neglect the non-positive term $\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}$, we obtain, for all $z\in\mathcal{U}_{2},k\geq 1$,
-$t_{k}(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z))$ $\leq t_{k}\Psi_{k}(z)+t_{k}\sigma_{k}C_{\mathcal{U}_{2}}$
-$\leq(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1})+t_{k}\sigma_{k}C_{\mathcal{U}_{2}}.$ (4.21)
-Summing (4.21) for $k=1,\ldots,K$ and dividing by $T_{K}$ yields, for all $z\in\mathcal{U}_{2}$,
-$\frac{1}{T_{K}}\sum_{k=1}^{K}t_{k}\left(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)\right)\leq\frac{1}{T_{K}}(E_{1}(z)+D_{1})+\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}C_{\mathcal{U}_{2}}.$ (4.22)
+$g_{1}(z)-g_{1}(z^{k+1/2})$ $\leq\operatorname{\mathsf{Var}}(g_{1}|\{z\}\times\mathcal{B}_{R})\triangleq b(z).$
+Thus, defining $C_{z}\triangleq a(z)+b(z)$ we can conclude that
+$H^{({\mathsf{F}}_{1},g_{1})}(z,z^{k+1/2})=\langle{\mathsf{F}}_{1}(z^{k+1/2}),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2})\leq C_{z}.$ (4.17)
+By (4.15) and monotonicity of ${\mathsf{F}}_{2}$ we have, for all $z\in\mathcal{D}_{1}\cap\mathcal{D}_{2}$ and $k\geq 1$,
+$\Psi_{k}(z)+\sigma_{k}C_{z}$ $\geq\Psi_{k}(z)+\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z,z^{k+1/2})$
+$=-H^{({\mathsf{F}}_{2},g_{2})}(z,z^{k+1/2})\geq H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z).$ (4.18)
+Multiplying by $t_{k}$, using (4.14) and (4.5), where we now neglect the non-positive quadratic terms, we obtain, for all $z\in\mathcal{D}_{1}\cap\mathcal{D}_{2}$ and $k\geq 1$,
+$t_{k}H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)$ $\leq t_{k}\Psi_{k}(z)+t_{k}\sigma_{k}C_{\mathcal{U}_{2}}\leq W_{k}(z)-W_{k+1}(z)+t_{k}\sigma_{k}C_{z}.$ (4.19)
+Summing (4.19) for $k=1,\ldots,K$ dividing by $T_{K}$, and applying $D_{1}=0$, yields
+$\sum_{k=1}^{K}\frac{t_{k}}{T_{K}}H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)\leq\frac{E_{1}(z)}{T_{K}}+\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}C_{z}\qquad\forall z\in\mathcal{D}_{1}\cap\mathcal{D}_{2}.$ (4.20)
 Using the definition of the ergodic trajectory (3.5), we can apply the Jensen inequality in order to finally arrive at
-$\langle{\mathsf{F}}_{2}(z),\bar{z}^{K}-z\rangle+g_{2}(\bar{z}^{K})-g_{2}(z)\leq\frac{1}{T_{K}}(E_{1}(z)+D_{1})+\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}C_{\mathcal{U}_{2}}\hskip 17.00024pt\forall z\in\mathcal{U}_{2}.$
-Thus, for any compact subset $\mathcal{U}_{2}\subset\operatorname{dom}g_{1}\cap\operatorname{dom}g_{2}\subset\mathcal{Z}$ we can use the definition of the localized gap function (2.4) specialized as the feasibility gap (2.6) together with $D_{1}=0$, to obtain (3.6).
-Moreover, under the step size condition (3.1) it follows that every weak accumulation point $\tilde{z}$ of the averaged trajectory $(\bar{z}^{k})$ satisfies
-$\langle{\mathsf{F}}_{2}(z),\tilde{z}-z\rangle+g_{2}(\tilde{z})-g_{2}(z)\leq 0\hskip 17.00024pt\forall z\in\mathcal{U}_{2}$
-In particular, $\Theta_{\rm Feas}(\tilde{z}|\mathcal{U}_{2})\leq 0$.
+$H^{({\mathsf{F}}_{2},g_{2})}(\bar{z}^{K},z)\leq\frac{E_{1}(z)}{T_{K}}+\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}C_{z}\hskip 17.00024pt\forall z\in\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2}).$ (4.21)
+Since $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is compact, we have that $b(z)\leq\operatorname{\mathsf{Var}}(g_{1}|\mathcal{U}_{2}\times\mathcal{B}_{R})$ and $C_{\mathcal{U}_{2}}\triangleq\sup_{z\in\mathcal{U}_{2}}C_{z}$, as well as $\sup_{z\in\mathcal{U}_{2}}E_{1}(z)$, are well-defined.
+Thus, for any compact subset $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})\subset\mathcal{Z}$ we can use the definition of the localized gap function (2.4) specialized as the feasibility gap (2.7), to obtain (3.6).
 Proof of (3.10)
-Choosing constant step $t_{k}=t$ and penalty sequence as in (3.9), we easily see that $\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}\leq\frac{a}{(1-\delta)(K+b)^{\delta}},$ so that the concrete complexity bound becomes
-$\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt}+\frac{aC_{\mathcal{U}_{2}}}{(1-\delta)(K+b)^{\delta}}.$
+Choosing constant step $t_{k}=t$ and penalty sequence as in (3.9), we easily see that $\frac{\sum_{k=1}^{K}t_{k}\sigma_{k}}{T_{K}}\leq\frac{a}{(1-\delta)(K+b)^{\delta}},$ so that (3.6) transforms into (3.10).
 Remark 4.1.
 Corollary 3.2 is formulated for regularization sequences $(\sigma_{k})_{k\geq 1}$ of the form (3.9) with $\delta\in(0,1)$.
-In the limiting case $\delta=1$ and constant step size $t_{k}=t$, the complexity bound
-$\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt}+\frac{a\log(K+b)}{K}C_{\mathcal{U}_{2}}.$
-Hence, even in this limiting case, not satisfying condition (3.1) explicitly requested in Algorithm 1, the ergodic average $(\bar{z}^{k})_{k}$ asymptotically approaches the feasible set $\mathcal{S}_{2}$.
+In the limiting case $\delta=1$ and constant step size $t_{k}=t$, the last term in (3.6) is upper bounded as $\frac{a\log(K+b)}{K}C_{\mathcal{U}_{2}}$.
+Hence, even in this limiting case, not satisfying condition (3.1) required in Algorithm 1, the ergodic average $(\bar{z}^{k})_{k\geq 1}$ asymptotically approaches the feasible set $\mathcal{S}_{2}$.
 $\Diamond$
 Establishing the rate on the optimality gap (3.7)
-Let $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be a compact set with $\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$.
-Using the monotonicity of ${\mathsf{F}}_{1}$ and ${\mathsf{F}}_{2}$, we see
-$\Psi_{k}(z)$ $=\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)$
-$+\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)$
-$\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)$
-$+\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right).$
-Hence, when choosing $z\in\mathcal{S}_{2}$, we are able to continue the above bound as
-$\Psi_{k}(z)$ $\geq\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)\qquad\forall k\geq 1.$
-Plugging this in (4.12) and dropping the non-positive term $\frac{1}{4}\norm{z^{k+1/2}-z^{k}}^{2}$, we arrive at
-$\sigma_{k}t_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)$ $\leq(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1})$
-$=W_{k}(z)-W_{k+1}(z).$
-To continue with our estimation, we proceed with the following simple result.
-Lemma 4.4.
-Let either Assumption 3 hold or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be compact. Then, the mapping $z\mapsto W_{k}(z)$ is bounded over bounded sets.
-Proof.
-Since
-$W_{k}(z)$ $=\frac{1}{2}\norm{z^{k}-z}^{2}+\frac{t_{k}^{2}}{2}\norm{{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2})}^{2}\leq\frac{1}{2}\norm{z^{k}-z}^{2}+\frac{t^{2}_{k}L^{2}_{k}}{2}\norm{z^{k+1/2}-z^{k-1/2}}^{2}$
-the result follows immediately from the boundedness of the sequences $(z^{k})_{k\geq 1}$, $(z^{k-1/2})_{k\geq 1}$ that hold both under Assumption 3 and when $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is compact.
-$\blacksquare$
-Since $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$, it follows from the assumptions made that $\mathcal{U}_{1}\cap\mathcal{S}_{2}\neq\varnothing$.
-Therefore, we can pick $z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}$ arbitrary, so that we can find a positive constant $\bar{\omega}(z)$ such that $W_{k}(z)\leq\bar{\omega}(z)$ for all $k\geq 1$.
-Using this bound, and $\sigma_{k+1}\leq\sigma_{k}$, we arrive that
-$\sum_{k=1}^{K}t_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)$ $\leq\frac{1}{\sigma_{1}}W_{1}(z)+W_{2}(z)(\frac{1}{\sigma_{2}}-\frac{1}{\sigma_{1}})+\ldots+W_{K}(z)(\frac{1}{\sigma_{K}}-\frac{1}{\sigma_{K-1}})$
+Let $\mathcal{U}_{1}\subset\mathcal{D}_{1}\cap\mathcal{D}_{2}$ be a compact set s.t.
+$\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$.
+By the monotonicity of ${\mathsf{F}}_{1}$ and ${\mathsf{F}}_{2}$ and choosing $z\in\mathcal{S}_{2}$, we see
+[…]
+$\geq\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)+H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)\geq\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z).$
+Plugging this in (4.5), dropping the non-positive quadratic terms, and recalling (4.14), we arrive at $\sigma_{k}t_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)\leq W_{k}(z)-W_{k+1}(z)$ for all $z\in\mathcal{S}_{2}$.
+Since by the $L_{k-1}$-Lipschitz property of ${\mathsf{V}}_{k-1}$,
+$W_{k}(z)$ $\leq\frac{1}{2}\norm{z^{k}-z}^{2}+t^{2}_{k-1}L^{2}_{k-1}\norm{z^{k-3/2}-z^{k-1/2}}^{2},$
+we immediately see that, for any fixed $z\in\mathcal{S}_{2}$, $W_{k}(z)$ is bounded in the sense that for some positive constant $\bar{\omega}(z)$, $W_{k}(z)\leq\bar{\omega}(z)$ for all $k\geq 1$.
+Using this bound, $W_{k}(z)\geq 0$, and $\sigma_{k+1}\leq\sigma_{k}$, we obtain that, for any $z\in\mathcal{S}_{2}$,
+$\sum_{k=1}^{K}t_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)$ $\leq\frac{1}{\sigma_{1}}W_{1}(z)+W_{2}(z)(\frac{1}{\sigma_{2}}-\frac{1}{\sigma_{1}})+\ldots+W_{K}(z)(\frac{1}{\sigma_{K}}-\frac{1}{\sigma_{K-1}})$
 $\leq\frac{\bar{\omega}(z)}{\sigma_{K}}.$
-Dividing both sides by $T_{K}$ and applying the Jensen inequality, we obtain
-$\langle{\mathsf{F}}_{1}(z),\bar{z}^{K}-z\rangle+g_{1}(\bar{z}^{K})-g_{1}(z)\leq\frac{\bar{\omega}(z)}{T_{K}\sigma_{K}}.$
-Thus, for any compact subset $\mathcal{U}_{1}\subset\mathcal{Z}$ with $\mathcal{U}_{1}\cap\mathcal{S}_{2}\neq\varnothing$, we can use the definition of the localized gap function (2.4) specialized as the optimality gap (2.7) to obtain
-$\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{T_{K}\sigma_{K}},$ (4.23)
-where $C_{\mathcal{U}_{1}}\triangleq\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\bar{\omega}(z)$.
-This is the upper bound in (3.7).
+Dividing both sides by $T_{K}$ and applying the Jensen inequality, we arrive at
+$H^{({\mathsf{F}}_{1},g_{1})}(\bar{z}^{K},z)\leq\frac{\bar{\omega}(z)}{T_{K}\sigma_{K}}\qquad\forall z\in\mathcal{S}_{2}.$ (4.22)
+Thus, for any compact subset $\mathcal{U}_{1}\subset\mathcal{Z}$ with $\mathcal{U}_{1}\cap\mathcal{S}_{2}\neq\varnothing$, we can use the definition of the localized gap function (2.4) specialized as the optimality gap (2.8) to obtain $\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{T_{K}\sigma_{K}}$, where $C_{\mathcal{U}_{1}}\triangleq\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\bar{\omega}(z)$.
+Note that the latter is well-defined since $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$ and $\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$ by the Theorem assumption.
+Thus, we have obtained the upper bound in (3.7).
 The lower bound in (3.7) holds since $\mathcal{S}_{1}\subseteq\mathcal{S}_{2}$:
@@
 Establishing the improved rate on the optimality gap (3.8)
-If additionally the lower level solution set enjoys $(\alpha,\rho)$ weak-sharpness we have from (2.9) and (3.6)
+If additionally the lower level solution set enjoys $(\alpha,\rho)$-weak sharpness we have from (2.10) and (3.6)
 $\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2}$ $)\geq-B_{\mathcal{U}_{1}}\operatorname{dist}(\bar{z}^{K},\mathcal{S}_{2})$
-$\geq-B_{\mathcal{U}_{1}}\left[\frac{\rho}{\alpha}\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\cap\mathcal{S}_{2})\right]^{1/\rho}\geq-B_{\mathcal{U}_{1}}\left[\frac{\rho}{\alpha}\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\right]^{1/\rho}$
+$\geq-B_{\mathcal{U}_{1}}\left[\frac{\rho}{\alpha}\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2}\cap\mathcal{S}_{2})\right]^{1/\rho}\geq-B_{\mathcal{U}_{1}}\left[\frac{\rho}{\alpha}\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\right]^{1/\rho}$
 $\geq-B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2T_{K}(\alpha/\rho)}+\frac{C_{\mathcal{U}_{2}}\sum_{k=1}^{K}t_{k}\sigma_{k}}{(\alpha/\rho)T_{K}}\right]^{1/\rho}.$
-This verifies (3.8).
 Proof of (3.11)
-Choosing the penalty sequence $(\sigma_{k})_{k\geq 1}$ according to (3.9) and the step size $t_{k}=t$ small enough, the lower and upper complexity bound on the gap function becomes
-$B_{\mathcal{U}_{1}}\left[\frac{\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2Kt(\alpha/\rho)}+\frac{aC_{\mathcal{U}_{2}}}{(\alpha/\rho)(1-\delta)(K+b)^{\delta}}\right]^{1/\rho}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{C_{\mathcal{U}_{1}}}{at(K+b)^{1-\delta}}.$
-Note that if $\delta=1$ in the choice of the penalty sequence (3.9), we do not obtain convergence in terms of the gap function to 0, but rather only an $O(1)$ upper bound and $o(1)$ lower bound.
+Choosing the penalty sequence $(\sigma_{k})_{k\geq 1}$ according to (3.9) and the step size $t_{k}=t$, the r.h.s.
+of (3.7) is upper bounded by the r.h.s.
+in (3.11) and the l.h.s.
+in (3.11) follows from (3.10).
+Note that if $\delta=1$ in the choice of the penalty sequence (3.9), we do not obtain convergence in terms of the gap function to 0, but rather only an $O(1)$ upper bound and a $o(1)$ lower bound.
+Proof of Theorem 3.3
+(i) From the prox-inequality (2.1) applied to the point $z^{k+1/2}$ and divided by $t_{k}>0$, we have, for any $z\in\mathcal{Z}$,
+$t_{k}^{-1}\langle z^{k+1/2}-z^{k},z^{k+1/2}-z\rangle$ $\leq\langle{\mathsf{V}}_{k}(z^{k-1/2}),z-z^{k+1/2}\rangle+G_{k}(z)-G_{k}(z^{k+1/2})$
+[…]
+$\hskip-17.00024pt\leq\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z,z^{k+1/2})-H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)+e_{k},$ (4.23)
+[…]
+$\lim_{j\to\infty}H^{({\mathsf{F}}_{1},g_{1})}(z,z^{k_{j}+1/2})=\lim_{j\to\infty}\langle{\mathsf{F}}_{1}(z^{k_{j}+1/2}),z-z^{k_{j}+1/2}\rangle+g_{1}(z)-g_{1}(z^{k_{j}+1/2})$
+[…]
+Thus, since $\sigma_{k}\to 0$ as $k\to\infty$, we have for any $z\in\operatorname{dom}(g_{2})\subseteq\operatorname{dom}(g_{1})$, that $\lim_{j\to\infty}\sigma_{k_{j}}H^{({\mathsf{F}}_{1},g_{1})}(z,z^{k_{j}+1/2})\leq 0.$ Thus, we obtained the limits of all the terms in (4.23) except the term with $H^{({\mathsf{F}}_{2},g_{2})}$.
+Using these limits when passing in (4.23) to the subsequence indexed by $k_{j}$ and then to the limit $j\to\infty$, using the lower semicontinuity of $g_{2}$, we obtain, for any $z\in\operatorname{dom}(g_{2})$
+$0$ $\leq\lim_{j\to\infty}-H^{({\mathsf{F}}_{2},g_{2})}(z^{k_{j}+1/2},z)=\lim_{j\to\infty}\langle{\mathsf{F}}_{2}(z),z-z^{k_{j}+1/2}\rangle+g_{2}(z)-g_{2}(z^{k_{j}+1/2})$
+[…]
+For $z\notin\operatorname{dom}(g_{2})$ the above inequality holds trivially.
+Thus, it holds for all $z\in\mathcal{Z}$.
+This shows that any weak limit point of $(z^{k+1/2})_{k\geq 1}$ is a Minty solution of $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$.
+Under Assumption 1, Minty solutions and Stampacchia solutions of $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$ (2.3) coincide (see also the proof of Lemma 2.1 in Appendix C), showing that $\tilde{z}\in\mathcal{S}_{2}$.
+(ii) By Lemma 4.2 (c), the sequence $(z^{k+1/2})_{k\geq 1}$ is bounded.
+Hence the ergodic average sequence $(\bar{z}^{k})_{k\geq 1}$ is bounded as well and thus admits weakly converging subsequences.
+Let $\tilde{z}$ be such that $\bar{z}^{k_{j}}\rightharpoonup\tilde{z}$ as $j\to\infty$.
+For any reference point $z\in\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})=\operatorname{dom}(g_{2})$ applying condition (3.1) to (4.21), we obtain that $\lim_{j\rightarrow\infty}H^{({\mathsf{F}}_{2},g_{2})}(\bar{z}^{k_{j}},z)\leq 0$.
+Thus, using the lower semi-continuity of $g_{2}$, $\tilde{z}$ satisfies $H^{({\mathsf{F}}_{2},g_{2})}(\tilde{z},z)\leq 0$.
+This inequality trivially holds for $z\notin\operatorname{dom}(g_{2})$, and hence holds for all $z\in\mathcal{Z}$.
+Thus, $\tilde{z}$ is a Minty solution of $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$, which again, under Assumption 1 is also a Stampacchia solution of $\operatorname{HVI}({\mathsf{F}}_{2},g_{2})$.
+This implies $\tilde{z}\in\mathcal{S}_{2}$.
+For any point $z\in\mathcal{S}_{2}$, (4.22) and (3.1) yield that $\lim_{j\rightarrow\infty}H^{({\mathsf{F}}_{1},g_{1})}(\bar{z}^{k_{j}},z)\leq 0.$ This, by the lower semi-continuity of $g_{1}$, implies
+$\langle{\mathsf{F}}_{1}(z),\tilde{z}-z\rangle+$ $g_{1}(\tilde{z})-g_{1}(z)\leq\lim_{j\to\infty}\left(\langle{\mathsf{F}}_{1}(z),\bar{z}^{k_{j}}-z\rangle+g_{1}(\bar{z}^{k_{j}})-g_{1}(z)\right)\leq 0.$
+Since $z\in\mathcal{S}_{2}$ is arbitrary and $\tilde{z}\in\mathcal{S}_{2}$, we have that $\tilde{z}$ is a Minty solution of (P), which is, under Assumption 1, also a Stampacchia solution of $\operatorname{HVI}({\mathsf{F}}_{1},g_{1})$, i.e., $\tilde{z}\in\mathcal{S}_{1}$.
+Lemma 4.2 (d) shows that the limit of $(\norm{z^{k+1/2}-z})_{k\geq 1}$ exists for all $z\in\mathcal{S}_{1}$.
+As just demonstrated, all weak accumulation points of $(\bar{z}^{k})_{k\geq 1}$ are points in $\mathcal{S}_{1}$.
+Lemma A.2 thus shows that $\bar{z}^{k}\rightharpoonup z_{\infty}$ for some $z_{\infty}\in\mathcal{S}_{1}$.
 5 Extensions
-This section contains several extensions.
-First, we develop estimates for the case in which the upper level $\operatorname{HVI}$ is strongly monotone, implying that the entire problem (P) has a unique solution.
-We then move on by explaining how Algorithm 1 can be extended to the Bregman setting, and Tseng’s modified extragradient method.
-5.1 Improved rates under stronger monotonicity assumptions
-On top the already formulated hypothesis, we impose the following:
+In this section, first we develop estimates for the case in which the upper-level $\operatorname{HVI}$ is strongly monotone, implying that the entire problem (P) has a unique solution.
+Second, we show how our method can be extended in a straightforward manner to the forward-backward-forward splitting method of [62].
+5.1 Improved rates under strong monotonicity assumption
+On top of the already formulated hypothesis, we impose the following:
 Assumption 6.
 The operator ${\mathsf{F}}_{1}:\mathcal{Z}\to\mathcal{Z}$ is $\mu$-strongly monotone.
-Since we assume that the over hierarchical equilibrium problem has a nonempty feasible set, Assumption 4 implies that $\mathcal{S}_{1}=\{z^{*}\}$ for some $z^{*}\in\mathcal{S}_{2}$.
+Since we assume that the whole hierarchical equilibrium problem (P) has a nonempty feasible set, Assumption 6 implies that $\mathcal{S}_{1}=\{z^{*}\}$ for some $z^{*}\in\mathcal{S}_{2}$.
 We again establish rates on gap functions in terms of the suitably adapted ergodic average
-$\bar{z}^{K}=\frac{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}z^{i+1/2}}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}},$ (5.1)
-where $\gamma_{k}\triangleq\frac{1}{\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)}$ and $\gamma_{0}\triangleq 1$.
+$\bar{z}^{K}=\frac{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}z^{i+1/2}}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}},\text{ where }\gamma_{k}\triangleq\frac{1}{\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)},\gamma_{0}\triangleq 1.$ (5.1)
 Theorem 5.1.
 Consider problem (P).
-Let Assumptions 1, 2, 4 be in place.
-Let additionally either Assumption 3 hold or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be compact.
-Let $(\sigma_{k})_{k\geq 0}$ be non-increasing and the sequences $(t_{k})_{k\geq 0}$ and $(\sigma_{k})_{k\geq 0}$ satisfy
-$4t^{2}_{k}L^{2}_{k}+2t_{k}\sigma_{k}\mu\leq 1,\qquad k\geq 1.$ (5.2)
+Let Assumptions 1, 2, 4, 6 hold.
+Let additionally either Assumption 3 hold, or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be compact.
+Let $(\bar{z}^{k})_{k\geq 1}$ as in (5.1) be generated by Algorithm 1 with
+$8t^{2}_{k}L^{2}_{k}+2t_{k}\sigma_{k}\mu\leq 1,\qquad k\geq 1.$ (5.2)
 Let $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ be a nonempty compact set. Then, there exists a constant $\tilde{C}_{\mathcal{U}_{2}}>0$ for which
@@
 Additionally, for any nonempty compact subset $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ with $\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$, there exists a constant $B_{\mathcal{U}_{1}}>0$ such that
-$B_{\mathcal{U}_{1}}\operatorname{dist}(\bar{z}^{K},\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{2\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$ (5.4)
-In particular, if the lower level solution set $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp, then
-$B_{\mathcal{U}_{1}}\left[\frac{\sigma_{1}\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{2(\alpha/\rho)\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}+\frac{\tilde{C}_{\mathcal{U}_{1}\cap\mathcal{S}_{2}}\sum_{i=1}^{k}t_{i}\sigma^{2}_{i}\gamma_{i}}{(\alpha/\rho)\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}\right]^{1/\rho}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{2\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$ (5.5)
+$B_{\mathcal{U}_{1}}\operatorname{dist}(\bar{z}^{K},\mathcal{S}_{2})\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{2\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$ (5.4)
+If additionally the lower level solution set $\mathcal{S}_{2}$ is $(\alpha,\rho)$-weakly sharp and $\mathcal{U}_{2}$ is s.t.
+$\mathcal{U}_{2}\cap\mathcal{S}_{2}\neq\varnothing$, then for the same set $\mathcal{U}_{1}$
+$\begin{split}-B_{\mathcal{U}_{1}}&\left[\frac{\sigma_{1}\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{2(\alpha/\rho)\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}+\frac{\tilde{C}_{\mathcal{U}_{2}}\sum_{i=1}^{k}t_{i}\sigma^{2}_{i}\gamma_{i}}{(\alpha/\rho)\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}\right]^{\frac{1}{\rho}}\leq\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2}).\end{split}$ (5.5)
 Proof.
-We start with a refined version of the recursion (4.12). Recall the notation (4.9) and the bound (4.1) that holds for any $z\in\mathcal{Z}$ and $k\geq 1$:
-$E_{k+1}(z)$ $\leq E_{k}(z)-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
+We start with a refined version of the recursion (4.5). Recall the bound (4.4) that implies for any $z\in\mathcal{Z}$ and $k\geq 1$:
+$E_{k+1}(z)$ $\leq E_{k}(z)-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+D_{k+1}$
 $+t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+t_{k}\left(G_{k}(z)-G_{k}(z^{k+1/2})\right).$ (5.6)
 For the second and third terms in the r.h.s. we have from (4.11)
-$\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq 2t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$
-$+\frac{1}{2}\left(4t_{k}^{2}L^{2}_{k}-1\right)\norm{z^{k+1/2}-z^{k}}^{2}.$ (5.7)
-Using the definition of $z^{k}$ and $z^{k-1/2}$, the non-expansiveness of the proximal mapping (LABEL:eq:nonexpansive), and the step size assumption (5.2), we further obtain
-$2t^{2}_{k}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}$ $\leq 2t^{2}_{k}t_{k-1}^{2}L_{k}^{2}\norm{{\mathsf{V}}_{k-1}(z^{k-1/2})-{\mathsf{V}}_{k-1}(z^{k-3/2})}^{2}\leq(1-t_{k}\sigma_{k}\mu)\frac{t_{k-1}^{2}}{2}\norm{{\mathsf{V}}_{k-1}(z^{k-1/2})-{\mathsf{V}}_{k-1}(z^{k-3/2})}^{2}.$ (5.8)
-Combining (5.7), (5.8), and using the notation (4.11), we obtain
-$\frac{t^{2}_{k}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq(1-t_{k}\sigma_{k}\mu)D_{k}-D_{k+1}+\frac{1}{2}\left(4t_{k}^{2}L^{2}_{k}-1\right)\norm{z^{k+1/2}-z^{k}}^{2}.$ (5.9)
-By the strong monotonicity, we obtain
+$D_{k+1}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq 4t_{k}^{2}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}-D_{k+1}$
+$+\left(4t_{k}^{2}L^{2}_{k}-\frac{1}{2}\right)\norm{z^{k+1/2}-z^{k}}^{2}.$ (5.7)
+Acting in the same way as for the proof of (4.13), but with the new step size assumption (5.2) that implies $4t_{k}^{2}L_{k}^{2}\leq 1-t_{k}\sigma_{k}\mu$, we further obtain
+$4t^{2}_{k}L^{2}_{k}\norm{z^{k}-z^{k-1/2}}^{2}$ $\leq 4t^{2}_{k}L_{k}^{2}t_{k-1}^{2}\norm{{\mathsf{V}}_{k-1}(z^{k-1/2})-{\mathsf{V}}_{k-1}(z^{k-3/2})}^{2}$
+$\leq(1-t_{k}\sigma_{k}\mu)D_{k}.$ (5.8)
+Combining (5.7), (5.8), we obtain
+$D_{k+1}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}$ $\leq(1-t_{k}\sigma_{k}\mu)D_{k}-D_{k+1}+\left(4t_{k}^{2}L^{2}_{k}-\frac{1}{2}\right)\norm{z^{k+1/2}-z^{k}}^{2}.$ (5.9)
+By the strong monotonicity of ${\mathsf{F}}_{1}$, we obtain
 $\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle$ $=\langle{\mathsf{F}}_{2}(z^{k+1/2}),z^{k+1/2}-z\rangle+\sigma_{k}\langle{\mathsf{F}}_{1}(z^{k+1/2}),z^{k+1/2}-z\rangle$
-$\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+\sigma_{k}\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+\sigma_{k}\mu\norm{z^{k+1/2}-z}^{2}$
-$\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+\sigma_{k}\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+\frac{\sigma_{k}\mu}{2}\norm{z^{k}-z}^{2}-\sigma_{k}\mu\norm{z^{k+1/2}-z^{k}}^{2},$ (5.10)
+$\hskip-42.50061pt\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+\sigma_{k}\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+\sigma_{k}\mu\norm{z^{k+1/2}-z}^{2}$
+$\hskip-42.50061pt\geq\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+\sigma_{k}\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle$
+$\hskip-34.00049pt+\frac{\sigma_{k}\mu}{2}\norm{z^{k}-z}^{2}-\sigma_{k}\mu\norm{z^{k+1/2}-z^{k}}^{2},$ (5.10)
 where the last inequality follows from the triangle inequality and $(a+b)^{2}/2\leq a^{2}+b^{2}\Leftrightarrow a^{2}\geq(a+b)^{2}/2-b^{2}$.
-Plugging (5.9) and (5.10) multiplied by 1-1 into (5.6), recalling the notation (4.9) and $G_{k}(\cdot)=g_{2}(\cdot)+\sigma_{k}g_{1}(\cdot)$, we obtain
-$E_{k+1}(z)$ $+D_{k+1}\leq(1-t_{k}\sigma_{k}\mu)(E_{k}(z)+D_{k})+(2t^{2}_{k}L^{2}_{k}+t_{k}\mu\sigma_{k}-\frac{1}{2})\norm{z^{k+1/2}-z^{k}}^{2}$
-$-t_{k}\left(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)+\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)\right)$
+Plugging (5.9) and (5.10) multiplied by 1-1 into (5.6), recalling $G_{k}=g_{2}+\sigma_{k}g_{1}$, we obtain
+$E_{k+1}(z)$ $+D_{k+1}$
+$\leq(1-t_{k}\sigma_{k}\mu)(E_{k}(z)+D_{k})+(4t^{2}_{k}L^{2}_{k}+t_{k}\mu\sigma_{k}-\frac{1}{2})\norm{z^{k+1/2}-z^{k}}^{2}$
+$-t_{k}(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z))$
+$-t_{k}\sigma_{k}(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z))$
 $\leq(1-t_{k}\sigma_{k}\mu)(E_{k}(z)+D_{k})$
-$-t_{k}\left(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z)+\sigma_{k}\left(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z)\right)\right)$ (5.11)
-where the last inequality uses (5.2).
-We now use this recursion to establish rates on the feasibility and optimality gap, starting with the feasibility gap.
-Since the analysis for the monotone setting holds also for strongly monotone case, we have that either Lemma 4.3 still holds under the same assumptions made there (i.e.
-either Assumption 3 or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ compact).
-Thus, in any case, for a compact subset $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$, there exists a constant $\tilde{C}_{\mathcal{U}_{2}}$ such that
-$\langle{\mathsf{F}}_{1}(z),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2})\leq\tilde{C}_{\mathcal{U}_{2}},\qquad\forall z\in\mathcal{U}_{2},k\geq 1.$
-Combining this with (5.11), we get
-$t_{k}(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z))$ $\leq\left(1-t_{k}\sigma_{k}\mu\right)(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1})+t_{k}\sigma_{k}\widetilde{C}_{\mathcal{U}_{2}}.$
-Multiplying both sides by $\sigma_{k}\gamma_{k}$ (see (5.1) for the definition of $\gamma_{k}$) and using that $\sigma_{k}\geq\sigma_{k+1}$, we obtain
-$t_{k}\sigma_{k}\gamma_{k}(\langle{\mathsf{F}}_{2}(z),z^{k+1/2}-z\rangle+g_{2}(z^{k+1/2})-g_{2}(z))$ $\leq\sigma_{k}\gamma_{k-1}(E_{k}(z)+D_{k})-\sigma_{k+1}\gamma_{k}(E_{k+1}(z)+D_{k+1})+t_{k}\sigma_{k}^{2}\gamma_{k}\widetilde{C}_{\mathcal{U}_{2}}.$
-Thus, telescoping these inequalities and using the convexity of $g_{2}$, for the ergodic average (5.1), we obtain
-$\langle{\mathsf{F}}_{2}(z),\bar{z}^{K}-z\rangle+g_{2}(\bar{z}^{K})-g_{2}(z)$ $\leq\frac{\sigma_{1}(E_{1}(z)+D_{1})+\widetilde{C}_{\mathcal{U}_{2}}\sum_{i=1}^{K}t_{i}\sigma_{i}^{2}\gamma_{i}}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$
-Taking the supremum over $\mathcal{U}_{2}$ on both sides and using the definition of the localized gap function (2.4) specialized as the feasibility gap (2.6) together with $D_{1}=0$, we obtain the upper bound (5.3).
+$-t_{k}\left(H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)+\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)\right),$ (5.11)
+where the last inequality uses (5.2) and (2.5). We now use this recursion to establish rates on the feasibility and optimality gap, starting with the feasibility gap.
+Since the analysis for the monotone setting holds also for strongly monotone case, we have that either Assumption 3 holds and implies by Lemma 4.2 that the sequences generated by the algorithm are bounded or $\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is compact.
+Thus, in any case, for a compact subset $\mathcal{U}_{2}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$, there exists a constant $\tilde{C}_{\mathcal{U}_{2}}$ such that, for all $z\in\mathcal{U}_{2},k\geq 1$, (cf.
+the derivation of (4.17))
+$H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)=\langle{\mathsf{F}}_{1}(z),z-z^{k+1/2}\rangle+g_{1}(z)-g_{1}(z^{k+1/2})\leq\tilde{C}_{\mathcal{U}_{2}}.$
+Combining this with (5.11) and recalling the notation (4.14), we get
+$t_{k}H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)$ $\leq\left(1-t_{k}\sigma_{k}\mu\right)W_{k}(z)-W_{k+1}(z)+t_{k}\sigma_{k}\widetilde{C}_{\mathcal{U}_{2}}.$
+Multiplying both sides by $\sigma_{k}\gamma_{k}$ (see (5.1) for the definition of $\gamma_{k}$) and using that $\sigma_{k}\geq\sigma_{k+1}$ and $W_{k}(z)\geq 0$ for $k\geq 1$, we obtain
+$t_{k}\sigma_{k}\gamma_{k}H^{({\mathsf{F}}_{2},g_{2})}(z^{k+1/2},z)$ $\leq\sigma_{k}\gamma_{k-1}W_{k}(z)-\sigma_{k+1}\gamma_{k}W_{k+1}(z)+t_{k}\sigma_{k}^{2}\gamma_{k}\widetilde{C}_{\mathcal{U}_{2}}.$
+Telescoping these inequalities and using the convexity of $g_{2}$ and $D_{1}=0$, $\gamma_{0}=1$, for the ergodic average (5.1), we obtain
+$H^{({\mathsf{F}}_{2},g_{2})}(\bar{z}^{K},z)$ $\leq\frac{\sigma_{1}E_{1}(z)+\widetilde{C}_{\mathcal{U}_{2}}\sum_{i=1}^{K}t_{i}\sigma_{i}^{2}\gamma_{i}}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$
+Taking the supremum over $z\in\mathcal{U}_{2}$ on both sides and using the definition of the feasibility gap (2.7), we obtain the upper bound (5.3).
 To obtain a bound in terms of the optimality gap, consider an arbitrary reference point $z\in\mathcal{S}_{2}$ and rearrange (5.11) to arrive at
-$t_{k}\sigma_{k}(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z))$ $\leq\left(1-t_{k}\sigma_{k}\mu\right)(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1}).$
+$t_{k}\sigma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)$ $\leq\left(1-t_{k}\sigma_{k}\mu\right)(E_{k}(z)+D_{k})-(E_{k+1}(z)+D_{k+1}).$
 Multiplying the previous inequality by $\gamma_{k}$ yields
-$t_{k}\sigma_{k}\gamma_{k}(\langle{\mathsf{F}}_{1}(z),z^{k+1/2}-z\rangle+g_{1}(z^{k+1/2})-g_{1}(z))$ $\leq\gamma_{k-1}(E_{k}(z)+D_{k})-\gamma_{k}(E_{k+1}(z)+D_{k+1}).$
+$t_{k}\sigma_{k}\gamma_{k}H^{({\mathsf{F}}_{1},g_{1})}(z^{k+1/2},z)$ $\leq\gamma_{k-1}(E_{k}(z)+D_{k})-\gamma_{k}(E_{k+1}(z)+D_{k+1}).$
 Telescoping these inequalities, using the convexity of $g_{1}$, $D_{1}=0$, $\gamma_{0}=1$, for the ergodic average (5.1), we obtain
-$\langle{\mathsf{F}}_{1}(z),\bar{z}^{K}-z\rangle+g_{1}(\bar{z}^{K})-g_{1}(z)$ $\leq\frac{E_{1}(z)}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$
+$H^{({\mathsf{F}}_{1},g_{1})}(\bar{z}^{K},z)$ $\leq\frac{E_{1}(z)}{\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}}.$
 Since $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})\cap\operatorname{dom}(g_{2})$ is a nonempty compact set with $\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$, we have that $\mathcal{U}_{1}\cap\mathcal{S}_{2}\neq\varnothing$.
@@
 The lower bound in (5.4) is the same as in (3.7).
-If additionally the lower level solution set enjoys $(\alpha,\rho)$ weak-sharpness we obtain (5.5) by combining (5.4), (2.9), and (5.3). $\blacksquare$
+If additionally the lower level solution set enjoys $(\alpha,\rho)$-weak sharpness we obtain (5.5) by combining (5.4), (2.10), and (5.3). $\blacksquare$
 To have a better understanding of the obtained rates, we now make a particular choice of the sequences $(t_{k})_{k\geq 1}$, $(\sigma_{k})_{k\geq 1}$.
@@
 $\Theta_{\rm Feas}(\bar{z}^{K}|\mathcal{U}_{2})\leq\frac{8L_{{\mathsf{F}}_{2}}(L_{{\mathsf{F}}_{1}}+\mu)\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{\mu K}+\tilde{C}_{\mathcal{U}_{2}}\frac{4L_{{\mathsf{F}}_{2}}(1+\ln K)}{\mu K},$ (5.13)
-$-B_{\mathcal{U}_{1}}\left[\frac{8L_{{\mathsf{F}}_{2}}(L_{{\mathsf{F}}_{1}}+\mu)\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{\mu K(\alpha/\rho)}+\tilde{C}_{\mathcal{U}_{2}}\frac{4L_{{\mathsf{F}}_{2}}(1+\ln K)}{\mu K(\alpha/\rho)}\right]^{1/\rho}\overset{(*)}{\leq}\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})$ (5.14)
-$\leq\frac{2(L_{{\mathsf{F}}_{1}}+\mu)\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{K},$ (5.15)
-where the inequality ()(*) holds under the $(\alpha,\rho)$-weakly sharpness assumption.
+$-B_{\mathcal{U}_{1}}$ $\left[\frac{8L_{{\mathsf{F}}_{2}}(L_{{\mathsf{F}}_{1}}+\mu)\sup_{z\in\mathcal{U}_{2}}\norm{z^{1}-z}^{2}}{\mu K(\alpha/\rho)}+\tilde{C}_{\mathcal{U}_{2}}\frac{4L_{{\mathsf{F}}_{2}}(1+\ln K)}{\mu K(\alpha/\rho)}\right]^{1/\rho}$ (5.14)
+$\overset{(*)}{\leq}\Theta_{\rm Opt}(\bar{z}^{K}|\mathcal{U}_{1}\cap\mathcal{S}_{2})\leq\frac{2(L_{{\mathsf{F}}_{1}}+\mu)\sup_{z\in\mathcal{U}_{1}\cap\mathcal{S}_{2}}\norm{z^{1}-z}^{2}}{K},$
+where the inequality ()(*) holds under the $(\alpha,\rho)$-weak sharpness assumption.
 Proof.
-First, we show that (5.2) is indeed satisfied:
-$4t_{k}^{2}L_{k}^{2}+2t_{k}\sigma_{k}\mu=4\frac{L_{k}^{2}}{16(L_{k}+\sigma_{k}\mu)^{2}}+\frac{2\sigma_{k}\mu}{4(L_{k}+\sigma_{k}\mu)}\leq\frac{1}{4}+\frac{1}{2}\leq 1.$ (5.16)
-Further,
-$t_{i}\sigma_{i}=\frac{1}{4(L_{{\mathsf{F}}_{2}}+\sigma_{i}(L_{{\mathsf{F}}_{1}}+\mu))}\cdot\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}=\frac{1}{L_{{\mathsf{F}}_{2}}+\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}(L_{{\mathsf{F}}_{1}}+\mu)}\cdot\frac{L_{{\mathsf{F}}_{2}}}{\mu i}=\frac{1}{i+\frac{4}{\mu}(L_{{\mathsf{F}}_{1}}+\mu)}\cdot\frac{1}{\mu}=\frac{1}{i+\varkappa}\cdot\frac{1}{\mu},$ (5.17)
-where we denoted $\varkappa=\frac{4}{\mu}(L_{{\mathsf{F}}_{1}}+\mu)$. This gives us
-$\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)=\prod_{i=1}^{k}(1-\frac{1}{i+\varkappa}\cdot\frac{1}{\mu}\cdot\mu)=\prod_{i=1}^{k}(1-\frac{1}{i+\varkappa})=\prod_{i=1}^{k}(\frac{i+\varkappa-1}{i+\varkappa})=\frac{\varkappa}{k+\varkappa}$ (5.18)
-and $\gamma_{k}=1/\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)=\frac{k+\varkappa}{\varkappa}$. Finally, we have
-$\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}=\sum_{i=1}^{K}\frac{1}{i+\varkappa}\cdot\frac{1}{\mu}\cdot\frac{i+\varkappa}{\varkappa}=\frac{K}{\mu\varkappa}=\frac{K}{4(L_{{\mathsf{F}}_{1}}+\mu)},$ (5.19)
-$\sum_{i=1}^{K}t_{i}\sigma_{i}^{2}\gamma_{i}=\sum_{i=1}^{K}\frac{1}{i+\varkappa}\cdot\frac{1}{\mu}\cdot\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}\cdot\frac{i+\varkappa}{\varkappa}=\frac{L_{{\mathsf{F}}_{2}}}{\mu(L_{{\mathsf{F}}_{1}}+\mu)}\sum_{i=1}^{K}\frac{1}{i}\leq\frac{L_{{\mathsf{F}}_{2}}(1+\ln K)}{\mu(L_{{\mathsf{F}}_{1}}+\mu)}.$ (5.20)
+First, it is easy to check that (5.2) is indeed satisfied.
+Further, $t_{i}\sigma_{i}=\frac{1}{L_{{\mathsf{F}}_{2}}+\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}(L_{{\mathsf{F}}_{1}}+\mu)}\cdot\frac{L_{{\mathsf{F}}_{2}}}{\mu i}=\frac{1}{i+\kappa}\cdot\frac{1}{\mu}$, where we denoted $\kappa=\frac{4}{\mu}(L_{{\mathsf{F}}_{1}}+\mu)$.
+This gives us
+$\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)=\prod_{i=1}^{k}\left(\frac{i+\kappa-1}{i+\kappa}\right)=\frac{\kappa}{k+\kappa}$
+and thus $\gamma_{k}=1/\prod_{i=1}^{k}(1-t_{i}\sigma_{i}\mu)=\frac{k+\kappa}{\kappa}$. Finally, we have
+$\sum_{i=1}^{K}t_{i}\sigma_{i}\gamma_{i}$ $=\sum_{i=1}^{K}\frac{1}{i+\kappa}\cdot\frac{1}{\mu}\cdot\frac{i+\kappa}{\kappa}=\frac{K}{\mu\kappa}=\frac{K}{4(L_{{\mathsf{F}}_{1}}+\mu)},$
+$\sum_{i=1}^{K}t_{i}\sigma_{i}^{2}\gamma_{i}$ $=\sum_{i=1}^{K}\frac{1}{i+\kappa}\cdot\frac{1}{\mu}\cdot\frac{4L_{{\mathsf{F}}_{2}}}{\mu i}\cdot\frac{i+\kappa}{\kappa}=\frac{L_{{\mathsf{F}}_{2}}\sum_{i=1}^{K}\frac{1}{i}}{\mu(L_{{\mathsf{F}}_{1}}+\mu)}\leq\frac{L_{{\mathsf{F}}_{2}}(1+\ln K)}{\mu(L_{{\mathsf{F}}_{1}}+\mu)}.$
 Substituting this into (5.3), we obtain (5.13). Substituting this into (5.4) and (5.5), we obtain (5.14). $\blacksquare$
 5.2 Tseng splitting
-Optimistic extragradient (also known as Popov’s or past extragradient) method saves one evaluation of the time-varying operator $V_{k}$ per iteration compared to the standard extragradient method.
+Optimistic extragradient (also known as Popov’s or past extragradient) method saves one evaluation of the time-varying operator ${\mathsf{V}}_{k}$ per iteration compared to the standard extragradient method.
 Yet, it still requires two evaluations of the proximal operator of $G_{k}$ per iteration.
-An attractive modified extragradient scheme, which uses past values of the operator but saves on one evaluation of the proximal operator, is the following optimistic version of Tseng’s Forward-Backward-Forward method [Tsen00, tran2024revisiting]
+An attractive modified extragradient scheme, which uses past values of the operator but saves on one evaluation of the proximal operator, is the following optimistic version of Tseng’s Forward-Backward-Forward method [62, 61]
 $z^{k+1/2}$ $=\operatorname{\textsf{prox}}_{t_{k}G_{k}}(z^{k}-t_{k}{\mathsf{V}}_{k}(z^{k-1/2})),$
@@
 By construction, there exits a $\xi^{k}\in\partial G_{k}(z^{k+1/2})$ such that
-$z^{k+1/2}$ $=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k-1/2})+\xi^{k}),$
-$z^{k+1}$ $=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k-1/2})+\xi^{k})-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}))$
-$=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k}).$
-This further gives us
-$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k+1}-z^{k}+z^{k}-z}^{2}$
-$=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+2\langle z^{k+1}-z^{k},z^{k+1}-z\rangle$
-$=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+\langle z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k})-z^{k},z^{k+1}-z\rangle$
-$=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1}-z^{k+1/2}\rangle$
-$\hskip 17.00024pt-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1/2}-z\rangle.$ (5.21)
+$z^{k+1/2}=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k-1/2})+\xi^{k}),\text{ and }z^{k+1}=z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k}).$
+Eq. (4.8) gives
+$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+2\langle z^{k+1}-z^{k},z^{k+1}-z\rangle$
+$\hskip-42.50061pt=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}+\langle z^{k}-t_{k}({\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k})-z^{k},z^{k+1}-z\rangle$
+$\hskip-42.50061pt=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1}-z^{k+1/2}\rangle$
+$\hskip-25.50037pt-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1/2}-z\rangle.$ (5.15)
 Now observe that $t_{k}({\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k})=z^{k}-z^{k+1}=z^{k}-z^{k+1/2}+t_{k}({\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2})).$ As a consequence, we get
-$2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})$ $+\xi^{k},z^{k+1}-z^{k+1/2}\rangle$
-$=2\langle z^{k}-z^{k+1/2},z^{k+1}-z^{k+1/2}\rangle+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle$
-$=\norm{z^{k}-z^{k+1/2}}^{2}+\norm{z^{k+1}-z^{k+1/2}}^{2}-\norm{z^{k+1}-z^{k}}$
-$\hskip 17.00024pt2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle.$
+$2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1}-z^{k+1/2}\rangle=2\langle z^{k}-z^{k+1/2},z^{k+1}-z^{k+1/2}\rangle$
+$+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle$
+$=\norm{z^{k}-z^{k+1/2}}^{2}+\norm{z^{k+1}-z^{k+1/2}}^{2}-\norm{z^{k+1}-z^{k}}^{2}$
+$+2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle.$
 Combining this estimate with (5.15) gives
-$\norm{z^{k+1}-z}^{2}$ $=\norm{z^{k}-z}^{2}-\norm{z^{k+1}-z^{k}}^{2}-\norm{z^{k}-z^{k+1/2}}^{2}+\norm{z^{k+1}-z^{k+1/2}}^{2}+\norm{z^{k+1}-z^{k}}^{2}$
-$\hskip 17.00024pt2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle$
-$\hskip 17.00024pt-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi_{k},z^{k+1/2}-z\rangle$
-Since $\xi^{k}\in\partial G_{k}(z^{k+1/2})$ it holds that $G_{k}(z^{k+1/2})-G_{k}(z)\leq\langle\xi^{k},z^{k+1/2}-z\rangle$ and therefore
-$-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi^{k},z^{k+1/2}-z\rangle$ $=-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1/2}-z\rangle-2t_{k}\langle\xi^{k},z^{k+1/2}-z\rangle$
-$\leq 2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+2t_{k}(G_{k}(z)-G_{k}(z^{k+1/2}))$ (5.22)
-Moreover, Young’s inequality yields
-$2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle\leq t_{k}^{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}+\norm{z^{k+1}-z^{k+1/2}}.$
-With this estimate, we arrive at (4.1)
-$\frac{1}{2}\norm{z^{k+1}-z}^{2}$ $\leq\frac{1}{2}\norm{z^{k}-z}^{2}-\frac{1}{2}\norm{z^{k+1/2}-z^{k}}^{2}+\frac{t_{k}^{2}}{2}\norm{{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2})}^{2}$ (5.23)
-$\hskip 17.00024pt+t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2}),z-z^{k+1/2}\rangle+t_{k}(G(z)-G(z^{k+1/2})).$
+$\norm{z^{k+1}-z}^{2}=\norm{z^{k}-z}^{2}-\norm{z^{k}-z^{k+1/2}}^{2}-\norm{z^{k+1}-z^{k+1/2}}^{2}$
+$+2t_{k}\langle{\mathsf{V}}_{k}(z^{k-1/2})-{\mathsf{V}}_{k}(z^{k+1/2}),z^{k+1}-z^{k+1/2}\rangle-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})+\xi_{k},z^{k+1/2}-z\rangle.$
+Since $\xi^{k}\in\partial G_{k}(z^{k+1/2})$ the subgradient inequality allows us to continue with
+$\norm{z^{k+1}-z}^{2}$ $\leq\norm{z^{k}-z}^{2}-\norm{z^{k+1/2}-z^{k}}^{2}-\norm{z^{k+1}-z^{k+1/2}}^{2}$
+$-2t_{k}\Psi_{k}(z)-2t_{k}\langle{\mathsf{V}}_{k}(z^{k+1/2})-{\mathsf{V}}_{k}(z^{k-1/2}),z^{k+1}-z^{k+1/2}\rangle.$
+where we have used the notation introduced in (4.2).
+This is exactly (4.10).
 Repeating the same steps as in the rest of Section 4.1, we obtain the same energy bound (4.5).
-Using this energy bound and repeating the same arguments as in Section 4, we obtain the same results as in Theorem 3.1 and Corollary 3.2.
+Using this energy bound and repeating the same arguments as in Section 4, we obtain the same results as reported in Section 3.1 for the optimistic extragradient method.
 6 Numerical Experiments
@@
 6.1 Hierarchical Nash equilibria
-To provide a simple illustration on the practical performance of our method, we’ve tested Algorithm 1 on a simple version of the hierarchical equilibrium problem described Section 1.1.4, taken from [LamSagSIOPT25].
+To provide a simple illustration on the practical performance of our method, we’ve tested Algorithm 1 on a simple version of the hierarchical equilibrium problem described in Section 1.1.3, taken from [38].
 We’ve chosen a deliberately simple example for which the equilibrium set of the game can be computed explicitly and we can monitor the evolution of the inexactness of the algorithm.
@@
 The player-cost functions are
-$h^{\ell}_{1}(y^{1},y^{-1})=0.5(y^{1})^{2}+y^{1}(y^{2}+2y^{3}+y^{4}-100),\;\varphi^{\ell}_{1}(y^{1})=\delta_{[-100,50]}(y^{1}),$
-$h^{\ell}_{2}(y^{2},y^{-2})=0.5(y^{2})^{2}+y^{2}(y^{1}+y^{3}+y^{4}-50),\;\varphi^{\ell}_{2}(y^{2})=\max\{-10(y^{2}-15),0\}+\delta_{[0,50]}(y^{2}),$
-$h^{\ell}_{3}(y^{3},y^{-3})=0.5(y^{3})^{2}+y^{3}(y^{2}+y^{4}-100),\;\varphi^{\ell}_{3}(y^{3})=\delta_{[0,100]}(y^{3})$
-$h^{\ell}_{4}(y^{4},y^{-4})=0.5(y^{4})^{2}+y^{4}(y^{1}+y^{2}+y^{2}+y^{3}-50),\;\varphi^{\ell}_{4}(y^{4})=\delta_{[0,50]}(y^{4}).$
+$h^{\ell}_{1}(y^{1},y^{-1})=0.5(y^{1})^{2}+y^{1}(y^{2}+2y^{3}+y^{4}-100),\;\varphi^{\ell}_{1}(y^{1})=\iota_{[-100,50]}(y^{1}),$
+$h^{\ell}_{2}(y^{2},y^{-2})=0.5(y^{2})^{2}+y^{2}(y^{1}+y^{3}+y^{4}-50),$
+$\varphi^{\ell}_{2}(y^{2})=\max\{-10(y^{2}-15),0\}+\iota_{[0,50]}(y^{2}),$
+$h^{\ell}_{3}(y^{3},y^{-3})=0.5(y^{3})^{2}+y^{3}(y^{2}+y^{4}-100),\;\varphi^{\ell}_{3}(y^{3})=\iota_{[0,100]}(y^{3})$
+$h^{\ell}_{4}(y^{4},y^{-4})=0.5(y^{4})^{2}+y^{4}(y^{1}+y^{2}+y^{2}+y^{3}-50),\;\varphi^{\ell}_{4}(y^{4})=\iota_{[0,50]}(y^{4}).$
 For the upper-level players, we assume
 $h^{u}_{1}(x_{1},x_{2})=(y^{2}-20)^{2}+(y^{4}-50)^{2}+(y^{2}+y^{4})(y^{1}+y^{3}),\varphi^{u}_{1}(x^{1})=0,$
-$h^{u}_{2}(x_{1},x_{2})=(y^{1})^{2}+y^{1}(y^{2}+y^{3})+(y^{3})^{2}+y^{3}(y^{2}+y^{4}),\varphi^{u}_{2}(x^{1})=0.$
+$h^{u}_{2}(x_{1},x_{2})=(y^{1})^{2}+y^{1}(y^{2}+y^{3})+(y^{3})^{2}+y^{3}(y^{2}+y^{4}),\varphi^{u}_{2}(x^{2})=0.$
 One can obtain an explicit expression for the lower-level equilibrium set: $\mathcal{S}_{2}=\{(-50,y^{2},50,50-y^{2}):15\leq y^{2}\leq 50\}$.
 Thus, the unique variational equilibrium of this game is (50,15,50,35)(-50,15,50,35).
-We’ve solved this game problem with Algorithm 1 using the regularization sequence $\sigma_{k}=\frac{1}{(k+3)^{1/2}}$ and constant step size.
-Figure 1 displays the temporal evolution of the strategies of the four players, as well as the distance to the unique variational equilibrium
-Figure 1: Evolution of the strategy profile and error plot.
+We have solved this game problem with Algorithm 1 using the regularization sequence $\sigma_{k}=\frac{1}{(k+3)^{1/2}}$ and constant step size $t_{k}=t$.
+Figure 1 displays the temporal evolution of the strategies of the four players, as well as the distance to the unique variational equilibrium.
+Figure 1: Evolution of the strategy profile and error plot for the hierarchical NEP.
 6.2 Constrained Min-Max problem
@@
 }(x,y)\in\operatorname*{argmin}_{(x^{\prime},y^{\prime})\in\mathcal{X}\times\mathcal{Y}}\frac{1}{2}\norm{{\mathbf{A}}x^{\prime}+{\mathbf{B}}y^{\prime}-c}^{2}_{\mathcal{Z}}$ (6.1)
-where ${\mathbf{A}}:\mathcal{X}\to\mathcal{Z}?$ and ${\mathbf{B}}:\mathcal{Y}\to\mathcal{Z}?$ are bounded linear operators.
+where ${\mathbf{A}}:\mathcal{X}\to\mathcal{Z}$ and ${\mathbf{B}}:\mathcal{Y}\to\mathcal{Z}$ are bounded linear operators.
 The saddle function $\mathcal{L}:\mathcal{Z}\triangleq\mathcal{X}\times\mathcal{Y}\to(-\infty,\infty]$ is of the form
@@
 $\partial g_{1}(x,y)=\partial\varphi(x)\times\partial\psi(y),\;g_{2}(x,y)=0.$
-Using the product space structure $z\triangleq(x,y)\in\mathcal{X}\times\mathcal{Y}\triangleq\mathcal{Z}$, the optimistic extragradient method can be applied to solve this family of equilibrium problems.
-Generalized Absolute value equations (GAVEs).
-The Generalized Absolute value equations (GAVEs) [mangasarian2006absolute] is an important non-smooth NP-hard problem in the form of
+Using the product space structure $z\triangleq(x,y)\in\mathcal{Z}$, the optimistic extragradient method can be applied to solve this family of equilibrium problems.
+Generalized Absolute value equations (GAVEs)
+The GAVEs [43] is an important non-smooth NP-hard problem in the form of
 […]
@@
 […]
-For our numerical investigations, we’ve implemented the following example from [Ebadi24]: Consider the boundary value problem
 […]
-$u(0)=-1,u(1)=0$
+For our numerical investigations, we’ve implemented the following example from [23]. Consider the boundary value problem
+[…]
+$u(0)=-1,u(1)=0,$
 where $f(x)=x^{2}-1$. This second-order ODE has the exact solution
@@
 […]
-Figure 2: Evolution of the approximate solution $u(k)$ after completion of the $k$-th iteration of the Algorithm (left), and the evolution of the residual relative to the true solution (right).
+Figure 2: Evolution of the approximate solution $u^{k}$ after completion of the $k$-th iteration of the Algorithm (left), and the evolution of the residual relative to the true solution (right).
 Using the five-point central difference formula for $u^{\prime\prime}(x_{i})$, we obtain a problem of the form (6.2) with
 $A=\frac{1}{12h^{2}}\begin{pmatrix}20&-6&-4&1&&&&&\\ -16&30&-16&1&&&&&\\ 1&-16&30&-16&1&&&&\\ &\ddots&\ddots&\ddots&&&&&\\ &&&&1&-16&30&-16&1\\ &&&&&1&-16&30&-16\\ &&&&&1&-4&-6&20\\ \end{pmatrix},\;b=\begin{pmatrix}f(x_{1})+\frac{11u_{0}}{12h^{2}}\\ f(x_{2})-\frac{u_{0}}{12h^{2}}\\ f(x_{3})\\ \vdots\\ f(x_{n-2})\\ f(x_{n-1})-\frac{u_{n-1}}{12h^{2}}\\ f(x_{n})+\frac{11u_{n+1}}{12h^{2}}\end{pmatrix}$.
-We’ve reconstructed the solution $u(x)$ using Algorithm 1 with regularization sequence $\sigma_{k}=\frac{1}{(k+3)^{1/2}}$ and constant step size.
-Figure LABEL:fig:GAVEODESolution displays the evolution of approximate solutions obtained with our method measured at the indicated snapshots of the algorithm.
+We’ve reconstructed the solution $u(x)$ using Algorithm 1 with regularization sequence $\sigma_{k}=\frac{1}{(k+3)^{1/2}}$ and constant step size $t_{k}=t$.
+Figure 2 displays the evolution of approximate solutions obtained with our method measured at the indicated snapshots of the algorithm.
 7 Conclusion
-This paper proves seminal results on the iteration complexity of hierarchical hemi-variational inequalities.
-Constructing a suitably defined optimistic extragradient method, we generalize the state-of-the art along several dimensions.
-In particular, we derive complexity guarantees in terms of suitably defined feasibility and optimality gaps.
-Our analysis reveals close connections to geometric conditions imposed on the operator defining the lower level equilibrium problems, conditions which have been so-far dominantly been used in the potential case.
-We belief that our proof strategy is going to be useful for several other problems currently under investigation in the literature.
-In particular, it will be very interesting to derive rates for stochastic formulation of our model template and derive rates in terms of expected gap functions.
-Further interesting directions for future research inlude the consideration of accelerated methods.
-We leave these important extensions to future research.
+This paper proves several results on the convergence rates for solving hierarchical HVI’s.
+Constructing a suitably defined optimistic extragradient method, we advance the state of the art along several dimensions.
+In particular, we derive convergence rates in terms of suitably defined feasibility and optimality gaps.
+Our analysis reveals close connections to geometric conditions imposed on the operator defining the lower level equilibrium problems, conditions which have so-far only been used in the context of hierarchical optimization.
+We believe that our proof strategy is going to be useful for several other problems currently under investigation in the literature.
+In particular, it will be very interesting to derive rates for the stochastic formulation of our model template and derive rates in terms of expected gap functions.
+Further interesting directions for future research include the consideration of accelerated methods.
+We leave these questions as important extensions to future research.
 Acknowledgements
@@
 MST’s research is supported by the Deutsche Forschungsgemeinschaft (DFG) - Projektnummer 556222748 "non-stationary hierarchical minimization".
-We thank Johannes-Carl Schnebel for his help on conducting the numerical experiments, and Enis Chenchene, Radu I.
-Bot, as well as David A.
-Hulett for fruitful discussions on this topic.
+PD’s research is supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689).
+We thank Johannes-Carl Schnebel for his help on conducting the numerical experiments, and Enis Chenchene, Radu I. Bot, as well as David A. Hulett for fruitful discussions on this topic.
 Appendix
-Appendix A The Fitzpatrick function
+Appendix A Auxiliary facts
+Lemma A.1 (Lemma 5.31, [9]).
+Let $(a_{k})_{k\geq 1},(b_{k})_{k\geq 1},(c_{k})_{k\geq 1}$ be nonnegative sequences such that $\sum_{k\geq 1}c_{k}<\infty$ and $a_{k+1}\leq a_{k}-b_{k}+c_{k}.$ Then, $\lim_{k\to\infty}a_{k}$ exists and $\sum_{k\geq 1}b_{k}<\infty$.
+Let $(x_{n})_{n\geq 1}$ be a sequence in $\mathcal{Z}$ and $(\lambda_{n})_{n\geq 1}$ a sequence of positive numbers such that $\sum_{n\geq 1}\lambda_{n}=+\infty$.
+Define the sequence of weighted averages $z_{n}=\frac{1}{\tau_{n}}\sum_{k=1}^{n}\lambda_{k}x_{k}$, where $\tau_{n}\triangleq\sum_{k=1}^{n}\lambda_{k}$.
+The next fact is Lemma 2.3 in [6].
+Lemma A.2 (Opial-Pasty).
+Let $\mathcal{S}$ be a nonempty subset of $\mathcal{Z}$ and assume that $\lim_{n\to\infty}\norm{x_{n}-x}$ exists for every $x\in\mathcal{S}$.
+If every weak cluster point of $(x_{n})_{n\geq 1}$ (respectively $(z_{n})_{n\geq 1}$) lies in $\mathcal{S}$, then $(x_{n})_{n\geq 1}$ (respectively $(z_{n})_{n\geq 1}$) converges weakly to an element of $\mathcal{S}$ as $n\to\infty$.
+Appendix B The Fitzpatrick function
 Let $\mathsf{M}:\mathcal{Z}\to 2^{\mathcal{Z}}$ be a maximally monotone operator.
-The Fitzpatrick function [fitzpatrick1988representing, BauCom16] $\mathcal{F}_{\mathsf{M}}:\mathcal{Z}\to(-\infty,+\infty]$, associated with the operator $\mathsf{M}$, is defined as
+The Fitzpatrick function [26, 9] $\mathcal{F}_{\mathsf{M}}:\mathcal{Z}\to(-\infty,+\infty]$, associated with the operator $\mathsf{M}$, is defined as
 $\mathcal{F}_{\mathsf{M}}(x,u)\triangleq\sup_{(y,v)\in\operatorname{graph}(\mathsf{M})}\{\langle x,v\rangle+\langle y,u\rangle-\langle y,v\rangle\}.$
-Following [Borwein:2016aa], we define the gap function $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)\triangleq\mathcal{F}_{\mathsf{M}}(x,0)$.
-$\operatorname{\mathsf{Gap}}_{\mathsf{M}}$ is convex, and in fact the smallest translation invariant gap function associated with the monotone operator $\mathsf{M}$ [Borwein:2016aa, Theorem 3.1].
+Following [12], we define the gap function $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)\triangleq\mathcal{F}_{\mathsf{M}}(x,0)$.
+$\operatorname{\mathsf{Gap}}_{\mathsf{M}}$ is convex, and in fact the smallest translation invariant gap function associated with the monotone operator $\mathsf{M}$ [12, Theorem 3.1].
 Importantly, this gives the properties $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)\geq 0$ and $\operatorname{\mathsf{Gap}}_{\mathsf{M}}(x)=0$ if and only if $x\in\operatorname{zer}(\mathsf{M})$.
-To make this concept concrete, observe that if $\mathsf{M}={\mathsf{F}}+\operatorname{\mathsf{NC}}_{\mathcal{C}}$, then the above definition of the gap function reduces to the well-known Auslender dual gap function [FacPan03, auslender1976optimisation]
+To make this concept concrete, observe that if $\mathsf{M}={\mathsf{F}}+\operatorname{\mathsf{NC}}_{\mathcal{C}}$, then the above definition of the gap function reduces to the well-known Auslender dual gap function [25, 7]
 $\operatorname{\mathsf{Gap}}_{{\mathsf{F}}+\operatorname{\mathsf{NC}}_{\mathcal{C}}}(x)=\sup_{y\in\mathcal{C}}\langle{\mathsf{F}}(y),x-y\rangle.$
@@
 $\operatorname{\mathsf{Gap}}_{{\mathsf{F}}+\partial g}(x)\leq\sup_{y\in\operatorname{dom}(g)}\langle{\mathsf{F}}(y),x-y\rangle+g(x)-g(y).$
-To the data $({\mathsf{F}},g)$, we thus associate the bifunction $H^{({\mathsf{F}},g)}:\mathcal{Z}\times\mathcal{Z}\to[-\infty,\infty]$ defined as
-$H^{({\mathsf{F}},g)}(x,y)\triangleq\langle{\mathsf{F}}(y),x-y\rangle+g(x)-g(y).$
+Recall the definition (2.5) of the bifunction $H^{({\mathsf{F}},g)}$.
 If ${\mathsf{F}}:\mathcal{Z}\to\mathcal{Z}$ is monotone and continuous on $\operatorname{dom}(g)$, it is easy to verify that $H^{({\mathsf{F}},g)}(x,y)\leq-H^{({\mathsf{F}},g)}(y,x)$ for all $(x,y)\in\operatorname{dom}(g)\times\operatorname{dom}(g)$ (i.e.
-$H^{({\mathsf{F}},g)}$ is a monotone bifunction [iusem2011maximal]).
+$H^{({\mathsf{F}},g)}$ is a monotone bifunction [32]).
 In the structured setting $\operatorname{HVI}({\mathsf{F}},g)$, we obtain the following bounds on the Fitzpatrick function, showing its close connection to the bifunction $H^{({\mathsf{F}},g)}$ and the Auslender dual gap function.
@@
 The function $\varphi^{({\mathsf{F}},g)}(x,u)$ is thus seen as the Fitzpatrick transform of $H^{({\mathsf{F}},g)}$. Second, since the function $x\mapsto H^{({\mathsf{F}},g)}(x,y)$ is convex, we see
-$\varphi^{({\mathsf{F}},g)}(x,u)\leq\sup_{y\in\operatorname{dom}(g)}\{\langle y,u\rangle-H^{({\mathsf{F}},g)}(y,x)\}=\left(H^{({\mathsf{F}},g)}(\bullet,x)\right)^{\ast}(u).$ (A.1)
+$\varphi^{({\mathsf{F}},g)}(x,u)\leq\sup_{y\in\operatorname{dom}(g)}\{\langle y,u\rangle-H^{({\mathsf{F}},g)}(y,x)\}=\left(H^{({\mathsf{F}},g)}(\bullet,x)\right)^{\ast}(u).$ (B.1)
 In particular,
-$\varphi^{({\mathsf{F}},g)}(x,u)\geq\langle x,u\rangle\hskip 17.00024pt\forall x\in\operatorname{dom}(g),$ (A.2)
-$\mathcal{F}_{{\mathsf{F}}+\partial g}(x,0)\leq\sup_{y\in\operatorname{dom}(g)}H^{({\mathsf{F}},g)}(x,y)=\varphi^{({\mathsf{F}},g)}(x,0).$ (A.3)
-Following this notation, we define for a given a subset $\mathcal{C}\subseteq\mathcal{Z}$, the restricted dual gap function as
-$\Theta(x|{\mathsf{F}},g,\mathcal{C})\triangleq\sup_{y\in\mathcal{C}}H^{({\mathsf{F}},g)}(x,y).$ (A.4)
-Appendix B Omitted proofs
+$\varphi^{({\mathsf{F}},g)}(x,u)\geq\langle x,u\rangle\hskip 17.00024pt\forall x\in\operatorname{dom}(g),$ (B.2)
+$\mathcal{F}_{{\mathsf{F}}+\partial g}(x,0)\leq\sup_{y\in\operatorname{dom}(g)}H^{({\mathsf{F}},g)}(x,y)=\varphi^{({\mathsf{F}},g)}(x,0).$ (B.3)
+Following this notation, we define the restricted dual gap function as in (2.4).
+Appendix C Omitted proofs
 Proof of Lemma 2.1.
+The convexity of $x\mapsto\Theta(x|{\mathsf{F}},g,\mathcal{C})$ follows from the fact that it is a supremum of convex functions. Since $\mathcal{C}$ is compact, this function is well-defined.
 The lower bound $\Theta(x|{\mathsf{F}},g,\mathcal{C})\geq 0$ is clear for $x\in\mathcal{C}$.
@@
 $\langle{\mathsf{F}}(y),y-x\rangle+g(y)-g(x)\geq\langle{\mathsf{F}}(x),y-x\rangle+g(y)-g(x)\geq 0\hskip 17.00024pt\forall y\in\mathcal{Z}.$
-Hence,
-$\Theta(x|{\mathsf{F}},g,\mathcal{C})=\sup_{y\in\mathcal{C}}\{\langle{\mathsf{F}}(y),x-y\rangle+g(x)-g(y)\}\leq 0.$
-This, together with already obtained inequality $\Theta(x|{\mathsf{F}},g,\mathcal{C})\geq 0$ implies $\Theta(x|{\mathsf{F}},g,\mathcal{C})=0$.
-Now assume $\Theta(x|{\mathsf{F}},g,\mathcal{C})=0$ for $x\in\mathcal{C}$. Then,
-$\langle{\mathsf{F}}(x^{\prime}),x-x^{\prime}\rangle+g(x)-g(x^{\prime})\leq 0\hskip 17.00024pt\forall x^{\prime}\in\mathcal{C}.$
-Equivalently,
-$\langle{\mathsf{F}}(x^{\prime}),x^{\prime}-x\rangle+g(x^{\prime})-g(x)\geq 0\hskip 17.00024pt\forall x^{\prime}\in\mathcal{C}.$
-This means $x\in\mathcal{C}$ is a Minty (weak) solution to the hemi-variational inequality with the data $({\mathsf{F}},g)$ over the set $\mathcal{C}$.
+Hence, $\Theta(x|{\mathsf{F}},g,\mathcal{C})=\sup_{y\in\mathcal{C}}\{\langle{\mathsf{F}}(y),x-y\rangle+g(x)-g(y)\}\leq 0$.
+This, together with already proved inequality $\Theta(x|{\mathsf{F}},g,\mathcal{C})\geq 0$ implies $\Theta(x|{\mathsf{F}},g,\mathcal{C})=0$.
+Now assume $\Theta(x|{\mathsf{F}},g,\mathcal{C})=0$ for $x\in\mathcal{C}$.
+Then, $\langle{\mathsf{F}}(x^{\prime}),x-x^{\prime}\rangle+g(x)-g(x^{\prime})\leq 0$ for all $x^{\prime}\in\mathcal{C}.$ Equivalently, $\langle{\mathsf{F}}(x^{\prime}),x^{\prime}-x\rangle+g(x^{\prime})-g(x)\geq 0$ for all $x^{\prime}\in\mathcal{C}.$ This means $x\in\mathcal{C}$ is a Minty (weak) solution to the $\operatorname{HVI}({\mathsf{F}},g)$ over the set $\mathcal{C}$.
 Let $w\in\mathcal{C}$ be arbitrary and consider v=tw+(1t)xv=tw+(1-t)x for $t\in[0,1]$.
@@
 Dividing both sides by $t$ and then letting $t\to 0^{+}$, the weak continuity of $\mathsf{F}$} implies
-$0\leq\langle{\mathsf{F}}(x),w-x\rangle+g(w)-g(x).$ (B.1)
+$0\leq\langle{\mathsf{F}}(x),w-x\rangle+g(w)-g(x).$ (C.1)
 Since $w\in\mathcal{C}$ has been chosen arbitrarily, we see that $x$ is a Stampacchia (strong) solution of the hemivariational inequality with the additional constraint $x\in\mathcal{C}$.
-We now show that if $\mathcal{C}\subset\operatorname{dom}(g)$, and there exists $\epsilon>0$ such that $\mathcal{U}\triangleq B(x,\epsilon)\cap\mathcal{C}=B(x,\epsilon)\cap\operatorname{dom}(g)$, then (B.1) holds also for $w\in\operatorname{dom}(g)$.
-To that end, assume to the contrary, there exits $z\in\operatorname{dom}(g)$ for which
-$\langle{\mathsf{F}}(x),z-x\rangle+g(z)-g(x)<0.$
-Then, by convexity of $\operatorname{dom}(g)$ and definition of $\mathcal{U}$ there exists $\lambda>0$ small enough for which $w=x+\lambda(z-x)\in\mathcal{U}$ and by convexity
+[…]
 $0\leq\langle{\mathsf{F}}(x),w-x\rangle+g(w)-g(x)\leq\lambda(\langle F(x),z-x\rangle+g(z)-g(x))<0,$
 thus arriving at a contradiction.
-Hence, $\langle F(x),z-x\rangle+g(z)-g(x)\geq 0$ for any $z\in\operatorname{dom}(g)$ and moreover for any $z\in\mathcal{Z}$, which finishes the proof that $x$ is a solution to (2.3).
+Hence, $\langle F(x),z-x\rangle+g(z)-g(x)\geq 0$ for any $z\in\operatorname{dom}(g)$ and moreover for any $z\in\mathcal{Z}$, which finishes the proof that $x$ solves (2.3).
 $\blacksquare$
 Proof of Proposition 2.3.
-Pick $z\in\operatorname{dom}(g)$.
-Without loss of generality, assume $z\notin\mathcal{S}$.
-Let $\bar{z}=\Pi_{\mathcal{S}}(z)$, so that there exists $p\in\operatorname{\mathsf{NC}}_{\mathcal{S}}(\bar{z})$ satisfying $p=z-\bar{z}$.
-Hence, $p^{\ast}\triangleq\frac{z-\bar{z}}{\norm{z-\bar{z}}}$ is a unit norm element of $\operatorname{\mathsf{NC}}_{\mathcal{S}}(\bar{z})$ satisfying
-$\langle p^{\ast},z-\bar{z}\rangle=\norm{z-\bar{z}}.$
 […]
@@
 By the convex sugradient inequality, this implies
-$\langle{\mathsf{F}}_{1}(z^{*}),z-z^{*}\rangle+g_{1}(z)-g_{1}(z^{*})\geq\langle-p^{*},z-z^{*}\rangle\hskip 17.00024pt\forall z\in\mathcal{Z}.$ (B.2)
-Take $z\in\mathcal{Z}$ and let $\hat{z}=\Pi_{\mathcal{S}_{2}}(z)$ be the orthogonal projector of $z$ onto $\mathcal{S}_{2}$. Thus, $\langle p^{*},\hat{z}-z^{*}\rangle\leq 0$, resulting in
-$\langle p^{*},\hat{z}-z\rangle\leq\langle p^{*},z^{*}-z\rangle,$
-which implies when combined with (B.2)
-$\langle{\mathsf{F}}_{1}(z^{*}),z-z^{*}\rangle+$ $g_{1}(z)-g_{1}(z^{*})\geq\langle p^{*},\hat{z}-z\rangle$
-$\geq-\norm{p^{*}}\cdot\norm{\hat{z}-z}=-\norm{p^{*}}\operatorname{dist}(z,\mathcal{S}_{2}).$
-Hence, for all compact $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})$ with $z^{*}\in\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$, we conclude $\Theta_{\rm Opt}(z|\mathcal{U}_{1}\cap\mathcal{S}_{1})\geq-B_{\mathcal{U}_{1}}\operatorname{dist}(z,\mathcal{S}_{2})$, where $B_{\mathcal{U}_{1}}=\sup_{\mathcal{U}_{1}\cap\mathcal{S}_{1}}\norm{p^{*}}$.
+$\langle{\mathsf{F}}_{1}(z^{*}),z-z^{*}\rangle+g_{1}(z)-g_{1}(z^{*})\geq\langle-p^{*},z-z^{*}\rangle\hskip 17.00024pt\forall z\in\mathcal{Z}.$ (C.2)
+Take $z\in\mathcal{Z}$ and let $\hat{z}=\Pi_{\mathcal{S}_{2}}(z)$.
+Thus, $\langle p^{*},\hat{z}-z^{*}\rangle\leq 0$, resulting in $\langle p^{*},\hat{z}-z\rangle\leq\langle p^{*},z^{*}-z\rangle,$ which implies when combined with (C.2)
+$\langle{\mathsf{F}}_{1}(z^{*}),z-z^{*}\rangle+$ $g_{1}(z)-g_{1}(z^{*})\geq\langle p^{*},\hat{z}-z\rangle\geq-\norm{p^{*}}\cdot\norm{\hat{z}-z}=-\norm{p^{*}}\operatorname{dist}(z,\mathcal{S}_{2}).$
+Hence, for all compact $\mathcal{U}_{1}\subset\operatorname{dom}(g_{1})$ with $z^{*}\in\mathcal{U}_{1}\cap\mathcal{S}_{1}\neq\varnothing$ and corresponding $p^{*}\in\operatorname{\mathsf{NC}}_{\mathcal{S}_{2}}(z^{*})$, we conclude $\Theta_{\rm Opt}(z|\mathcal{U}_{1}\cap\mathcal{S}_{1})\geq-B_{\mathcal{U}_{1}}\operatorname{dist}(z,\mathcal{S}_{2})$, where $B_{\mathcal{U}_{1}}\triangleq\norm{p^{*}}$.
 To show (2.10), we can directly use Definition 2.4, to conclude